### v2

# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

Abstract
CLIP is one of the most important multimodal foundational models today, aligning visual and textual signals into a shared feature space using a simple contrastive learning loss on large-scale image-text pairs. What powers CLIP’s capabilities? The rich supervision signals provided by natural language — the carrier of human knowledge — shape a powerful cross-modal representation space. As a result, CLIP supports a variety of tasks, including zero-shot classification, detection, segmentation, and cross-modal retrieval, significantly influencing the entire multimodal domain. However, with the rapid advancements in large language models (LLMs) like GPT-4 and LLaMA, the boundaries of language comprehension and generation are continually being pushed. This raises an intriguing question: can the capabilities of LLMs be harnessed to further improve multimodal representation learning? The potential benefits of incorporating LLMs into CLIP are clear. LLMs’ strong textual understanding can fundamentally improve CLIP’s ability to handle image captions, drastically enhancing its ability to process long and complex texts — a well-known limitation of vanilla CLIP. Moreover, LLMs are trained on a vast corpus of text, possessing open-world knowledge. This allows them to expand on caption information during training, increasing the efficiency of the learning process. However, realizing this potential is challenging. Despite LLMs’ powerful internal comprehension, their autoregressive nature hides this capability within the model, leading to output features with poor discriminability. Our experiments show that directly integrating LLMs into CLIP results in catastrophic performance drops. In this paper, we propose LLM2CLIP, a novel approach that embraces the power of LLMs to unlock CLIP’s potential. By fine-tuning the LLM in the caption space with contrastive learning, we extract its textual capabilities into the output embeddings, significantly improving the output layer’s textual discriminability. We then design an efficient training process where the fine-tuned LLM acts as a powerful teacher for CLIP’s visual encoder. Thanks to the LLM’s presence, we can now incorporate longer and more complex captions without being restricted by vanilla CLIP text encoder’s context window and ability limitations. Our experiments demonstrate that this approach brings substantial improvements in cross-modal tasks. Our method directly boosted the performance of the previously SOTA EVA02 model by 16.5% on both long-text and short-text retrieval tasks, transforming a CLIP model trained solely on English data into a state-of-the-art cross-lingual model. Moreover, when integrated into multimodal training with models like Llava 1.5, it consistently outperformed CLIP across nearly all benchmarks, demonstrating comprehensive performance improvements.
Figure 1: LLM2CLIP Overview. After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training. We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision. Experimental results demonstrate that LLM2CLIP can make any SOTA CLIP model even more SOTA ever.
1 Introduction
CLIP (Radford et al., 2021) is one of the most important multimodal foundational models today. It aligns vision and language signals into a shared feature space by employing a simple contrastive learning loss on large-scale image-text pairs. As a retriever, CLIP supports a wide range of tasks, including zero-shot classification (Qian & Hu, 2024), detection (Lin & Gong, 2023), segmentation (Zhou et al., 2023), and image-text retrieval (Lülf et al., 2024; Koukounas et al., 2024). As a feature extractor, it has become dominant in virtually all cross-modal representation tasks, such as image understanding, video understanding, and text-to-image/video generation. For instance, works like LLaVA (Liu et al., 2023) and Qwen-VL (Bai et al., 2023) leverage CLIP as a feature extractor to obtain visual features for text models, while models like Stable Diffusion (Rombach et al., 2021) and DALL·E 2 (Ramesh et al., 2022) use CLIP’s text encoder to extract textual features for visual models.
What makes CLIP so powerful, particularly as a vision encoder? The core of its strength lies in its unprecedented ability to align visual pretraining with natural language — the carrier of human knowledge. Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text. This alignment with language is what sets CLIP apart and unlocks its vast potential. However, since CLIP’s introduction, large language models (LLMs) have advanced significantly. Models like GPT-4 (Achiam et al., 2023) and Llama (Dubey et al., 2024) now demonstrate remarkable language capabilities, yet these advancements have not translated to corresponding improvements in visual representation learning. This prompts the question: can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
The potential benefits of incorporating LLMs into CLIP are clear. LLMs’ strong textual understanding can fundamentally improve CLIP’s ability to handle image captions, drastically enhancing its ability to process long and complex texts — a well-known limitation of vanilla CLIP. Moreover, LLMs are trained on a vast corpus of text, possessing open-world knowledge. This allows them to expand on caption information during training, increasing the efficiency of the learning process.
In this work, we aim to leverage large language models (LLMs) to enable CLIP to learn more powerful, fine-grained, and rich visual representations. Currently, CLIP is often criticized for its bag-of-words-like perception and the limitations of its text encoder, which suffers from a constrained model size, limited context length, and is trained predominantly on image captioning data, lacking exposure to diverse world corpora. A natural approach would be to replace CLIP’s text encoder with an LLM that embeds rich human knowledge. However, this presents significant challenges. In the cross-modal contrastive learning framework employed by CLIP, the text encoder functions as a set of knowledge anchors in the shared latent space, guiding the alignment of the vision encoder with human knowledge of the physical world. The structure, richness, and discriminability of these knowledge anchors are critical to the visual model’s effectiveness. In contrast, LLMs are primarily designed to predict the next word rather than generate explicit representations of the knowledge they contain. Their textual comprehension abilities and open-world knowledge are latent within the model, rather than present in the output embeddings, making them difficult to utilize in the same explicit manner as CLIP’s text encoder. As a result, using LLMs as a text encoder may not produce linearly separable features, which are crucial for effective feature alignment.
Table 1: Comparison of top-1 Caption Retrieval Accuracy (CRA) for various language models in MS COCO 5K testing set.
Language Model CRA
CLIP-L/14 25.3
EVA02-L/14 27.1
Llama3-8B 5.2
Llama3.2-1B 5.6
Llama3-8B-CC 29.2
Llama3.2-1B-CC 28.0
Figure 2: Real examples of top-1 results from the caption-to-caption retrieval experiment. Before fine-tuning, Llama3’s results were often completely unrelated.
To validate our hypothesis, we designed a caption-to-caption retrieval experiment, as shown in Table 1 and Figure 2. Each image in the MS-COCO dataset has five human-annotated captions. We selected the first two captions as positive samples and performed retrieval across the entire validation set. Using the caption retrieval accuracy (CRA), we evaluated the text model’s ability to differentiate between captions, helping us determine which language model is better suited for CLIP. We found that Llama-3 8B achieved only 18.4% top-1 accuracy, while the standard CLIP-ViT-L reached 66.0% top-1 accuracy. As illustrated in Figure 2, the top-1 caption retrieved by original Llama-3 can be entirely unrelated to the query caption, clearly obstructing effective CLIP learning. Therefore, directly using an LLM to guide CLIP’s visual encoder training is highly constrained.
We believe that enhancing the discriminative power of LLM output tokens through fine-tuning is vital for the success of our proposed approach, allowing the latent capabilities of LLMs to surface. Encouragingly, we found that this can be achieved very efficiently. Specifically, we designed a caption contractive (CC) fine-tuning strategy, applying lightweight fine-tuning to the output tokens of Llama-3 8B using LoRA on the CC3M (Sharma et al., 2018) image captioning dataset. The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions. We utilized a supervised SimCSE (Gao et al., 2021; BehnamGhader et al., 2024) contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner (Chen et al., 2023) were treated as positive pairs, pulling them closer. In contrast, all other captions formed a negative sample set that the model learned to push away. Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 18.4% to 73%, which is a 7% improvement over the original CLIP-ViT-L text encoder. This successful fine-tuning process enables us to more effectively harness the open-world capabilities of LLMs for CLIP training.
In a nutshell, we present LLM2CLIP, a novel approach for enhancing visual representation learning through the integration of large language models (LLMs) as shown in Figure 1. This method takes a straightforward yet audacious step by replacing the original CLIP text encoder and augmenting the CLIP visual encoder with the vast knowledge embedded in LLMs. We have identified key obstacles associated with this innovative idea and proposed a cost-effective fine-tuning strategy to overcome them. Our experiments demonstrate that leveraging LLMs as teachers for CLIP training yields substantial improvements, with LLM2CLIP significantly outperforming state-of-the-art pre-trained CLIP models. Our method increased the performance of the previously SOTA EVA02 model by 16.5% on both long-text and short-text retrieval tasks, transforming a CLIP model trained solely on English data into a state-of-the-art cross-lingual model. Furthermore, when incorporated into multimodal model training, such as with Llava 1.5, it consistently achieved comprehensive improvements over EVA02 across nearly all benchmarks. Additionally, the efficient training method proposed by LLM2CLIP ensures that the training cost is nearly identical to fine-tuning the original CLIP. We have also demonstrated that using more powerful language models and larger training datasets can further boost LLM2CLIP’s performance, showcasing the immense potential of our approach. These promising outcomes affirm that we have successfully transformed CLIP into a more general-purpose foundational model. The enhanced LLM2CLIP model possesses richer knowledge and exhibits a remarkable capacity for distinguishing fine-grained and complex long-text semantics. This advancement not only broadens the range of supported downstream tasks but also propels progress across the entire vision domain.
2 Related Works
Contrastive Language-Image Pre-training (CLIP).
CLIP (Radford et al., 2021) is one of the most important multimodal models today. By training on web-scale data, image captions serve as rich, fine-grained supervision signals for learning representations that align closely with image, freeing CLIP from the limitations of manually defined labels. The richness of the captions endows CLIP with many remarkable capabilities; besides zero-shot image-text retrieval, CLIP features can even achieve region-level correspondence, supporting zero-shot detection and segmentation tasks. Furthermore, CLIP is the most widely used encoder in current image and video understanding tasks. In Tong et al. (2024)’s analysis, it stands out among various visual encoders in MLLMs training as the most significant model, demonstrating its strong capabilities. For generative tasks, CLIP is equally foundational. Its text and image encoders are used to provide text control (Rombach et al., 2021; Ramesh et al., 2022) and image control signals (Blattmann et al., 2023; Zhang et al., 2023) for conditioning generative models. Additionally, the CLIP score (Hessel et al., 2021) has become a crucial metric for assessing the relevance of generated images to text. While the original CLIP only utilized image and text modalities, there have been massive extensions of CLIP to incorporate various other modalities (Dai et al., 2024; He et al., 2023; Moon et al., 2022; Zhang et al., 2022). CLIP definitely has a significant impact on the field of vision and multimodal, and its influence is still growing.
CLIP meets Stronger Language Models.
Several works have explored the integration of LLMs into CLIP. JinaCLIP (Koukounas et al., 2024) employed Jina-embeddings-v2 (Günther et al., 2023) as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts. Though achieving similar visual performance to EVA-CLIP (Sun et al., 2023), its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs. T5-V (Jiang et al., 2024) took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B (Li et al., 2024a) while freezing the ViT gradient to focus on adjusting the MLLM output. However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP. MATE (Jang et al., 2024) designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs. They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks. However, they did not recognize the critical issue we propose in this paper: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training. Our work aims to thoroughly explore the impact of LLMs on CLIP as a foundational model, going beyond the influence on long-text retrieval.
CLIP meets Longer Captions.
It has been widely recognized that the quality of CLIP’s text embedding is coarse and limited to only 77 tokens. Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly. For instance, DCI (Urbanek et al., 2024) hired annotators to expand captions based on SAM (Kirillov et al., 2023)’s target hints. LaCLIP (Fan et al., 2024) adopted ChatGPT, Bard, and human rewriting. DreamLIP (Zheng et al., 2024) leveraged ShareCaptioner (Chen et al., 2023) and InstructBLIP to augment 30M captions. Recap-DataComp-1B (Li et al., 2024b) used Llama3-trained LLAVA1.5 to extend 1B captions. In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones (Urbanek et al., 2024), splitting captions into multiple segments (Fan et al., 2024; Zheng et al., 2024), or finetuning model’s positional encoding to support longer token inputs (Zhang et al., 2024). LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions. It not only resolves the limition of token length but also allows for better understanding using LLMs’ open-world knowledge. As a result, our model achieves a breakthrough in performance.
3 Methods
The contributions of our methodology are threefold: First, we designed experiments to analyze the key reason preventing LLMs from directly participating in multimodal representation learning — the weak discriminability of their output features. Second, we introduced the caption contrastive fine-tuning method, which significantly improves feature discriminability. Third, we developed the LLM2CLIP training framework, which has been proven to be an efficient and effective method for leveraging LLMs to deliver substantial performance improvements to pretrained CLIP models.
3.1 Native LLMs are Ineffective Text Encoders for CLIP
Large Language Models (LLMs), such as Llama-3, are trained on massive world corpora through auto-regression, equipping them with open-world knowledge and enabling them to perform various tasks. We initially hoped to leverage the capabilities of LLMs to directly retrain a CLIP model.
Although LLMs exhibit strong text comprehension, they are difficult to use directly as text embedding models. This is because their knowledge is encapsulated within the model, and their output features are heavily skewed towards individual word predictions. As generative models, they are not trained to ensure good linear separability of output features, making them less effective when used to interpret captions for CLIP. As highlighted by  Chen et al. (2022), cross-modal contrastive learning in CLIP requires that each modality possesses strong internal discriminability.
To evaluate the effectiveness of various language models in terms of text discriminability, and to test whether the native output features of LLMs, as hypothesized, struggle to distinguish image captions, we introduce a new metric: the MS COCO Caption Retrieve Accracy (CRA). MS COCO is a widely used multimodal dataset, containing over 330K images, each with five captions. These captions are written by different human annotators and provide diverse descriptions for each image. In our evaluation on the MS COCO validation set, we use only the first two captions of each image and treat the captions of the same image as positive pairs, while all other captions serve as negative samples. We then perform caption-to-caption retrieval and assess Top-1 accuracy using different language models, defining the result as their CRA score. Higher CRA scores indicate better discriminability of the language models on image captions.
As shown in Table 1, using a pure LLM results in a CRA score of only 18.4%, indicating that the majority of captions cannot be well-separated in the output space. In fact, captions with similar distances may be entirely unrelated, as illustrated in Figure 2. However, the text encoder from the original state-of-the-art CLIP model achieves a CRA score of 66%, proving the inadequacy of native LLM output features in caption discriminability. Consequently, it is challenging to apply LLMs directly in CLIP model training. Subsequent experiments in Table 6 also confirm that replacing CLIP’s text encoder and the corresponding ViT with Llama-3 8B for contrastive learning significantly underperforms the original CLIP.
3.2 A Crucial Lesson: LLM Learning for Image Caption Discrimination
In this section, we aim to fine-tune the LLM’s token outputs to better capture features that can distinguish between image captions.
The process of improving the discriminability of LLM output features on caption text is quite straightforward: we need the distances between captions of the same image to be closer, and those of different images to be further apart. Therefore, we apply a caption contrastive (CC) fine-tuning to the LLM’s output features, treating different captions of the same image as positive samples and the rest of the captions as negative samples.
To obtain enough varied descriptions for the same image, we use the ShareCaptioner (Zheng et al., 2024; Chen et al., 2023) modified CC-3M (Sharma et al., 2018) dataset, which provides both original captions and augmented dense captions for each image. These can be treated as positive pairs. We followed the training methodology of LLM2Vec (BehnamGhader et al., 2024), first expanding the LLM’s attention mechanism to bidirectional attention, and employing Masked Next Token Prediction (MNTP) for initialization to achieve better results. Specifically:
First, we transform the LLM’s causal attention mechanism into bidirectional attention, as we no longer need it to retain generative capabilities but rather function as an encoder. Since autoregressive training is no longer required, switching to bidirectional attention improves its ability to capture contextual information. Second, we employ MNTP to train the newly added bidirectional attention mechanism, providing a strong initialization. For a given sequence of $N$ tokens, we mask a subset and predict their values, similar to BERT (Devlin et al., 2018). However, unlike BERT, we adapt this process to fit the nature of LLMs by predicting tokens just before the masked token. We train on both image captions and pure text with equal weighting. In addition to CC-3M, we also use the Wikitext-103 (Merity et al., 2016) dataset to preserve the LLM’s text capabilities and prevent divergence from its original strengths. Finally, we perform the actual caption contrastive fine-tuning, using a supervised SimCSE loss to bring captions of the same image closer together and push apart captions of different images. We use two prompt templates: “Given a caption, retrieve a detailed relevant caption” and “Given a detailed caption, retrieve a short relevant caption,” which are prepended to the query (either original or dense caption) to retrieve the corresponding dense or original caption. Similarly, we use a 1.5M-paired general text dataset curated from  Springer et al. (2024) to maintain strong performance in pure language tasks. All training is efficiently conducted using LoRA and is completed in just one epoch, ensuring low computational cost.
A remarkable result followed: after fine-tuning the LLM as a caption encoder for only one epoch on CC-3M, the CRA score for Llama-3 8B jumped from 18.4% to 73.0%, surpassing the previous state-of-the-art CLIP and EVA models’ text encoders trained on the massive Laion-2B (Schuhmann et al., 2022) and COYO-700M (Byeon et al., 2022) image-text datasets. As shown in Table 6, subsequent experiments demonstrated that after CC fine-tuning, the LLM finally unleashed its powerful capabilities, significantly boosting the performance of the previously state-of-the-art CLIP model, in stark contrast to the results without CC fine-tuning. This breakthrough uncovers the potential of LLMs in CLIP training and removes a major obstacle in leveraging LLMs to advance vision foundation models.
3.3 LLM2CLIP: Makes SOTA CLIP more SOTA ever
With the modifications to the LLM discussed earlier, we have now obtained a super text encoder that is well-suited for CLIP training. The next step is to use this LLM in conjunction with the pretrained state-of-the-art CLIP visual encoder to reconstruct a more powerful cross-modal feature space.
As shown in  Figure 1, in the LLM2CLIP training phase, we freeze the gradients of the LLM to preserve its inherent capabilities, for two primary reasons. First, this significantly reduces the computational cost and memory footprint of fine-tuning. CLIP training requires a very large batch size to maintain the effectiveness of negative samples. Allocating memory to the LLM could compromise CLIP’s performance. Second, by freezing the LLM, we ensure that the open-world knowledge it has acquired from large-scale corpora remains intact during the multimodal alignment process.
To compensate for the frozen LLM, and inspired by methods like FuseMix (Vouitsis et al., 2023) and APE (Rosenfeld et al., 2022), we introduce several new linear layers after the LLM as adapters. These layers serve as learnable parameters to improve the alignment between the LLM and CLIP visual encoder. Following the original design of CLIP, we also employ a projector layer to align the dimensions of the two encoders, facilitating the use of CLIP loss for training.
With this powerful LLM-based super text encoder, we achieve a substantial qualitative leap in CLIP’s language comprehension capabilities. The LLM’s open-world knowledge allows the CLIP visual encoder to learn more structured and globally informed visual representations that are aligned with human knowledge. Moreover, this approach enables us to fully utilize high-quality, long, and dense caption datasets without requiring any special architectural adjustments, which previous works like DCI, DreamLip, and Recaption struggled to effectively leverage. As shown in  Table 2, LLM2CLIP makes any existing SOTA CLIP model even more SOTA, significantly surpassing the performance of previous one.
3.4 Overview and Efficiency Discussion
We propose LLM2CLIP as a method that efficiently incorporates large language models (LLMs) into CLIP training, leveraging the capabilities of LLMs to make cross-modal representation learning significantly more powerful. In our experiments, We evaluated large language models, including Llama with 1B and 8B parameters, as well as Mistral-Nemo with 12B parameters.It might seem that incorporating such large LLMs would greatly increase the computational burden of training CLIP, especially since CLIP itself is computationally expensive, requiring a large batch size. However, our proposed LLM2CLIP is remarkably lightweight. The training overhead is nearly identical to fine-tuning the original CLIP model, with minimal additional cost, yet the LLM provides much stronger supervision.
Here, we highlight some of the design details that significantly improve training efficiency: 1). During the caption contrastive fine-tuning stage, we employ LoRA training for the LLM. Even for a 12B LLM, training with a batch size of 512 only requires around 70GB of GPU memory, making it feasible to run on a single 80GB 8 A100 GPU node. 2). In the LLM2CLIP stage, we freeze the LLM’s gradients and only train the learnable adapter, CLIP’s original Vision encoder, and two projectors. The additional trainable parameters are roughly equivalent to those in the original CLIP, minimizing the overhead. To further reduce the inference cost of using LLMs, we pre-extract all the text features from the training data and store them in memory. This ensures that, even though LLMs provide powerful textual supervision, the memory and computational costs during training remain nearly identical to those of standard CLIP training.
For instance, when we trained LLM2CLIP using a Mistral-Nemo [^1] https://mistral.ai/news/mistral-nemo/ 12B model integrated with the commonly used EVA ViT-L/14-224, with a batch size of 4096 on 8 H100 GPUs, the memory usage per GPU was only 30GB, and the entire training process took just 9 hours. Despite this efficient training cost, LLM2CLIP brought transformative improvements in downstream tasks such as long and short text retrieval, cross-language retrieval, and LLAVA training.
4 EXPERIMENTS
4.1 Experiment Setting
4.1.1 Caption Contrastive Fine-tuning
Training Dataset.
We utilized the ShareCaptioner-modified CC-3M dataset (Zheng et al., 2024; Chen et al., 2023), which provides both original captions and augmented dense captions for each image, for contrastive learning. For the Masked Next Token Prediction and caption contrastive fine-tuning stages, we employed the Wikitext-103 dataset (Merity et al., 2016) and the E5 dataset from  Springer et al. (2024) to ensure that the general text domain was not excessively biased.
Training Setting.
We trained the language model using LoRA, applying lightweight training with 1 epoch on all datasets. We adopted the average of all LLM output tokens as the text global embedding for a caption. All training parameters follow the design of  BehnamGhader et al. (2024).
4.1.2 LLM2CLIP Fine-tuning
Training Dataset.
To compare different configurations, we designed three experimental settings based on dataset size: LLM2CLIP-3M: This configuration only uses the CC-3M dataset for training, representing our lightweight setting. LLM2CLIP-15M: This configuration uses both CC-3M and CC-12M datasets, which is our default setting. All LLM2CLIP models without a specific dataset size label are trained with this 15M data. LLM2CLIP-60M: This configuration scales up the dataset, combining CC-3M, CC-12M, YFCC-15M, and a randomly selected 30M subset from Recaption-1B (Li et al., 2024b). All data used are dense captions rewritten by multimodal large language models (MLLMs). The CC-3M, CC-12M, and YFCC datasets are sourced from  Zheng et al. (2024) using the ShareCaptioner for rewriting, while Recaption data was rewritten using Llama3-LLAVA1.5. We primarily used a mix of original captions and dense captions, with the default mixing ratio being 1:1. LLM2CLIP-S represents a setting where only original data is used for training, matching the original CLIP pretraining distribution to analyze the model’s benefits separately.
Training Setting.
All experiments with different datasets were conducted for 4 epochs. We froze the LLM gradients by default. During training, we pre-extracted the text features from the captions using the LLM and stored them in memory to avoid repeated LLM inference, reducing additional computational overhead. We trained the original CLIP vision encoder, the LLM’s learnable adapter, and the projectors for both encoders. Training on the 3M, 15M, and 60M datasets for CLIP ViT-L/14-224 required only 30G per GPU memory and 2, 9, and 45 hours on 8 H100 GPUs, respectively.
4.1.3 Evaluation
Evaluation Dataset.
For short-text datasets, we used the commonly available COCO 2014 5k test set and the Flickr 1k test set as our test datasets. For long-text datasets, we used the 1K subset of the ShareGPT4V (Chen et al., 2023) dataset and the Urban1k (Zhang et al., 2024) dataset, both provided by LongCLIP, along with the DOCCI (Onoe et al., 2024) dataset. The ShareGPT4V-1M dataset was generated using annotations from GPT-4V and ShareCaptioner, with images from Laion, CC, SBU (Ordonez et al., 2011) and MS COCO. We used a randomly selected 1K subset of this dataset. Urban1k consists of captions generated by GPT-4V for 1,000 busy urban view images from Visual Genome. Each caption is a long and complete sentence that describes the image, including types, colors, and relative locations of various attributes. The model can only successfully match the images with the correct captions if it accurately understands and models the detailed attributes in both modalities. The DOCCI dataset contains approximately 1.5K high-resolution images with detailed human-annotated descriptive captions. DOCCI is divided into a training set of 9.6K pairs and a test set of 5.1K pairs. We used the test set for image-lengthy caption retrieval experiments. For the Chinese retrieval tasks, we used the FlickrCN (Lan et al., 2017) and CNCOCO (Li et al., 2018) datasets, which are translated versions of Flickr30K and MS COCO-1K, respectively. These datasets were tested in Chinese. For Llava (Liu et al., 2023) training, we used the standard training and test sets from LLAVA 1.5.
Evaluation Setting.
The experiments in this paper primarily focus on how LLM2CLIP enhances the performance of widely used vanilla CLIP across various dimensions. We mainly compare our approach with the EVA (Fang et al., 2023) and OpenAI CLIP models as baselines, as they are the most widely used SOTA vision encoders in the open-source community. The CLIP models we use for comparison are ViT-B/16222https://huggingface.co/openai/clip-vit-base-patch16, ViT-L/14333https://huggingface.co/openai/clip-vit-large-patch14, and ViT-L/14-336444https://huggingface.co/openai/clip-vit-large-patch14-336. For EVA02 models, we use EVA02 ViT-B/16555https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02_CLIP_B_psz16_s8B.pt, EVA02 ViT-L/14666https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02_CLIP_L_psz14_s4B.pt, and EVA02 ViT-L/14-336777https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02_CLIP_L_336_psz14_s6B.pt, making up a total of six models. For language models, we experimented with four different models: Jina-Embeddings-V2888https://huggingface.co/jinaai/jina-embeddings-v2-base-en, and three popular LLMs from the Llama3 (Dubey et al., 2024) family, namely Llama 3.2 1B, Llama 3 8B, and Mistral-Nemo 12B. To avoid any misunderstanding, the experiments in this paper, unless otherwise noted, use EVA02 ViT-L/14 as the default baseline for comparison. The default language model used is LLaMA 3 8B, trained on a dataset of 15M setting mentioned above.
4.2 Main Results
Directly Replacing CLIP’s Text Encoder with a Vanilla LLM is Harmful.
As shown in the experiments in Table 4, directly replacing the text encoder of EVA02 ViT-L/14 with Llama3-8B significantly degrades retrieval performance. For example, performance on the DOCCI benchmark nearly halves, dropping from 75.0/73.4 to 51.7/50.6. Additionally, the CRA score of the vanilla Llama3-8B is particularly low, indicating that its output features exhibit very poor discriminability for captions. These results clearly show that such an approach imposes a significant burden on CLIP’s learning process.
Table 2: Systematic Comparison Experiment Demonstrating the Performance Improvements of LLM2CLIP.
Methods Flickr30k COCO ShareGPT4V Urban-1k DOCCI
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
ViT-B/16
ALIGN 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1
BLIP 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5
Jina-CLIP 80.6 67.4 55.6 41.1 - - 87.7 88.0 78.7 80.0
Long-CLIP 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4
CLIP 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1
+LLM2CLIP 89.2 78.1 62.2 48.7 98.1 97.4 86.1 90.0 84.1 85.0
EVA02 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0
+LLM2CLIP 88.5 78.0 63.6 49.8 98.0 98.1 84.7 89.7 85.5 86.8
ViT-L/14
Long-CLIP 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6
CLIP 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8
+LLM2CLIP 92.6 81.7 64.9 52.5 98.4 98.4 87.6 92.0 87.6 88.7
EVA02 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0
+LLM2CLIP-3M 89.6 77.3 59.7 48.0 98.3 98.6 87.1 91.1 84.9 87.8
+LLM2CLIP 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4
+LLM2CLIP-30M 92.0 83.5 69.0 55.3 98.9 98.8 93.1 95.0 89.3 91.2
+LLM2CLIP-60M 94.4 83.2 70.4 55.7 99.2 99.4 94.1 95.2 90.2 92.0
ViT-L/14-336
CLIP 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7
+LLM2CLIP 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0
+LLM2CLIP-60M 93.9 82.3 68.5 54.8 98.9 99.1 94.6 95.9 89.6 90.6
EVA02 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4
+LLM2CLIP 93.9 83.8 68.7 55.7 98.8 99.2 89.5 94.2 89.2 91.3
Table 3: Retrieval Performance across Flickr30K-CN and COCO-CN.
Methods Flickr-CN COCO-CN
I2T@1 I2T@5 I2T@10 T2I@1 T2I@5 T2I@10 I2T@1 I2T@5 I2T@10 T2I@1 T2I@5 T2I@10
ViT-L/14-336
Wukong 76.1 94.8 97.5 51.7 78.9 86.3 53.4 80.2 90.1 55.2 81.0 90.6
CN-CLIP 80.2 96.6 98.2 68.0 90.7 95.4 63.4 84.2 92.9 64.0 89.2 94.4
JinaCLIP 3.30 9.90 15.1 0.7 3.5 6.0 2.9 8.9 13.7 1.0 4.9 8.2
EVA02 4.40 11.8 16.7 0.94 2.9 4.8 2.7 9.8 15.2 1.0 3.7 7.3
+LLM2CLIP 86.9 98.1 99.3 75.1 92.9 96.0 69.1 92.5 97.2 70.0 92.6 96.7
Improving the Discriminability of LLM Output Features is Key to Integrating LLMs with CLIP.
To enhance the discriminability of LLM output features, we applied Caption Contrastive fine-tuning. Llama3-8B-TC and Llama3-8B-CC represent the results of using pure text corpora and a mix of text and augmented CC3M corpora from ShareCaptioner, respectively, both trained with supervised SimCSE loss. As shown in the Table 6, the contrastive learning on mixed caption corpora yields higher CRA scores compared to training on generic text, with noticeable improvements across almost all benchmarks. This difference stems from the fact that LLMs trained on caption-distributed data have stronger discriminability for caption features. These findings highlight the importance of text model discriminability for CLIP training and underscore the rationale behind the caption contrastive fine-tuning in LLM2CLIP.
LLM2CLIP Makes Pretrained SOTA CLIP Even More SOTA.
We applied the LLM2CLIP fine-tuning method to both EVA02 and CLIP models and observed significant performance improvements in Table 2. Even with lightweight fine-tuning, the results substantially surpassed those of the original models pretrained on datasets like Laion2B. Compared to other methods such as LongCLIP and JinaCLIP, which also attempt to fine-tune pretrained CLIP models, our performance gains were transformative. This demonstrates the effectiveness of LLM2CLIP as a method for introducing LLMs to enhance CLIP’s performance.
Table 4: Ablation Study of LLM2CLIP. Here LLM2CLIP-S refers to the results trained on the original short caption dataset.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0
+ Jina-Bert 88.1 77.7 60.5 51.1 83.3 81.0 66.9 68.5 68.9 71.2
       ++ Dense Caption 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9
+ Llama3-8B-S 87.9 75.6 56.7 41.8 55.1 46.1 37.2 35.1 39.3 32.3
       ++ CC Finetuning 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6
        $+$ Dense Caption 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4
LLM’s Built-in Knowledge Directly Benefits CLIP.
In Table 4’s ablation study, we examine the impact of different steps in LLM2CLIP on the EVA02 model. When using only the original captions, there was minimal improvement, as these captions are from the same distribution as those used in EVA02 CLIP’s pretraining. However, we observed that the LLM, through its open-world knowledge, is able to provide a deeper understanding of the original captions, empowering the CLIP model with enhanced capabilities and further boosting performance. Naturally, when dense captions are introduced, the performance improvement becomes even more pronounced.
LLM Improves CLIP’s Ability to Handle Long and Dense Captions.
Thanks to the LLM’s inherent long context windows and strong text understanding capabilities, CLIP can process long caption datasets more effectively without requiring any architectural modifications. This stands in contrast to methods like LongCaption, which require fine-tuning of positional encodings, or Dreamlip, which splits captions into sub-captions. Jina CLIP, for example, enables long-text input by replacing its text encoder with Jina BERT. To verify the LLM’s natural advantage with long texts, we compared our method to Jina CLIP to assess whether the LLM provides a better understanding of long captions. As shown in Table 4, when dense captions are added during training, LLM2CLIP demonstrates significantly greater performance improvements than Jina CLIP, proving that LLMs are better suited to processing long and dense captions.
Larger LLMs Lead to Better Performance.
LLM2CLIP explores the method of replacing CLIP’s text encoder with an LLM and continuing training. In Table X, we thoroughly investigate the performance differences that various versions of LLMs bring to CLIP. In the experiments shown in Table 6, we swapped the text encoder in CLIP with different language models to explore their impact. We maintained the original LLM2CLIP design, with the only change being the initialization of the text encoder. As expected, within the same Llama family, the 8B model significantly outperforms the 1B model. Mistral Namo also shows a slight improvement over the 8B model. It’s important to note that during CLIP training, the LLM’s gradients remain frozen, further proving that larger LLMs, with their richer knowledge, can deliver substantial performance improvements to CLIP. Even so, the 1B LLaMA3.2 model still achieved an impressive performance boost for EVA02 CLIP, with an average improvement of 11.1% — an impressive feat considering that EVA02 is already one of the SOTA open-source CLIP models.
Table 5: Comparison Experiment of Different Ratios of Dense Captions in the LLM2CLIP Training Process.
Ratio Flickr30k COCO ShareGPT4v Urban-1k DOCCI
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
100% 85.5 72.7 60.1 46.9 98.7 99.0 88.7 93.9 88.0 90.5
75% 92.4 82.6 68.5 54.2 98.7 99.3 89.0 94.3 88.1 90.2
50% 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4
25% 93.0 82.8 68.1 54.8 98.4 98.7 87.7 92.9 87.9 90.0
0% 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6
Table 6: Comparison of various text encoders.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average CRA
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
EVA02 Vit-L/14 89.8 73.3 63.8 63.8 89.3 91.9 68.5 73.3 75.0 73.4 76.2 69.8
+Jina Bert 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 78.2 74.2
+Llama3-8B 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 66.3 18.4
+Llama3-8B-TC 92.7 82.1 68.1 54.6 97.7 98.2 88.9 93.8 85.0 87.8 84.8 71.3
+Llama3-8B-CC 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 85.6 73.0
+Llama3.2-1B-CC 91.6 81.3 65.8 52.5 98.3 98.2 84.5 91.9 83.4 86.4 83.4 72.8
+Mistral-Nemo-12B-CC 93.5 83.7 68.5 54.7 98.6 98.9 90.4 94.3 88.0 89.7 86.0 73.3
Table 7: Performance of Llava 1.5. The best results are highlighted in bold.We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
MODEL VQA Datasets Pope Metrics MM Benchmarks Seed Benchmarks
VQAv2 GQA VizWiz SQA-IMG TextVQA Random Adv. Popular MME MMBench MMBench-CN LlavaBench MMVet All IMG Video
Llava (Paper) 78.5 62.0 50.0 66.8 58.2 87.3 86.1 84.2 1510.7 64.3 58.3 65.4 31.1 58.6 66.1 37.3
Llava (Rep.) 79.04 62.86 50.57 67.97 57.48 87.7 84.85 86.3 1476.69 66.66 60.39 58.0 34.3 59.86 66.95 39.71
+LLM2CLIP 79.80 63.15 52.37 69.92 58.35 88.55 82.76 87.75 1505.82 68.29 60.40 62.7 34.8 60.96 68.80 38.96
Like a miracle: LLM enables English CLIP to learn Chinese without exposure to Chinese data.
To validate the knowledge transfer capabilities of the LLM, we designed an out-of-distribution (OOD) experiment by conducting an image-text retrieval task in a completely unfamiliar language. This is an especially challenging experiment because all of our training was performed on purely English text, yet now we are testing the model on Chinese data. As shown in Table 3, models like EVA02 CLIP and Jina CLIP, which performed well on English datasets, achieved near-zero accuracy on this task. However, the magical power of LLM2CLIP became evident: the LLM’s capabilities allowed the model to achieve impressive performance on Chinese retrieval tasks, even surpassing models that were originally trained on hundreds millions of Chinese data, such as Wukong (Gu et al., 2022) and CN-CLIP (Yang et al., 2022). This result once again demonstrates that LLM2CLIP can effectively integrate the inherent abilities of the LLM into CLIP, enabling it to handle tasks far beyond its original scope.
LLM2CLIP Can Improve the Performance of VLLM.
As shown in Table 7, we conducted experiments using the Llava 1.5 Liu et al. (2023) VLLM (Vision-Language Large Model) training framework. Llava incorporates a CLIP visual encoder into the LLM for multimodal instruction learning, meaning the quality of the visual encoder can significantly impact Llava’s performance. We compared the original CLIP with a version enhanced by our LLM2CLIP fine-tuning, running two versions of the experiment according to Llava’s official implementation for a fair comparison. The results showed that in over 87.5% of the benchmarks, we achieved substantial performance improvements, with the remaining benchmarks showing results that were very close. This demonstrates the potential of LLM2CLIP for complex image reasoning and related tasks.
4.3 Ablation
Impact of Different Data Ratios.
With the integration of an LLM, LLM2CLIP enhances our ability to understand dense captions. In  Table 5, we present an ablation study that examines the impact of different ratios of dense captions versus original captions on CLIP’s training performance. The “Ratio” column represents the proportion of dense captions used during training. Overall, dense captions contribute to a noticeable improvement in performance, but more is not always better. For example, when using 0% dense captions, the performance on datasets like COCO and Flickr30K remains decent, but long-text benchmarks such as ShareGPT4V, Urban-1K, and DOCCI show poor results. Conversely, with 100% dense captions, performance on short-text retrieval benchmarks is the worst. It’s important to note that the dense captions we used were generated by the ShareCaptioner model, and there may be some distribution differences and noise compared to real data captions, which could somehow explain why using 100% dense data is suboptimal. Our findings indicate that the best performance is achieved when dense captions make up 50% to 75% of the training data, striking a balance between both short-text and long-text retrieval tasks.
Impact of Different Data Size.
As shown in  Table 2, larger training datasets consistently yield positive results for LLM2CLIP, with clear improvements in both long-text and short-text retrieval tasks. Our 60M dataset version has already pushed the limits of what CLIP could previously achieve. Even the 3M version, though lightweight, still delivers significant performance gains, demonstrating the efficiency of the LLM2CLIP approach. It’s worth noting that training on the 3M dataset takes only about 3 hours on an 8 H100 GPUs machine, yet results in a transformative leap in performance for a well-pretrained CLIP ViT model.
5 Limitations and Future Work
While LLMs possess open-world knowledge, aligning them with CLIP may require specific data to fully unlock their potential. LLM2CLIP is a method that requires only a small amount of data for fine-tuning in conjunction with a pretrained CLIP visual encoder, and we have already demonstrated significant performance gains. However, we did not intentionally select fine-tuning data based on specific characteristics. By focusing on aspects such as data distribution, length, or categories, we could further tailor the LLM to address CLIP’s limitations, allowing the LLM to act as a more comprehensive teacher for various tasks. Moreover, in this work, we froze the gradients of the LLM during fine-tuning to maintain a large batch size for CLIP training. However, it would be worth exploring approaches that allow for updating the LLM’s gradients, as there may be better engineering practices that can balance the two objectives. Re-training LLM2CLIP from scratch on datasets like Laion-2B (Schuhmann et al., 2022) and Recaption-1B (Li et al., 2024b) is also a promising direction that we did not pursue due to time constraints.
6 Conclusion
This paper presents a method that enables LLMs to assist in CLIP training. Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation learning, allowing language — the carrier of human knowledge — to play a pivotal role in spreading the intelligence of LLMs across multiple modalities and applications. The core contributions of this work are threefold: First, we identified the primary reason why LLMs struggle to directly participate in multimodal representation learning — the lack of discriminability in their output features. Second, we introduced the caption contrastive fine-tuning method, which significantly improves feature discriminability, removing the biggest obstacle to utilizing LLMs in CLIP training. Third, we designed the LLM2CLIP training framework, which has proven to be an efficient and effective approach for significantly enhancing the performance of pretrained CLIP models. We experimented with LLM2CLIP on both CLIP and EVA models and validated it across different Llama3 models, including 1B, 8B, and 13B, all of which yielded transformative improvements in CLIP performance. LLM2CLIP offers several advantages, such as seamless compatibility with long and complex text inputs, and the ability to incorporate the open-world knowledge of LLMs into CLIP training — as demonstrated by its ability to match Chinese text even when trained solely on English data. We firmly believe that the potential of LLMs extends far beyond what has been shown, and we hope this work serves as a starting point for using LLMs to enhance various vision foundation models, ultimately spreading their general capabilities and open-world knowledge across a wide range of modalities and applications.

### v3

@@
 Abstract
-CLIP is one of the most important multimodal foundational models today, aligning visual and textual signals into a shared feature space using a simple contrastive learning loss on large-scale image-text pairs.
-What powers CLIP’s capabilities?
-The rich supervision signals provided by natural language — the carrier of human knowledge — shape a powerful cross-modal representation space.
-As a result, CLIP supports a variety of tasks, including zero-shot classification, detection, segmentation, and cross-modal retrieval, significantly influencing the entire multimodal domain.
-However, with the rapid advancements in large language models (LLMs) like GPT-4 and LLaMA, the boundaries of language comprehension and generation are continually being pushed.
-This raises an intriguing question: can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
+CLIP is a foundational multimodal model that aligns image and text features into a shared space using contrastive learning on large-scale image-text pairs.
+Its strength lies in leveraging natural language as a rich supervisory signal.
+With the rapid progress of large language models (LLMs), we explore their potential to further enhance CLIP’s multimodal representation learning.
+This work introduces a fine-tuning approach that integrates LLMs with the pretrained CLIP visual encoder, leveraging LLMs’ advanced text understanding and open-world knowledge to improve CLIP’s ability to process long and complex captions.
+To address the challenge of LLMs’ autoregressive nature, we propose a caption-to-caption contrastive learning framework to enhance the discriminative power of their outputs.
+Our method achieves substantial performance gains on various downstream tasks, demonstrating the effectiveness of combining LLMs with CLIP for enhanced multimodal learning.
+1 Introduction
+CLIP [32] is one of the most influential multimodal foundational models today.
+It aligns visual and language signals into a shared feature space by employing a simple contrastive learning loss on large-scale image-text pairs.
+As a retriever, CLIP supports a wide range of tasks, including zero-shot classification [31], detection [22], segmentation [46], and image-text retrieval [27, 17].
+As a feature extractor, it has become dominant in virtually all cross-modal representation tasks, such as image understanding, video understanding, and text-to-image/video generation.
+For instance, works like LLaVA [24] and Qwen-VL [2] leverage CLIP as a feature extractor to obtain visual features for text models, while models like Stable Diffusion [34] and DALL·E 2 [33] use CLIP’s text encoder to extract textual features for visual models.
+Figure 1: Zero-shot multimodal retrieval radar chart.
+LLM2CLIP significantly improves the performance of original models like CLIP and EVA02 through fine-tuning, substantially outperforming other fine-tuning-based methods such as LongCLIP.
+Figure 2: By adjusting the output space of the LLM, we enable the LLM to more effectively act as a private tutor for CLIP, allowing the fine-tuning process of LLM2CLIP to efficiently adjust the pretrained CLIP’s cross-modal space.
+Figure 3: LLM2CLIP Overview.
+After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training.
+We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision.
+What makes CLIP particularly powerful as a vision encoder?
+The core of its strength lies in its unprecedented ability to align visual pretraining with natural language—the carrier of human knowledge.
+Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text.
+This alignment with language is what sets CLIP apart and unlocks its vast potential.
+However, since CLIP’s introduction, large language models (LLMs) have advanced significantly.
+Models like GPT-4 [1] and Llama [8] now demonstrate remarkable language capabilities, yet these advancements have not translated into corresponding improvements in visual representation learning.
+This raises the question: Can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
 The potential benefits of incorporating LLMs into CLIP are clear.
@@
 This allows them to expand on caption information during training, increasing the efficiency of the learning process.
-However, realizing this potential is challenging.
-Despite LLMs’ powerful internal comprehension, their autoregressive nature hides this capability within the model, leading to output features with poor discriminability.
-Our experiments show that directly integrating LLMs into CLIP results in catastrophic performance drops.
-In this paper, we propose LLM2CLIP, a novel approach that embraces the power of LLMs to unlock CLIP’s potential.
-By fine-tuning the LLM in the caption space with contrastive learning, we extract its textual capabilities into the output embeddings, significantly improving the output layer’s textual discriminability.
-We then design an efficient training process where the fine-tuned LLM acts as a powerful teacher for CLIP’s visual encoder.
-Thanks to the LLM’s presence, we can now incorporate longer and more complex captions without being restricted by vanilla CLIP text encoder’s context window and ability limitations.
-Our experiments demonstrate that this approach brings substantial improvements in cross-modal tasks.
-Our method directly boosted the performance of the previously SOTA EVA02 model by 16.5% on both long-text and short-text retrieval tasks, transforming a CLIP model trained solely on English data into a state-of-the-art cross-lingual model.
-Moreover, when integrated into multimodal training with models like Llava 1.5, it consistently outperformed CLIP across nearly all benchmarks, demonstrating comprehensive performance improvements.
-Figure 1: LLM2CLIP Overview.
-After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training.
-We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision.
-Experimental results demonstrate that LLM2CLIP can make any SOTA CLIP model even more SOTA ever.
-1 Introduction
-CLIP (Radford et al., 2021) is one of the most important multimodal foundational models today.
-It aligns vision and language signals into a shared feature space by employing a simple contrastive learning loss on large-scale image-text pairs.
-As a retriever, CLIP supports a wide range of tasks, including zero-shot classification (Qian & Hu, 2024), detection (Lin & Gong, 2023), segmentation (Zhou et al., 2023), and image-text retrieval (Lülf et al., 2024; Koukounas et al., 2024).
-As a feature extractor, it has become dominant in virtually all cross-modal representation tasks, such as image understanding, video understanding, and text-to-image/video generation.
-For instance, works like LLaVA (Liu et al., 2023) and Qwen-VL (Bai et al., 2023) leverage CLIP as a feature extractor to obtain visual features for text models, while models like Stable Diffusion (Rombach et al., 2021) and DALL·E 2 (Ramesh et al., 2022) use CLIP’s text encoder to extract textual features for visual models.
-What makes CLIP so powerful, particularly as a vision encoder?
-The core of its strength lies in its unprecedented ability to align visual pretraining with natural language — the carrier of human knowledge.
-Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text.
-This alignment with language is what sets CLIP apart and unlocks its vast potential.
-However, since CLIP’s introduction, large language models (LLMs) have advanced significantly.
-Models like GPT-4 (Achiam et al., 2023) and Llama (Dubey et al., 2024) now demonstrate remarkable language capabilities, yet these advancements have not translated to corresponding improvements in visual representation learning.
-This prompts the question: can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
-The potential benefits of incorporating LLMs into CLIP are clear.
-LLMs’ strong textual understanding can fundamentally improve CLIP’s ability to handle image captions, drastically enhancing its ability to process long and complex texts — a well-known limitation of vanilla CLIP.
-Moreover, LLMs are trained on a vast corpus of text, possessing open-world knowledge.
-This allows them to expand on caption information during training, increasing the efficiency of the learning process.
-In this work, we aim to leverage large language models (LLMs) to enable CLIP to learn more powerful, fine-grained, and rich visual representations.
-Currently, CLIP is often criticized for its bag-of-words-like perception and the limitations of its text encoder, which suffers from a constrained model size, limited context length, and is trained predominantly on image captioning data, lacking exposure to diverse world corpora.
-A natural approach would be to replace CLIP’s text encoder with an LLM that embeds rich human knowledge.
-However, this presents significant challenges.
-In the cross-modal contrastive learning framework employed by CLIP, the text encoder functions as a set of knowledge anchors in the shared latent space, guiding the alignment of the vision encoder with human knowledge of the physical world.
-The structure, richness, and discriminability of these knowledge anchors are critical to the visual model’s effectiveness.
-In contrast, LLMs are primarily designed to predict the next word rather than generate explicit representations of the knowledge they contain.
-Their textual comprehension abilities and open-world knowledge are latent within the model, rather than present in the output embeddings, making them difficult to utilize in the same explicit manner as CLIP’s text encoder.
-As a result, using LLMs as a text encoder may not produce linearly separable features, which are crucial for effective feature alignment.
-Table 1: Comparison of top-1 Caption Retrieval Accuracy (CRA) for various language models in MS COCO 5K testing set.
+In this work, we aim to design a method that incorporates the powerful capabilities of LLMs to fine-tune pretrained state-of-the-art (SOTA) CLIP models, further enhancing their performance.
+We propose replacing the weak text encoder of CLIP with an LLM, followed by cross-modal contrastive learning training on image-text datasets.
+While LLMs have already been introduced into CLIP in the InternVL [6] framework, this approach requires large-scale retraining, and the actual performance improvement solely attributable to the LLM remains unclear.
+In our experiments, we found that directly using LLMs to fine-tune CLIP led to failure in all tasks, with a significant drop in performance across various downstream tasks as in Table 6.
+We discovered that the output features of LLMs for different texts (or even the same text) were very similar, which led to poor differentiation.
+This low discriminative power interferes with CLIP’s contrastive learning task.
+To investigate this further, we designed a caption-to-caption retrieval experiment, as shown in Table 1 and Figure 4.
+In this experiment, each image in the MS-COCO dataset has five human-annotated captions.
+We selected the first two captions as positive samples and performed retrieval across the entire test set.
+Using caption retrieval accuracy (CRA), we evaluated the text model’s ability to differentiate between captions, helping us determine which language model is better suited for CLIP.
+We found that Llama-3 8B achieved only 5.2% top-1 accuracy, while the standard CLIP-ViT-L reached 25.2% top-1 accuracy.
+As illustrated in Figure 4, the top-1 caption retrieved by the original Llama-3 model can be entirely unrelated to the query caption, clearly obstructing effective CLIP learning.
+Therefore, directly using an LLM to guide CLIP’s visual encoder training is highly constrained.
+Recent work on LLM2Vec also validates this idea, leading us to realize the necessity of fine-tuning the output space of LLMs.
 Language Model CRA
-CLIP-L/14 25.3
+CLIP-L/14 25.2
 EVA02-L/14 27.1
@@
 Llama3.2-1B 5.6
-Llama3-8B-CC 29.2
-Llama3.2-1B-CC 28.0
-Figure 2: Real examples of top-1 results from the caption-to-caption retrieval experiment. Before fine-tuning, Llama3’s results were often completely unrelated.
-To validate our hypothesis, we designed a caption-to-caption retrieval experiment, as shown in Table 1 and Figure 2.
-Each image in the MS-COCO dataset has five human-annotated captions.
-We selected the first two captions as positive samples and performed retrieval across the entire validation set.
-Using the caption retrieval accuracy (CRA), we evaluated the text model’s ability to differentiate between captions, helping us determine which language model is better suited for CLIP.
-We found that Llama-3 8B achieved only 18.4% top-1 accuracy, while the standard CLIP-ViT-L reached 66.0% top-1 accuracy.
-As illustrated in Figure 2, the top-1 caption retrieved by original Llama-3 can be entirely unrelated to the query caption, clearly obstructing effective CLIP learning.
-Therefore, directly using an LLM to guide CLIP’s visual encoder training is highly constrained.
-We believe that enhancing the discriminative power of LLM output tokens through fine-tuning is vital for the success of our proposed approach, allowing the latent capabilities of LLMs to surface.
-Encouragingly, we found that this can be achieved very efficiently.
-Specifically, we designed a caption contractive (CC) fine-tuning strategy, applying lightweight fine-tuning to the output tokens of Llama-3 8B using LoRA on the CC3M (Sharma et al., 2018) image captioning dataset.
+Llama3-8B-CC 29.5
+Llama3.2-1B-CC 29.4
+Table 1: Comparison of top-1 Caption Retrieval Accuracy (CRA) for various language models on the MS COCO 5K test set.
+Figure 4: Real examples of top-1 results from the caption-to-caption retrieval experiment in MS COCO 5K test set. Before fine-tuning, Llama3’s results were often completely unrelated.
+We designed a caption contractive (CC) fine-tuning strategy, as Figure 2, pplying lightweight fine-tuning to the output tokens of Llama-3 8B using LoRA on the CC3M [36] image captioning dataset.
 The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions.
-We utilized a supervised SimCSE (Gao et al., 2021; BehnamGhader et al., 2024) contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner (Chen et al., 2023) were treated as positive pairs, pulling them closer.
+We utilized a supervised SimCSE [11, 3] contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner [5] were treated as positive pairs, pulling them closer.
 In contrast, all other captions formed a negative sample set that the model learned to push away.
-Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 18.4% to 73%, which is a 7% improvement over the original CLIP-ViT-L text encoder.
-This successful fine-tuning process enables us to more effectively harness the open-world capabilities of LLMs for CLIP training.
-In a nutshell, we present LLM2CLIP, a novel approach for enhancing visual representation learning through the integration of large language models (LLMs) as shown in Figure 1.
-This method takes a straightforward yet audacious step by replacing the original CLIP text encoder and augmenting the CLIP visual encoder with the vast knowledge embedded in LLMs.
-We have identified key obstacles associated with this innovative idea and proposed a cost-effective fine-tuning strategy to overcome them.
-Our experiments demonstrate that leveraging LLMs as teachers for CLIP training yields substantial improvements, with LLM2CLIP significantly outperforming state-of-the-art pre-trained CLIP models.
-Our method increased the performance of the previously SOTA EVA02 model by 16.5% on both long-text and short-text retrieval tasks, transforming a CLIP model trained solely on English data into a state-of-the-art cross-lingual model.
-Furthermore, when incorporated into multimodal model training, such as with Llava 1.5, it consistently achieved comprehensive improvements over EVA02 across nearly all benchmarks.
-Additionally, the efficient training method proposed by LLM2CLIP ensures that the training cost is nearly identical to fine-tuning the original CLIP.
-We have also demonstrated that using more powerful language models and larger training datasets can further boost LLM2CLIP’s performance, showcasing the immense potential of our approach.
-These promising outcomes affirm that we have successfully transformed CLIP into a more general-purpose foundational model.
-The enhanced LLM2CLIP model possesses richer knowledge and exhibits a remarkable capacity for distinguishing fine-grained and complex long-text semantics.
-This advancement not only broadens the range of supported downstream tasks but also propels progress across the entire vision domain.
+Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 5.2% to 29.5%.
+This successful fine-tuning process enables us to more effectively harness the capabilities of LLMs for CLIP training.
+Building on fine-tuning of LLMs, we designed an efficient approach, LLM2CLIP, to enhance the performance of pretrained CLIP.
+As in  Figure 3, This method uses an LLM to further fine-tune the pretrained CLIP, allowing the LLM to impart knowledge and capabilities to CLIP’s vision encoder that were lacking during its initial pretraining.
+LLM2CLIP consists of two stages: in the first stage, we apply caption contrastive (CC) fine-tuning on the LLM with a small dataset to adjust the LLM’s output space.
+In the second stage, we fine-tune the CLIP vision encoder with the CC-fine-tuned LLM, with LLM gradients frozen, to further improve CLIP’s performance.
+Experimental results demonstrate that training on only 15M open-source data substantially boosts CLIP’s capabilities, with even 3M data yielding impressive results.
+The main contributions of this paper are:
+  1. 1.
+    We identify the core limitation preventing LLMs from fully leveraging their capabilities within the CLIP training framework—the insufficient separability of LLM output features.
+  2. 2.
+    We introduce a straightforward approach by applying caption-to-caption contrastive fine-tuning on the LLM’s output layer, enhancing the separability of the output space.
+This approach shows experimentally that involving LLMs in the CLIP training process can provide substantial benefits.
+  3. 3.
+    We propose LLM2CLIP, an efficient and effective LLM-assisted fine-tuning method for CLIP, enabling low-cost fine-tuning of pretrained CLIP models and achieving significant performance improvements.
 2 Related Works
-Contrastive Language-Image Pre-training (CLIP).
-CLIP (Radford et al., 2021) is one of the most important multimodal models today.
-By training on web-scale data, image captions serve as rich, fine-grained supervision signals for learning representations that align closely with image, freeing CLIP from the limitations of manually defined labels.
-The richness of the captions endows CLIP with many remarkable capabilities; besides zero-shot image-text retrieval, CLIP features can even achieve region-level correspondence, supporting zero-shot detection and segmentation tasks.
-Furthermore, CLIP is the most widely used encoder in current image and video understanding tasks.
-In Tong et al.
-(2024)’s analysis, it stands out among various visual encoders in MLLMs training as the most significant model, demonstrating its strong capabilities.
-For generative tasks, CLIP is equally foundational.
-Its text and image encoders are used to provide text control (Rombach et al., 2021; Ramesh et al., 2022) and image control signals (Blattmann et al., 2023; Zhang et al., 2023) for conditioning generative models.
-Additionally, the CLIP score (Hessel et al., 2021) has become a crucial metric for assessing the relevance of generated images to text.
-While the original CLIP only utilized image and text modalities, there have been massive extensions of CLIP to incorporate various other modalities (Dai et al., 2024; He et al., 2023; Moon et al., 2022; Zhang et al., 2022).
-CLIP definitely has a significant impact on the field of vision and multimodal, and its influence is still growing.
 CLIP meets Stronger Language Models.
 Several works have explored the integration of LLMs into CLIP.
-JinaCLIP (Koukounas et al., 2024) employed Jina-embeddings-v2 (Günther et al., 2023) as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts.
-Though achieving similar visual performance to EVA-CLIP (Sun et al., 2023), its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs.
-T5-V (Jiang et al., 2024) took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B (Li et al., 2024a) while freezing the ViT gradient to focus on adjusting the MLLM output.
+JinaCLIP [17] employed Jina-embeddings-v2 [13] as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts.
+Though achieving similar visual performance to EVA-CLIP [38], its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs.
+E5-V [15] took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B [19] while freezing the ViT gradient to focus on adjusting the MLLM output.
 However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP.
-MATE (Jang et al., 2024) designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs.
+MATE [14] designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs.
 They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks.
@@
 Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly.
-For instance, DCI (Urbanek et al., 2024) hired annotators to expand captions based on SAM (Kirillov et al., 2023)’s target hints.
-LaCLIP (Fan et al., 2024) adopted ChatGPT, Bard, and human rewriting.
-DreamLIP (Zheng et al., 2024) leveraged ShareCaptioner (Chen et al., 2023) and InstructBLIP to augment 30M captions.
-Recap-DataComp-1B (Li et al., 2024b) used Llama3-trained LLAVA1.5 to extend 1B captions.
-In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones (Urbanek et al., 2024), splitting captions into multiple segments (Fan et al., 2024; Zheng et al., 2024), or finetuning model’s positional encoding to support longer token inputs (Zhang et al., 2024).
+For instance, DCI [40] hired annotators to expand captions based on SAM [16]’s target hints.
+LaCLIP [9] adopted ChatGPT, Bard, and human rewriting.
+DreamLIP [45] leveraged ShareCaptioner [5] and InstructBLIP to augment 30M captions.
+Recap-DataComp-1B [20] used Llama3-trained LLAVA1.5 to extend 1B captions.
+In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones [40], splitting captions into multiple segments [9, 45], or finetuning model’s positional encoding to support longer token inputs [44].
 LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions.
@@
 3 Methods
-The contributions of our methodology are threefold: First, we designed experiments to analyze the key reason preventing LLMs from directly participating in multimodal representation learning — the weak discriminability of their output features.
-Second, we introduced the caption contrastive fine-tuning method, which significantly improves feature discriminability.
-Third, we developed the LLM2CLIP training framework, which has been proven to be an efficient and effective method for leveraging LLMs to deliver substantial performance improvements to pretrained CLIP models.
-3.1 Native LLMs are Ineffective Text Encoders for CLIP
-Large Language Models (LLMs), such as Llama-3, are trained on massive world corpora through auto-regression, equipping them with open-world knowledge and enabling them to perform various tasks.
-We initially hoped to leverage the capabilities of LLMs to directly retrain a CLIP model.
-Although LLMs exhibit strong text comprehension, they are difficult to use directly as text embedding models.
-This is because their knowledge is encapsulated within the model, and their output features are heavily skewed towards individual word predictions.
-As generative models, they are not trained to ensure good linear separability of output features, making them less effective when used to interpret captions for CLIP.
-As highlighted by  Chen et al.
-(2022), cross-modal contrastive learning in CLIP requires that each modality possesses strong internal discriminability.
-To evaluate the effectiveness of various language models in terms of text discriminability, and to test whether the native output features of LLMs, as hypothesized, struggle to distinguish image captions, we introduce a new metric: the MS COCO Caption Retrieve Accracy (CRA).
-MS COCO is a widely used multimodal dataset, containing over 330K images, each with five captions.
-These captions are written by different human annotators and provide diverse descriptions for each image.
-In our evaluation on the MS COCO validation set, we use only the first two captions of each image and treat the captions of the same image as positive pairs, while all other captions serve as negative samples.
-We then perform caption-to-caption retrieval and assess Top-1 accuracy using different language models, defining the result as their CRA score.
-Higher CRA scores indicate better discriminability of the language models on image captions.
-As shown in Table 1, using a pure LLM results in a CRA score of only 18.4%, indicating that the majority of captions cannot be well-separated in the output space.
-In fact, captions with similar distances may be entirely unrelated, as illustrated in Figure 2.
-However, the text encoder from the original state-of-the-art CLIP model achieves a CRA score of 66%, proving the inadequacy of native LLM output features in caption discriminability.
-Consequently, it is challenging to apply LLMs directly in CLIP model training.
-Subsequent experiments in Table 6 also confirm that replacing CLIP’s text encoder and the corresponding ViT with Llama-3 8B for contrastive learning significantly underperforms the original CLIP.
-3.2 A Crucial Lesson: LLM Learning for Image Caption Discrimination
+LLM2CLIP is a fine-tuning strategy for CLIP, consisting of two stages.
+The first stage aims to improve the separability of the LLM’s output space, and the second stage involves fine-tuning by replacing CLIP’s text model with the LLM.
+3.1 Caption Contrastive Fine-tuning
 In this section, we aim to fine-tune the LLM’s token outputs to better capture features that can distinguish between image captions.
@@
 Therefore, we apply a caption contrastive (CC) fine-tuning to the LLM’s output features, treating different captions of the same image as positive samples and the rest of the captions as negative samples.
-To obtain enough varied descriptions for the same image, we use the ShareCaptioner (Zheng et al., 2024; Chen et al., 2023) modified CC-3M (Sharma et al., 2018) dataset, which provides both original captions and augmented dense captions for each image.
+To obtain enough varied descriptions for the same image, we use the ShareCaptioner [45, 5] modified CC-3M [36] dataset, which provides both original captions and augmented dense captions for each image.
 These can be treated as positive pairs.
-We followed the training methodology of LLM2Vec (BehnamGhader et al., 2024), first expanding the LLM’s attention mechanism to bidirectional attention, and employing Masked Next Token Prediction (MNTP) for initialization to achieve better results.
+We followed the training methodology of LLM2Vec [3], first expanding the LLM’s attention mechanism to bidirectional attention, and employing Masked Next Token Prediction (MNTP) for initialization to achieve better results.
 Specifically:
-First, we transform the LLM’s causal attention mechanism into bidirectional attention, as we no longer need it to retain generative capabilities but rather function as an encoder.
+  1. 1.
+    Bidirectional Attention Transformation: First, we transform the LLM’s causal attention mechanism into bidirectional attention, as we no longer need it to retain generative capabilities but rather function as an encoder.
 Since autoregressive training is no longer required, switching to bidirectional attention improves its ability to capture contextual information.
-Second, we employ MNTP to train the newly added bidirectional attention mechanism, providing a strong initialization.
-For a given sequence of $N$ tokens, we mask a subset and predict their values, similar to BERT (Devlin et al., 2018).
+  2. 2.
+    Masked Next Token Prediction (MNTP): Second, we employ MNTP to train the newly added bidirectional attention mechanism, providing a strong initialization.
+For a given sequence of $N$ tokens, we mask a subset and predict their values, similar to BERT [7].
 However, unlike BERT, we adapt this process to fit the nature of LLMs by predicting tokens just before the masked token.
 We train on both image captions and pure text with equal weighting.
-In addition to CC-3M, we also use the Wikitext-103 (Merity et al., 2016) dataset to preserve the LLM’s text capabilities and prevent divergence from its original strengths.
-Finally, we perform the actual caption contrastive fine-tuning, using a supervised SimCSE loss to bring captions of the same image closer together and push apart captions of different images.
+In addition to CC-3M, we also use the Wikitext-103 [28] dataset to preserve the LLM’s text capabilities and prevent divergence from its original strengths.
+  3. 3.
+    Caption Contrastive Learning: Finally, we perform the actual caption contrastive fine-tuning, using a supervised SimCSE loss to bring captions of the same image closer together and push apart captions of different images.
 We use two prompt templates: “Given a caption, retrieve a detailed relevant caption” and “Given a detailed caption, retrieve a short relevant caption,” which are prepended to the query (either original or dense caption) to retrieve the corresponding dense or original caption.
-Similarly, we use a 1.5M-paired general text dataset curated from  Springer et al.
-(2024) to maintain strong performance in pure language tasks.
+Similarly, we use a 1.5M-paired general text dataset curated from [37] to maintain strong performance in pure language tasks.
 All training is efficiently conducted using LoRA and is completed in just one epoch, ensuring low computational cost.
-A remarkable result followed: after fine-tuning the LLM as a caption encoder for only one epoch on CC-3M, the CRA score for Llama-3 8B jumped from 18.4% to 73.0%, surpassing the previous state-of-the-art CLIP and EVA models’ text encoders trained on the massive Laion-2B (Schuhmann et al., 2022) and COYO-700M (Byeon et al., 2022) image-text datasets.
-As shown in Table 6, subsequent experiments demonstrated that after CC fine-tuning, the LLM finally unleashed its powerful capabilities, significantly boosting the performance of the previously state-of-the-art CLIP model, in stark contrast to the results without CC fine-tuning.
-This breakthrough uncovers the potential of LLMs in CLIP training and removes a major obstacle in leveraging LLMs to advance vision foundation models.
-3.3 LLM2CLIP: Makes SOTA CLIP more SOTA ever
+In our experiments, we found that the most critical step was the third step with the SimCSE loss, which significantly enhanced the caption discriminability of the LLM.
+3.2 LLM2CLIP: Use LLM to Finetuning
 With the modifications to the LLM discussed earlier, we have now obtained a super text encoder that is well-suited for CLIP training.
@@
 Second, by freezing the LLM, we ensure that the open-world knowledge it has acquired from large-scale corpora remains intact during the multimodal alignment process.
-To compensate for the frozen LLM, and inspired by methods like FuseMix (Vouitsis et al., 2023) and APE (Rosenfeld et al., 2022), we introduce several new linear layers after the LLM as adapters.
+To compensate for the frozen LLM, and inspired by methods like FuseMix [41] and APE [35], we introduce several new linear layers after the LLM as adapters.
 These layers serve as learnable parameters to improve the alignment between the LLM and CLIP visual encoder.
 Following the original design of CLIP, we also employ a projector layer to align the dimensions of the two encoders, facilitating the use of CLIP loss for training.
-With this powerful LLM-based super text encoder, we achieve a substantial qualitative leap in CLIP’s language comprehension capabilities.
+With this powerful LLM-based super text encoder, we achieve a substantial qualitative leap in CLIP’s language comprehension capabilities as results in Table 2.
 The LLM’s open-world knowledge allows the CLIP visual encoder to learn more structured and globally informed visual representations that are aligned with human knowledge.
 Moreover, this approach enables us to fully utilize high-quality, long, and dense caption datasets without requiring any special architectural adjustments, which previous works like DCI, DreamLip, and Recaption struggled to effectively leverage.
-As shown in  Table 2, LLM2CLIP makes any existing SOTA CLIP model even more SOTA, significantly surpassing the performance of previous one.
 3.3 Overview and Efficiency Discussion
@@
 The additional trainable parameters are roughly equivalent to those in the original CLIP, minimizing the overhead.
-To further reduce the inference cost of using LLMs, we pre-extract all the text features from the training data and store them in memory.
+To further reduce the inference cost of using LLMs, we pre-extract all the text features from the training data and store them in disk memory.
 This ensures that, even though LLMs provide powerful textual supervision, the memory and computational costs during training remain nearly identical to those of standard CLIP training.
@@
 Despite this efficient training cost, LLM2CLIP brought transformative improvements in downstream tasks such as long and short text retrieval, cross-language retrieval, and Llava training.
+Methods Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
+I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
+ViT-B/16
+ALIGN 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
+BLIP 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
+Jina-CLIP 80.6 67.4 55.6 41.1 - - 87.7 88.0 78.7 80.0 75.7 69.1
+Long-CLIP 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
+CLIP 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
++LLM2CLIP 89.2 78.1 62.2 48.7 98.1 97.4 86.1 90.0 84.1 85.0 83.9 79.8
+EVA02 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
++LLM2CLIP 88.5 78.0 63.6 49.8 98.0 98.1 84.7 89.7 85.5 86.8 84.1 80.5
+ViT-L/14
+Long-CLIP 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
+CLIP 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
++LLM2CLIP 92.6 81.7 64.9 52.5 98.4 98.4 87.6 92.0 87.6 88.7 86.2 82.7
+EVA02 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
++LLM2CLIP-3M 89.6 77.3 59.7 48.0 98.3 98.6 87.1 91.1 84.9 87.8 83.9 80.6
++LLM2CLIP 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
++LLM2CLIP-30M 92.0 83.5 69.0 55.3 98.9 98.8 93.1 95.0 89.3 91.2 88.5 84.8
++LLM2CLIP-60M 94.4 83.2 70.4 55.7 99.2 99.4 94.1 95.2 90.2 92.0 89.7 85.1
+ViT-L/14-336
+CLIP 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
++LLM2CLIP 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0 86.6 83.3
++LLM2CLIP-60M 93.9 82.3 68.5 54.8 98.9 99.1 94.6 95.9 89.6 90.6 89.1 84.5
+EVA02 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
++LLM2CLIP 93.9 83.8 68.7 55.7 98.8 99.2 89.5 94.2 89.2 91.3 88.0 84.8
+Table 2: Systematic Comparison Experiment Demonstrating the Performance Improvements of LLM2CLIP. Results with a blue background represent the outcomes of LLM2CLIP.
 4 EXPERIMENTS
@@
 Training Dataset.
-We utilized the ShareCaptioner-modified CC-3M dataset (Zheng et al., 2024; Chen et al., 2023), which provides both original captions and augmented dense captions for each image, for contrastive learning.
-For the Masked Next Token Prediction and caption contrastive fine-tuning stages, we employed the Wikitext-103 dataset (Merity et al., 2016) and the E5 dataset from  Springer et al.
-(2024) to ensure that the general text domain was not excessively biased.
+For contrastive learning, we utilized the ShareCaptioner-modified CC-3M dataset [45, 5], which includes both original captions and augmented dense captions for each image.
+In addition, we used the Wikitext-103 dataset [28] and the E5 dataset [37] during the Masked Next Token Prediction and caption contrastive fine-tuning stages to ensure coverage of the general text domain without excessive bias.
 Training Setting.
-We trained the language model using LoRA, applying lightweight training with 1 epoch on all datasets.
-We adopted the average of all LLM output tokens as the text global embedding for a caption.
-All training parameters follow the design of  BehnamGhader et al.
-(2024).
+The language model was trained with LoRA for lightweight training, completing one epoch across all datasets.
+We used the average of all LLM output tokens as the global text embedding for a caption.
+All training parameters followed the configuration detailed in [3].
+The design of the Adaptor is the same as FuseMix [41], using a simple inverted bottleneck MLP architecture.
 4.1.2 LLM2CLIP Fine-tuning
 Training Dataset.
-To compare different configurations, we designed three experimental settings based on dataset size: LLM2CLIP-3M: This configuration only uses the CC-3M dataset for training, representing our lightweight setting.
-LLM2CLIP-15M: This configuration uses both CC-3M and CC-12M datasets, which is our default setting.
-All LLM2CLIP models without a specific dataset size label are trained with this 15M data.
-LLM2CLIP-60M: This configuration scales up the dataset, combining CC-3M, CC-12M, YFCC-15M, and a randomly selected 30M subset from Recaption-1B (Li et al., 2024b).
-All data used are dense captions rewritten by multimodal large language models (MLLMs).
-The CC-3M, CC-12M, and YFCC datasets are sourced from  Zheng et al.
-(2024) using the ShareCaptioner for rewriting, while Recaption data was rewritten using Llama3-LLAVA1.5.
-We primarily used a mix of original captions and dense captions, with the default mixing ratio being 1:1.
-LLM2CLIP-S represents a setting where only original data is used for training, matching the original CLIP pretraining distribution to analyze the model’s benefits separately.
+We defined three experimental configurations based on dataset size:
+  • LLM2CLIP-3M: This configuration uses only the CC-3M dataset, serving as our lightweight setting.
+  • LLM2CLIP-15M: Our default configuration, combining CC-3M and CC-12M datasets.
+  • LLM2CLIP-60M: A larger configuration that combines CC-3M, CC-12M, YFCC-15M, and a 30M randomly selected subset from Recaption-1B [20].
+All datasets included dense captions rewritten by multimodal large language models (MLLMs).
+Specifically, the CC-3M, CC-12M, and YFCC datasets were sourced from [45] and rewritten using ShareCaptioner, while Recaption data were rewritten with Llama3-LLAVA1.5.
+A 1:1 mix of original captions and dense captions was used by default.
 Training Setting.
-All experiments with different datasets were conducted for 4 epochs.
-We froze the LLM gradients by default.
-During training, we pre-extracted the text features from the captions using the LLM and stored them in memory to avoid repeated LLM inference, reducing additional computational overhead.
-We trained the original CLIP vision encoder, the LLM’s learnable adapter, and the projectors for both encoders.
-Training on the 3M, 15M, and 60M datasets for CLIP ViT-L/14-224 required only 30G per GPU memory and 2, 9, and 45 hours on 8 H100 GPUs, respectively.
-4.1.3 Evaluation
-Evaluation Dataset.
-For short-text datasets, we used the commonly available COCO 2014 5k test set and the Flickr 1k test set as our test datasets.
-For long-text datasets, we used the 1K subset of the ShareGPT4V (Chen et al., 2023) dataset and the Urban1k (Zhang et al., 2024) dataset, both provided by LongCLIP, along with the DOCCI (Onoe et al., 2024) dataset.
-The ShareGPT4V-1M dataset was generated using annotations from GPT-4V and ShareCaptioner, with images from Laion, CC, SBU (Ordonez et al., 2011) and MS COCO.
-We used a randomly selected 1K subset of this dataset.
-Urban1k consists of captions generated by GPT-4V for 1,000 busy urban view images from Visual Genome.
-Each caption is a long and complete sentence that describes the image, including types, colors, and relative locations of various attributes.
-The model can only successfully match the images with the correct captions if it accurately understands and models the detailed attributes in both modalities.
-The DOCCI dataset contains approximately 1.5K high-resolution images with detailed human-annotated descriptive captions.
-DOCCI is divided into a training set of 9.6K pairs and a test set of 5.1K pairs.
-We used the test set for image-lengthy caption retrieval experiments.
-For the Chinese retrieval tasks, we used the FlickrCN (Lan et al., 2017) and CNCOCO (Li et al., 2018) datasets, which are translated versions of Flickr30K and MS COCO-1K, respectively.
-These datasets were tested in Chinese.
-For Llava (Liu et al., 2023) training, we used the standard training and test sets from LLAVA 1.5.
-Evaluation Setting.
-The experiments in this paper primarily focus on how LLM2CLIP enhances the performance of widely used vanilla CLIP across various dimensions.
-We mainly compare our approach with the EVA (Fang et al., 2023) and OpenAI CLIP models as baselines, as they are the most widely used SOTA vision encoders in the open-source community.
-The CLIP models we use for comparison are ViT-B/16222https://huggingface.co/openai/clip-vit-base-patch16, ViT-L/14333https://huggingface.co/openai/clip-vit-large-patch14, and ViT-L/14-336444https://huggingface.co/openai/clip-vit-large-patch14-336.
-For EVA02 models, we use EVA02 ViT-B/16555https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02_CLIP_B_psz16_s8B.pt, EVA02 ViT-L/14666https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02_CLIP_L_psz14_s4B.pt, and EVA02 ViT-L/14-336777https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02_CLIP_L_336_psz14_s6B.pt, making up a total of six models.
-For language models, we experimented with four different models: Jina-Embeddings-V2888https://huggingface.co/jinaai/jina-embeddings-v2-base-en, and three popular LLMs from the Llama3 (Dubey et al., 2024) family, namely Llama 3.2 1B, Llama 3 8B, and Mistral-Nemo 12B.
-To avoid any misunderstanding, the experiments in this paper, unless otherwise noted, use EVA02 ViT-L/14 as the default baseline for comparison.
-The default language model used is LLaMA 3 8B, trained on a dataset of 15M setting mentioned above.
-4.2 Main Results
-Directly Replacing CLIP’s Text Encoder with a Vanilla LLM is Harmful.
-As shown in the experiments in Table 4, directly replacing the text encoder of EVA02 ViT-L/14 with Llama3-8B significantly degrades retrieval performance.
-For example, performance on the DOCCI benchmark nearly halves, dropping from 75.0/73.4 to 51.7/50.6.
-Additionally, the CRA score of the vanilla Llama3-8B is particularly low, indicating that its output features exhibit very poor discriminability for captions.
-These results clearly show that such an approach imposes a significant burden on CLIP’s learning process.
-Table 2: Systematic Comparison Experiment Demonstrating the Performance Improvements of LLM2CLIP.
-Methods Flickr30k COCO ShareGPT4V Urban-1k DOCCI
-I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-ViT-B/16
-ALIGN 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1
-BLIP 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5
-Jina-CLIP 80.6 67.4 55.6 41.1 - - 87.7 88.0 78.7 80.0
-Long-CLIP 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4
-CLIP 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1
-+LLM2CLIP 89.2 78.1 62.2 48.7 98.1 97.4 86.1 90.0 84.1 85.0
-EVA02 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0
-+LLM2CLIP 88.5 78.0 63.6 49.8 98.0 98.1 84.7 89.7 85.5 86.8
-ViT-L/14
-Long-CLIP 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6
-CLIP 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8
-+LLM2CLIP 92.6 81.7 64.9 52.5 98.4 98.4 87.6 92.0 87.6 88.7
-EVA02 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0
-+LLM2CLIP-3M 89.6 77.3 59.7 48.0 98.3 98.6 87.1 91.1 84.9 87.8
-+LLM2CLIP 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4
-+LLM2CLIP-30M 92.0 83.5 69.0 55.3 98.9 98.8 93.1 95.0 89.3 91.2
-+LLM2CLIP-60M 94.4 83.2 70.4 55.7 99.2 99.4 94.1 95.2 90.2 92.0
-ViT-L/14-336
-CLIP 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7
-+LLM2CLIP 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0
-+LLM2CLIP-60M 93.9 82.3 68.5 54.8 98.9 99.1 94.6 95.9 89.6 90.6
-EVA02 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4
-+LLM2CLIP 93.9 83.8 68.7 55.7 98.8 99.2 89.5 94.2 89.2 91.3
-Table 3: Retrieval Performance across Flickr30K-CN and COCO-CN.
+Each configuration was trained for four epochs.
+We froze the LLM gradients to reduce computational overhead and pre-extracted text features using the LLM, storing them on disk to avoid repeated inference during training.
+Inspired by Fusemix’s efficient training strategy, we added a four-layer linear adapter after the LLM and trained with gradients enabled for the CLIP ViT model to better integrate LLM knowledge.
+Additionally, we introduced two projector layers after each encoder for dimensional alignment.
+Training the CLIP ViT-L/14-224 model required approximately 30GB of GPU memory per GPU and took 2, 9, and 45 hours on 8 H100 GPUs for the 3M, 15M, and 60M datasets, respectively.
+We employed the AdamW [26] optimizer, utilizing a total batch size of 4096, and set the learning rate to 1e-5 with a cosine decay rate of 0.05.
+For language models, we tested four configurations: Jina-Embeddings-V2, and three popular models from the Llama3 family [8]: Llama 3.2 1B, Llama 3 8B, and Mistral-Nemo 12B.
+Unless otherwise noted, experiments used EVA02 ViT-L/14 as the baseline and Llama 3 8B as the default language model.
 Methods Flickr-CN COCO-CN
@@
 +LLM2CLIP 86.9 98.1 99.3 75.1 92.9 96.0 69.1 92.5 97.2 70.0 92.6 96.7
+Table 3: Retrieval Performance across Flickr30K-CN and COCO-CN.
+Methods Imagenet
+0-shot Linear
+CLIP L/14-336 76.6 84.8
++ LLM2CLIP 74.1 85.2
+CLIP L/14 75.5 83.9
++ LLM2CLIP 71.8 84.4
+CLIP B16 70.2 79.9
++ LLM2CLIP 67.3 80.6
+Table 4: Imagenet 0-shot classification and linear probe results.
+MODEL VQA Datasets Pope Metrics MM Benchmarks Seed Benchmarks
+VQAv2 GQA VizWiz SQA-IMG TextVQA Random Adv. Popular MME MMBench MMBench-CN LlavaBench MMVet All IMG Video
+Llava (Paper) 78.5 62.0 50.0 66.8 58.2 87.3 86.1 84.2 1510.7 64.3 58.3 65.4 31.1 58.6 66.1 37.3
+Llava (Rep.) 79.04 62.86 50.57 67.97 57.48 87.7 84.85 86.3 1476.69 66.66 60.39 58.0 34.3 59.86 66.95 39.71
++LLM2CLIP 79.80 63.15 52.37 69.92 58.35 88.55 82.76 87.75 1505.82 68.29 60.40 62.7 34.8 60.96 68.80 38.96
+Table 5: Performance of Llava 1.5 benchmarks. The best results are highlighted in bold.We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
+Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average CRA
+I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
+EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5 27.1
++Jina Bert 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0 29.68
++Llama3-8B 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0 5.22
++Llama3-8B-TC(LLM2Vec) 92.7 82.1 68.1 54.6 97.7 98.2 88.9 93.8 85.0 87.8 86.5 83.3 28.42
++Llama3-8B-CC 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2 29.46
++Llama3.2-1B-CC 91.6 81.3 65.8 52.5 98.3 98.2 84.5 91.9 83.4 86.4 84.7 82.1 29.42
++Mistral-Nemo-12B-CC 93.5 83.7 68.5 54.7 98.6 98.9 90.4 94.3 88.0 89.7 87.8 84.3 29.96
+Table 6: Comparison of various text encoders.
+Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
+I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
+EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
++ Jina-Bert 88.1 77.7 60.5 51.1 83.3 81.0 66.9 68.5 68.9 71.2 73.5 70.0
+       ++ Dense Caption 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0
++ Llama3-8B 87.9 75.6 56.7 41.8 55.1 46.1 37.2 35.1 39.3 32.3 55.2 46.2
+       ++ Dense Caption 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0
++ Llama3-8B-CC 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
+       ++ Dense Caption 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
+Table 7: Ablation analysis of performance variations between short simple captions and long & dense captions using different text encoders.
+We conducted experiments on the EVA02 ViT-L/14 model, replacing its original text encoder with Jina-BERT (+ Jina-BERT), the original Llama3 (+ Llama3-8B), and Llama3 fine-tuned with Caption Contrastive learning (+ Llama3-8B-CC).
+Fine-tuning was performed on datasets containing only short simple captions and datasets with a 50% probability mix of long dense captions.
+The objective of this experiment was to isolate the effects of LLMs and long dense captions, and to demonstrate that the design of LLM2CLIP enhances the efficiency of dense caption utilization.
+The results with a red background indicate that the LLM without fine-tuning has a significant negative effect on CLIP.
+Ratio Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
+I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
+100% 85.5 72.7 60.1 46.9 98.7 99.0 88.7 93.9 88.0 90.5 84.2 80.6
+75% 92.4 82.6 68.5 54.2 98.7 99.3 89.0 94.3 88.1 90.2 87.3 84.1
+50% 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
+25% 93.0 82.8 68.1 54.8 98.4 98.7 87.7 92.9 87.9 90.0 87.0 83.8
+0% 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
+Table 8: Comparison Experiment of Different Ratios of Dense Captions in the LLM2CLIP Training Process.
+Adaptor Layers Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
+I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
+0 91.3 79.9 64.0 50.2 97.1 97.7 88.1 90.2 81.0 86.1 84.3 80.8
+1 92.1 80.6 65.9 51.8 97.8 98.0 90.9 91.8 84.3 87.9 86.2 82.0
+4 92.1 79.9 66.0 52.2 97.4 98.0 93.0 92.1 85.0 88.0 86.7 82.0
+Table 9: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different adaptor layer numbers.
+Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks.
+The best result in each column is highlighted in bold.
+All experiments were conducted using OpenAI’s CLIP model with a ViT architecture at 224 resolution, trained on 15M data points.
+Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
+I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
+Bidirectional Attention + MNTP + Sup. SimCLR 83.3 73.5 55.3 45.6 98.1 98.0 81.5 87.9 78.8 82.6 79.4 77.5
+Bidirectional Attention + Sup. SimCLR 84.6 74.5 55.6 46.2 98.0 97.6 79.3 88.5 80.0 82.6 79.5 77.9
+Sup. SimCLR 82.0 73.5 54.6 44.8 97.6 97.6 82.2 87.9 76.9 81.5 78.7 77.1
+Table 10: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different methods.
+Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks.
+The best result in each column is highlighted in bold.
+The experiments were conducted using the Llama 3.2 1B and the EVA02 ViT-L/14 model on 3M data.
+It can be observed that MNTP has no significant effect, while bidirectional attention provides a slight performance improvement.
+4.2 Main Results
+Evaluation Setting.
+Our experiments assessed LLM2CLIP’s ability to enhance vanilla CLIP performance across multiple dimensions.
+We compared our approach against state-of-the-art (SOTA) models, including EVA [10] and OpenAI CLIP.
+Specifically, we evaluated CLIP ViT-B/16, ViT-L/14, and ViT-L/14-336, along with EVA02 ViT-B/16, EVA02 ViT-L/14, and EVA02 ViT-L/14-336, yielding a total of six comparison models.
+4.2.1 Short- and Long-Text Retrieval
+Evaluation Dataset.
+For short-text retrieval, we used the MSCOCO [23] 5K test set and the Flickr [43] 1K test set.
+For long-text retrieval, we employed datasets from LongCLIP, including a 1K subset of ShareGPT4V [5], the Urban1K [44] dataset, and the DOCCI [29] dataset.
+The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU [30], and MS COCO.
+Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions.
+DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation.
+LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval.
+As shown in Table 2, LLM2CLIP fine-tuning yielded significant performance improvements, even surpassing models pretrained on large datasets like LAION2B.
+Compared to LongCLIP and JinaCLIP, our method achieved transformative gains, showcasing its efficiency in leveraging LLM knowledge to enhance pretrained CLIP models.
+4.2.2 Cross-Lingual Retrieval
+Evaluation Dataset.
+For Chinese retrieval tasks, we used FlickrCN [18] and CNCOCO [21], which are translated versions of Flickr30K and MS COCO-1K, respectively.
+LLAVA training and evaluation were conducted using the standard splits from LLAVA 1.5 [24].
+LLM Enables English CLIP to Learn Chinese Without Exposure to Chinese Data.
+To test LLM knowledge transfer, we conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text.
+As shown in Table 3, EVA02 CLIP and JinaCLIP showed near-zero performance, while LLM2CLIP achieved remarkable results, even outperforming models trained on extensive Chinese data like Wukong [12] and CN-CLIP [42].
+This demonstrates the power of LLMs in transferring knowledge to cross-modal spaces.
+4.2.3 Zero-shot Imagenet Classification and Linear Probe Performance
+Evaluation Dataset.
+We conduct zero-shot retrieval testing on Imagenet, following the CLIP method.
+We also perform linear probing, fine-tuning the model with the exact same hyperparameters: batch size of 1024, learning rate of 0.1, momentum 0.9, weight decay 0 and total 50 epochs using SGD.
+Imagenet Performance Better In Linear Probe.
+Our method shows a slight drop in performance on Imagenet for zero-shot classification.
+We hypothesize that this is due to insufficient alignment of the data, resulting in some categories having very low word frequency, as well as the LLM’s limited ability to distinguish short captions.
+However, in the linear probe experiment, we observe that after supplementing the category information, LLM2CLIP actually leads to a greater improvement.
+This demonstrates that LLM2CLIP can indeed provide better visual features.
+One of the current limitations is that scaling the dataset to a larger size might be necessary to resolve the zero-shot classification issue.
+4.2.4 Multimodal Large Language Models Performance
+Evaluation Dataset.
+Following LLAVA1.5 [24], we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B.
+Pretraining included 558K image-caption pairs and 665K visual instruction samples.
+We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction.
+LLM2CLIP Improves VLLM Performance.
+As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks.
+This highlights the potential of LLM2CLIP for improving complex image reasoning and understanding tasks.
+4.3 Main Conclusions
+Directly Replacing CLIP’s Text Encoder with a Vanilla LLM is Harmful.
+Directly replacing the text encoder of EVA02 ViT-L/14 with Llama3-8B significantly degrades retrieval performance, as shown by the red-highlighted numbers in Table 7 and Table 6, with average performance drops of 10%-20%.
+Additionally, the CRA score of the vanilla Llama3-8B is particularly low, indicating that its output features lack sufficient discriminability for captions.
+These results clearly demonstrate that such an approach imposes a significant burden on CLIP’s learning process.
 Improving the Discriminability of LLM Output Features is Key to Integrating LLMs with CLIP.
 To enhance the discriminability of LLM output features, we applied Caption Contrastive fine-tuning.
-Llama3-8B-TC and Llama3-8B-CC represent the results of using pure text corpora and a mix of text and augmented CC3M corpora from ShareCaptioner, respectively, both trained with supervised SimCSE loss.
+Llama3-8B-TC and Llama3-8B-CC represent the results of using pure text corpora(same as LLM2Vec) and a mix of text and augmented CC3M corpora from ShareCaptioner, respectively, both trained with supervised SimCSE loss.
 As shown in the Table 6, the contrastive learning on mixed caption corpora yields higher CRA scores compared to training on generic text, with noticeable improvements across almost all benchmarks.
@@
 These findings highlight the importance of text model discriminability for CLIP training and underscore the rationale behind the caption contrastive fine-tuning in LLM2CLIP.
-LLM2CLIP Makes Pretrained SOTA CLIP Even More SOTA.
-We applied the LLM2CLIP fine-tuning method to both EVA02 and CLIP models and observed significant performance improvements in Table 2.
-Even with lightweight fine-tuning, the results substantially surpassed those of the original models pretrained on datasets like Laion2B.
-Compared to other methods such as LongCLIP and JinaCLIP, which also attempt to fine-tune pretrained CLIP models, our performance gains were transformative.
-This demonstrates the effectiveness of LLM2CLIP as a method for introducing LLMs to enhance CLIP’s performance.
-Table 4: Ablation Study of LLM2CLIP. Here LLM2CLIP-S refers to the results trained on the original short caption dataset.
-Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI
-I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0
-+ Jina-Bert 88.1 77.7 60.5 51.1 83.3 81.0 66.9 68.5 68.9 71.2
-       ++ Dense Caption 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9
-+ Llama3-8B-S 87.9 75.6 56.7 41.8 55.1 46.1 37.2 35.1 39.3 32.3
-       ++ CC Finetuning 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6
-        $+$ Dense Caption 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4
 LLM’s Built-in Knowledge Directly Benefits CLIP.
-In Table 4’s ablation study, we examine the impact of different steps in LLM2CLIP on the EVA02 model.
-When using only the original captions, there was minimal improvement, as these captions are from the same distribution as those used in EVA02 CLIP’s pretraining.
-However, we observed that the LLM, through its open-world knowledge, is able to provide a deeper understanding of the original captions, empowering the CLIP model with enhanced capabilities and further boosting performance.
-Naturally, when dense captions are introduced, the performance improvement becomes even more pronounced.
+In Table 7, we observe that directly using LLM2CLIP leads to significant performance improvements even with short caption datasets, which is quite challenging, as CLIP has already seen most of this training data.
+However, the results from LLM2CLIP with +Llama3-8B-CC still show a clear performance gain, which can be attributed purely to the improvement in the text encoder.
+On the other hand, +Jina-Bert fails to achieve similar results, and in fact, leads to a performance decline.
 LLM Improves CLIP’s Ability to Handle Long and Dense Captions.
 Thanks to the LLM’s inherent long context windows and strong text understanding capabilities, CLIP can process long caption datasets more effectively without requiring any architectural modifications.
-This stands in contrast to methods like LongCaption, which require fine-tuning of positional encodings, or Dreamlip, which splits captions into sub-captions.
 Jina CLIP, for example, enables long-text input by replacing its text encoder with Jina BERT.
@@
 LLM2CLIP explores the method of replacing CLIP’s text encoder with an LLM and continuing training.
-In Table X, we thoroughly investigate the performance differences that various versions of LLMs bring to CLIP.
+We thoroughly investigate the performance differences that various versions of LLMs bring to CLIP.
 In the experiments shown in Table 6, we swapped the text encoder in CLIP with different language models to explore their impact.
 We maintained the original LLM2CLIP design, with the only change being the initialization of the text encoder.
-As expected, within the same Llama family, the 8B model significantly outperforms the 1B model.
-Mistral Namo also shows a slight improvement over the 8B model.
+As expected, within the same Llama family, the 8B model significantly outperfoWrms the 1B model.
+Mistral Memo also shows a slight improvement over the 8B model.
 It’s important to note that during CLIP training, the LLM’s gradients remain frozen, further proving that larger LLMs, with their richer knowledge, can deliver substantial performance improvements to CLIP.
-Even so, the 1B LLaMA3.2 model still achieved an impressive performance boost for EVA02 CLIP, with an average improvement of 11.1% — an impressive feat considering that EVA02 is already one of the SOTA open-source CLIP models.
-Table 5: Comparison Experiment of Different Ratios of Dense Captions in the LLM2CLIP Training Process.
-Ratio Flickr30k COCO ShareGPT4v Urban-1k DOCCI
-I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-100% 85.5 72.7 60.1 46.9 98.7 99.0 88.7 93.9 88.0 90.5
-75% 92.4 82.6 68.5 54.2 98.7 99.3 89.0 94.3 88.1 90.2
-50% 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4
-25% 93.0 82.8 68.1 54.8 98.4 98.7 87.7 92.9 87.9 90.0
-0% 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6
-Table 6: Comparison of various text encoders.
-Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average CRA
-I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-EVA02 Vit-L/14 89.8 73.3 63.8 63.8 89.3 91.9 68.5 73.3 75.0 73.4 76.2 69.8
-+Jina Bert 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 78.2 74.2
-+Llama3-8B 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 66.3 18.4
-+Llama3-8B-TC 92.7 82.1 68.1 54.6 97.7 98.2 88.9 93.8 85.0 87.8 84.8 71.3
-+Llama3-8B-CC 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 85.6 73.0
-+Llama3.2-1B-CC 91.6 81.3 65.8 52.5 98.3 98.2 84.5 91.9 83.4 86.4 83.4 72.8
-+Mistral-Nemo-12B-CC 93.5 83.7 68.5 54.7 98.6 98.9 90.4 94.3 88.0 89.7 86.0 73.3
-Table 7: Performance of Llava 1.5. The best results are highlighted in bold.We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
-MODEL VQA Datasets Pope Metrics MM Benchmarks Seed Benchmarks
-VQAv2 GQA VizWiz SQA-IMG TextVQA Random Adv. Popular MME MMBench MMBench-CN LlavaBench MMVet All IMG Video
-Llava (Paper) 78.5 62.0 50.0 66.8 58.2 87.3 86.1 84.2 1510.7 64.3 58.3 65.4 31.1 58.6 66.1 37.3
-Llava (Rep.) 79.04 62.86 50.57 67.97 57.48 87.7 84.85 86.3 1476.69 66.66 60.39 58.0 34.3 59.86 66.95 39.71
-+LLM2CLIP 79.80 63.15 52.37 69.92 58.35 88.55 82.76 87.75 1505.82 68.29 60.40 62.7 34.8 60.96 68.80 38.96
-Like a miracle: LLM enables English CLIP to learn Chinese without exposure to Chinese data.
-To validate the knowledge transfer capabilities of the LLM, we designed an out-of-distribution (OOD) experiment by conducting an image-text retrieval task in a completely unfamiliar language.
-This is an especially challenging experiment because all of our training was performed on purely English text, yet now we are testing the model on Chinese data.
-As shown in Table 3, models like EVA02 CLIP and Jina CLIP, which performed well on English datasets, achieved near-zero accuracy on this task.
-However, the magical power of LLM2CLIP became evident: the LLM’s capabilities allowed the model to achieve impressive performance on Chinese retrieval tasks, even surpassing models that were originally trained on hundreds millions of Chinese data, such as Wukong (Gu et al., 2022) and CN-CLIP (Yang et al., 2022).
-This result once again demonstrates that LLM2CLIP can effectively integrate the inherent abilities of the LLM into CLIP, enabling it to handle tasks far beyond its original scope.
-LLM2CLIP Can Improve the Performance of VLLM.
-As shown in Table 7, we conducted experiments using the Llava 1.5 Liu et al.
-(2023) VLLM (Vision-Language Large Model) training framework.
-Llava incorporates a CLIP visual encoder into the LLM for multimodal instruction learning, meaning the quality of the visual encoder can significantly impact Llava’s performance.
-We compared the original CLIP with a version enhanced by our LLM2CLIP fine-tuning, running two versions of the experiment according to Llava’s official implementation for a fair comparison.
-The results showed that in over 87.5% of the benchmarks, we achieved substantial performance improvements, with the remaining benchmarks showing results that were very close.
-This demonstrates the potential of LLM2CLIP for complex image reasoning and related tasks.
 4.4 Ablation
@@
 It’s important to note that the dense captions we used were generated by the ShareCaptioner model, and there may be some distribution differences and noise compared to real data captions, which could somehow explain why using 100% dense data is suboptimal.
-Our findings indicate that the best performance is achieved when dense captions make up 50% to 75% of the training data, striking a balance between both short-text and long-text retrieval tasks.
+Our findings indicate that the best performance is achieved when dense captions make up 50% to 75% of the training data.
 Impact of Different Data Size.
@@
 It’s worth noting that training on the 3M dataset takes only about 3 hours on an 8 H100 GPUs machine, yet results in a transformative leap in performance for a well-pretrained CLIP ViT model.
+Adaptor design.
+In Table 9, we explore the impact of the Adaptor and its layer count.
+We find that adding at least one layer of the Adaptor is beneficial, resulting in an average performance improvement of two points.
+However, increasing it to four layers provides minimal additional gains.
+Caption Contrastive Fine-tuning Ablation.
+In Table 10, we explore the impact of the key steps in LLM2Vec’s Caption Contrastive fine-tuning, namely MNTP and bidirectional attention.
+Through comparisons, we observe that MNTP offers almost no positive effect, while bidirectional attention provides some performance gains but is not decisive.
+Therefore, we conclude that the most significant contribution comes from the supervised SimCSE contrastive learning applied to the captions.
 5 Limitations and Future Work
+This paper focuses solely on the impact of LLMs on CLIP.
+In reality, there are many other areas where text can play a transformative role, such as diffusion models [34, 4, 39], Segment Anything [16] and GroundingDINO [25], both of which could potentially see significant performance improvements by applying our method.
 While LLMs possess open-world knowledge, aligning them with CLIP may require specific data to fully unlock their potential.
@@
 By focusing on aspects such as data distribution, length, or categories, we could further tailor the LLM to address CLIP’s limitations, allowing the LLM to act as a more comprehensive teacher for various tasks.
-Moreover, in this work, we froze the gradients of the LLM during fine-tuning to maintain a large batch size for CLIP training.
-However, it would be worth exploring approaches that allow for updating the LLM’s gradients, as there may be better engineering practices that can balance the two objectives.
-Re-training LLM2CLIP from scratch on datasets like Laion-2B (Schuhmann et al., 2022) and Recaption-1B (Li et al., 2024b) is also a promising direction that we did not pursue due to time constraints.
 6 Conclusion
-This paper presents a method that enables LLMs to assist in CLIP training.
 Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation learning, allowing language — the carrier of human knowledge — to play a pivotal role in spreading the intelligence of LLMs across multiple modalities and applications.
@@
 Third, we designed the LLM2CLIP training framework, which has proven to be an efficient and effective approach for significantly enhancing the performance of pretrained CLIP models.
-We experimented with LLM2CLIP on both CLIP and EVA models and validated it across different Llama3 models, including 1B, 8B, and 13B, all of which yielded transformative improvements in CLIP performance.
-LLM2CLIP offers several advantages, such as seamless compatibility with long and complex text inputs, and the ability to incorporate the open-world knowledge of LLMs into CLIP training — as demonstrated by its ability to match Chinese text even when trained solely on English data.
-We firmly believe that the potential of LLMs extends far beyond what has been shown, and we hope this work serves as a starting point for using LLMs to enhance various vision foundation models, ultimately spreading their general capabilities and open-world knowledge across a wide range of modalities and applications.

### v4

@@
 Abstract
-CLIP is a foundational multimodal model that aligns image and text features into a shared space using contrastive learning on large-scale image-text pairs.
-Its strength lies in leveraging natural language as a rich supervisory signal.
-With the rapid progress of large language models (LLMs), we explore their potential to further enhance CLIP’s multimodal representation learning.
-This work introduces a fine-tuning approach that integrates LLMs with the pretrained CLIP visual encoder, leveraging LLMs’ advanced text understanding and open-world knowledge to improve CLIP’s ability to process long and complex captions.
-To address the challenge of LLMs’ autoregressive nature, we propose a caption-to-caption contrastive learning framework to enhance the discriminative power of their outputs.
-Our method achieves substantial performance gains on various downstream tasks, demonstrating the effectiveness of combining LLMs with CLIP for enhanced multimodal learning.
+CLIP is a foundational multimodal model that aligns image and text features into a shared representation space via contrastive learning on large-scale image-text pairs.
+Its effectiveness primarily stems from the use of natural language as rich supervision.
+Motivated by the remarkable advancements in large language models (LLMs), this work explores how LLMs’ superior text understanding and extensive open-world knowledge can enhance CLIP’s capability, especially for processing longer and more complex image captions.
+We propose an efficient post-training strategy that integrates LLMs into pretrained CLIP.
+To address the challenge posed by the autoregressive nature of LLMs, we introduce a caption-to-caption contrastive fine-tuning framework, significantly enhancing the discriminative quality of LLM outputs.
+Extensive experiments demonstrate that our approach outperforms LoRA-based methods, achieving nearly fourfold faster training with superior performance.
+Furthermore, we validate substantial improvements over state-of-the-art models such as CLIP, EVA02, and SigLip2 across various zero-shot multimodal retrieval tasks, cross-lingual retrieval tasks, and multimodal language model pretraining.
 1 Introduction
-CLIP [32] is one of the most influential multimodal foundational models today.
-It aligns visual and language signals into a shared feature space by employing a simple contrastive learning loss on large-scale image-text pairs.
-As a retriever, CLIP supports a wide range of tasks, including zero-shot classification [31], detection [22], segmentation [46], and image-text retrieval [27, 17].
-As a feature extractor, it has become dominant in virtually all cross-modal representation tasks, such as image understanding, video understanding, and text-to-image/video generation.
-For instance, works like LLaVA [24] and Qwen-VL [2] leverage CLIP as a feature extractor to obtain visual features for text models, while models like Stable Diffusion [34] and DALL·E 2 [33] use CLIP’s text encoder to extract textual features for visual models.
-Figure 1: Zero-shot multimodal retrieval radar chart.
-LLM2CLIP significantly improves the performance of original models like CLIP and EVA02 through fine-tuning, substantially outperforming other fine-tuning-based methods such as LongCLIP.
-Figure 2: By adjusting the output space of the LLM, we enable the LLM to more effectively act as a private tutor for CLIP, allowing the fine-tuning process of LLM2CLIP to efficiently adjust the pretrained CLIP’s cross-modal space.
 Figure 1: LLM2CLIP Overview.
@@
 We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision.
+Figure 2: By adjusting the output space of the LLM, we enable the LLM to more effectively act as a private tutor for CLIP, allowing the fine-tuning process of LLM2CLIP to efficiently adjust the pretrained CLIP’s cross-modal space.
+CLIP [28] is one of the most influential multimodal foundational models today.
 What makes CLIP particularly powerful as a vision encoder?
@@
 This allows them to expand on caption information during training, increasing the efficiency of the learning process.
-In this work, we aim to design a method that incorporates the powerful capabilities of LLMs to fine-tune pretrained state-of-the-art (SOTA) CLIP models, further enhancing their performance.
-We propose replacing the weak text encoder of CLIP with an LLM, followed by cross-modal contrastive learning training on image-text datasets.
-While LLMs have already been introduced into CLIP in the InternVL [6] framework, this approach requires large-scale retraining, and the actual performance improvement solely attributable to the LLM remains unclear.
-In our experiments, we found that directly using LLMs to fine-tune CLIP led to failure in all tasks, with a significant drop in performance across various downstream tasks as in Table 6.
-We discovered that the output features of LLMs for different texts (or even the same text) were very similar, which led to poor differentiation.
-This low discriminative power interferes with CLIP’s contrastive learning task.
-To investigate this further, we designed a caption-to-caption retrieval experiment, as shown in Table 1 and Figure 4.
-In this experiment, each image in the MS-COCO dataset has five human-annotated captions.
-We selected the first two captions as positive samples and performed retrieval across the entire test set.
-Using caption retrieval accuracy (CRA), we evaluated the text model’s ability to differentiate between captions, helping us determine which language model is better suited for CLIP.
-We found that Llama-3 8B achieved only 5.2% top-1 accuracy, while the standard CLIP-ViT-L reached 25.2% top-1 accuracy.
-As illustrated in Figure 4, the top-1 caption retrieved by the original Llama-3 model can be entirely unrelated to the query caption, clearly obstructing effective CLIP learning.
-Therefore, directly using an LLM to guide CLIP’s visual encoder training is highly constrained.
-Recent work on LLM2Vec also validates this idea, leading us to realize the necessity of fine-tuning the output space of LLMs.
-Language Model CRA
+This paper successfully explores a efficient method of leveraging LLM capabilities in CLIP pretraining.
+LLMs, as text generation models, are primarily pretrained to capture the distribution of natural language, thus their outputs naturally lie close to textual spaces.
+However, what CLIP pretraining actually requires is a robust text embedding model.
+Previous works such as LLM2Vec [2] and NV-Embed-v2 [17] have demonstrated that fine-tuning LLMs significantly enhances their ability as text encoders.
+Nevertheless, it remains unclear how directly applying LLMs affects CLIP performance.
+Our preliminary results indicate that directly integrating LLMs into CLIP pretraining can be detrimental because LLMs struggle to differentiate effectively between image descriptions, as illustrated by example cases in Figure 3.
+Moreover, as shown in Table 1, we conducted caption-to-caption retrieval experiments on the COCO 5K test set (each image has five captions; we used the first two captions as positive pairs) and observed that large language models like Llama 3 8B achieve very low accuracy, barely distinguishing between different captions of the same image.
+Language Model Top 1 Acc.
 CLIP-L/14 25.2
@@
 Llama3.2-1B-CC 29.4
-Table 1: Comparison of top-1 Caption Retrieval Accuracy (CRA) for various language models on the MS COCO 5K test set.
+Table 1: Comparison of top-1 retrieval accuracy for various language models on the MS COCO [22] 5K test set.
 Figure 3: Real examples of top-1 results from the caption-to-caption retrieval experiment in MS COCO 5K test set. Before fine-tuning, Llama3’s results were often completely unrelated.
-We designed a caption contractive (CC) fine-tuning strategy, as Figure 2, pplying lightweight fine-tuning to the output tokens of Llama-3 8B using LoRA on the CC3M [36] image captioning dataset.
+We designed a caption contractive (CC) fine-tuning strategy, applying lightweight fine-tuning to the output tokens of LLM using image captioning dataset as in  Figure 2.
 The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions.
@@
 In the second stage, we fine-tune the CLIP vision encoder with the CC-fine-tuned LLM, with LLM gradients frozen, to further improve CLIP’s performance.
-Experimental results demonstrate that training on only 15M open-source data substantially boosts CLIP’s capabilities, with even 3M data yielding impressive results.
 The main contributions of this paper are:
@@
   3. 3.
-    We propose LLM2CLIP, an efficient and effective LLM-assisted fine-tuning method for CLIP, enabling low-cost fine-tuning of pretrained CLIP models and achieving significant performance improvements.
+    We successfully verified that LLM2CLIP can substantially enhance the performance of CLIP and EVA02 across multiple tasks, including short-text retrieval, long-text retrieval, cross-language retrieval, and multimodal large language model pretraining, through a simple post-training strategy.
+Remarkably, LLM2CLIP can further improve the short-text retrieval accuracy of state-of-the-art large-scale pretrained models like SigLip2 [33] by +1 and +1.9 points, achieve impressive gains of +14.8 and +15.8 points in long-text retrieval, and deliver significant improvements of +11.9 and +15.2 points in multilingual tasks.
 2 Related Works
@@
 As a result, our model achieves a breakthrough in performance.
-3 Methods
-LLM2CLIP is a fine-tuning strategy for CLIP, consisting of two stages.
-The first stage aims to improve the separability of the LLM’s output space, and the second stage involves fine-tuning by replacing CLIP’s text model with the LLM.
-3.1 Caption Contrastive Fine-tuning
-In this section, we aim to fine-tune the LLM’s token outputs to better capture features that can distinguish between image captions.
-The process of improving the discriminability of LLM output features on caption text is quite straightforward: we need the distances between captions of the same image to be closer, and those of different images to be further apart.
-Therefore, we apply a caption contrastive (CC) fine-tuning to the LLM’s output features, treating different captions of the same image as positive samples and the rest of the captions as negative samples.
-To obtain enough varied descriptions for the same image, we use the ShareCaptioner [45, 5] modified CC-3M [36] dataset, which provides both original captions and augmented dense captions for each image.
-These can be treated as positive pairs.
-We followed the training methodology of LLM2Vec [3], first expanding the LLM’s attention mechanism to bidirectional attention, and employing Masked Next Token Prediction (MNTP) for initialization to achieve better results.
-Specifically:
-  1. 1.
-    Bidirectional Attention Transformation: First, we transform the LLM’s causal attention mechanism into bidirectional attention, as we no longer need it to retain generative capabilities but rather function as an encoder.
-Since autoregressive training is no longer required, switching to bidirectional attention improves its ability to capture contextual information.
-  2. 2.
-    Masked Next Token Prediction (MNTP): Second, we employ MNTP to train the newly added bidirectional attention mechanism, providing a strong initialization.
-For a given sequence of $N$ tokens, we mask a subset and predict their values, similar to BERT [7].
-However, unlike BERT, we adapt this process to fit the nature of LLMs by predicting tokens just before the masked token.
-We train on both image captions and pure text with equal weighting.
-In addition to CC-3M, we also use the Wikitext-103 [28] dataset to preserve the LLM’s text capabilities and prevent divergence from its original strengths.
-  3. 3.
-    Caption Contrastive Learning: Finally, we perform the actual caption contrastive fine-tuning, using a supervised SimCSE loss to bring captions of the same image closer together and push apart captions of different images.
-We use two prompt templates: “Given a caption, retrieve a detailed relevant caption” and “Given a detailed caption, retrieve a short relevant caption,” which are prepended to the query (either original or dense caption) to retrieve the corresponding dense or original caption.
-Similarly, we use a 1.5M-paired general text dataset curated from [37] to maintain strong performance in pure language tasks.
-All training is efficiently conducted using LoRA and is completed in just one epoch, ensuring low computational cost.
-In our experiments, we found that the most critical step was the third step with the SimCSE loss, which significantly enhanced the caption discriminability of the LLM.
-3.2 LLM2CLIP: Use LLM to Finetuning
-With the modifications to the LLM discussed earlier, we have now obtained a super text encoder that is well-suited for CLIP training.
-The next step is to use this LLM in conjunction with the pretrained state-of-the-art CLIP visual encoder to reconstruct a more powerful cross-modal feature space.
-As shown in  Figure 3, in the LLM2CLIP training phase, we freeze the gradients of the LLM to preserve its inherent capabilities, for two primary reasons.
-First, this significantly reduces the computational cost and memory footprint of fine-tuning.
-CLIP training requires a very large batch size to maintain the effectiveness of negative samples.
-Allocating memory to the LLM could compromise CLIP’s performance.
-Second, by freezing the LLM, we ensure that the open-world knowledge it has acquired from large-scale corpora remains intact during the multimodal alignment process.
-To compensate for the frozen LLM, and inspired by methods like FuseMix [41] and APE [35], we introduce several new linear layers after the LLM as adapters.
-These layers serve as learnable parameters to improve the alignment between the LLM and CLIP visual encoder.
-Following the original design of CLIP, we also employ a projector layer to align the dimensions of the two encoders, facilitating the use of CLIP loss for training.
-With this powerful LLM-based super text encoder, we achieve a substantial qualitative leap in CLIP’s language comprehension capabilities as results in Table 2.
-The LLM’s open-world knowledge allows the CLIP visual encoder to learn more structured and globally informed visual representations that are aligned with human knowledge.
-Moreover, this approach enables us to fully utilize high-quality, long, and dense caption datasets without requiring any special architectural adjustments, which previous works like DCI, DreamLip, and Recaption struggled to effectively leverage.
-3.3 Overview and Efficiency Discussion
-We propose LLM2CLIP as a method that efficiently incorporates large language models (LLMs) into CLIP training, leveraging the capabilities of LLMs to make cross-modal representation learning significantly more powerful.
-In our experiments, We evaluated large language models, including Llama with 1B and 8B parameters, as well as Mistral-Nemo with 12B parameters.It might seem that incorporating such large LLMs would greatly increase the computational burden of training CLIP, especially since CLIP itself is computationally expensive, requiring a large batch size.
-However, our proposed LLM2CLIP is remarkably lightweight.
-The training overhead isnearly identical to fine-tuning the original CLIP model, with minimal additional cost, yet the LLM provides much stronger supervision.
-Here, we highlight some of the design details that significantly improve training efficiency: 1).
-During the caption contrastive fine-tuning stage, we employ LoRA training for the LLM.
-Even for a 12B LLM, training with a batch size of 512 only requires around 70GB of GPU memory, making it feasible to run on a single 80GB 8 A100 GPU node.
-2).
-In the LLM2CLIP stage, we freeze the LLM’s gradients and only train the learnable adapter, CLIP’s original Vision encoder, and two projectors.
-The additional trainable parameters are roughly equivalent to those in the original CLIP, minimizing the overhead.
-To further reduce the inference cost of using LLMs, we pre-extract all the text features from the training data and store them in disk memory.
-This ensures that, even though LLMs provide powerful textual supervision, the memory and computational costs during training remain nearly identical to those of standard CLIP training.
-For instance, when we trained LLM2CLIP using a Mistral-Nemo [^1] https://mistral.ai/news/mistral-nemo/ 12B model integrated with the commonly used EVA ViT-L/14-224, with a batch size of 4096 on 8 H100 GPUs, the memory usage per GPU was only 30GB, and the entire training process took just 9 hours.
-Despite this efficient training cost, LLM2CLIP brought transformative improvements in downstream tasks such as long and short text retrieval, cross-language retrieval, and Llava training.
-Methods Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
+Methods Resolution Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
 I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-ViT-B/16
-ALIGN 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
-BLIP 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
-Jina-CLIP 80.6 67.4 55.6 41.1 - - 87.7 88.0 78.7 80.0 75.7 69.1
-Long-CLIP 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
-CLIP 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
-+LLM2CLIP 89.2 78.1 62.2 48.7 98.1 97.4 86.1 90.0 84.1 85.0 83.9 79.8
-EVA02 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
-+LLM2CLIP 88.5 78.0 63.6 49.8 98.0 98.1 84.7 89.7 85.5 86.8 84.1 80.5
-ViT-L/14
-Long-CLIP 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
-CLIP 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
-+LLM2CLIP 92.6 81.7 64.9 52.5 98.4 98.4 87.6 92.0 87.6 88.7 86.2 82.7
-EVA02 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
-+LLM2CLIP-3M 89.6 77.3 59.7 48.0 98.3 98.6 87.1 91.1 84.9 87.8 83.9 80.6
-+LLM2CLIP 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
-+LLM2CLIP-30M 92.0 83.5 69.0 55.3 98.9 98.8 93.1 95.0 89.3 91.2 88.5 84.8
-+LLM2CLIP-60M 94.4 83.2 70.4 55.7 99.2 99.4 94.1 95.2 90.2 92.0 89.7 85.1
-ViT-L/14-336
-CLIP 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
-+LLM2CLIP 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0 86.6 83.3
-+LLM2CLIP-60M 93.9 82.3 68.5 54.8 98.9 99.1 94.6 95.9 89.6 90.6 89.1 84.5
-EVA02 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
-+LLM2CLIP 93.9 83.8 68.7 55.7 98.8 99.2 89.5 94.2 89.2 91.3 88.0 84.8
-Table 2: Systematic Comparison Experiment Demonstrating the Performance Improvements of LLM2CLIP. Results with a blue background represent the outcomes of LLM2CLIP.
-4 EXPERIMENTS
-4.1 Experiment Setting
-4.1.1 Caption Contrastive Fine-tuning
-Training Dataset.
-For contrastive learning, we utilized the ShareCaptioner-modified CC-3M dataset [45, 5], which includes both original captions and augmented dense captions for each image.
-In addition, we used the Wikitext-103 dataset [28] and the E5 dataset [37] during the Masked Next Token Prediction and caption contrastive fine-tuning stages to ensure coverage of the general text domain without excessive bias.
-Training Setting.
-The language model was trained with LoRA for lightweight training, completing one epoch across all datasets.
-We used the average of all LLM output tokens as the global text embedding for a caption.
-All training parameters followed the configuration detailed in [3].
-The design of the Adaptor is the same as FuseMix [41], using a simple inverted bottleneck MLP architecture.
-4.1.2 LLM2CLIP Fine-tuning
-Training Dataset.
-We defined three experimental configurations based on dataset size:
-  • LLM2CLIP-3M: This configuration uses only the CC-3M dataset, serving as our lightweight setting.
-  • LLM2CLIP-15M: Our default configuration, combining CC-3M and CC-12M datasets.
-  • LLM2CLIP-60M: A larger configuration that combines CC-3M, CC-12M, YFCC-15M, and a 30M randomly selected subset from Recaption-1B [20].
-All datasets included dense captions rewritten by multimodal large language models (MLLMs).
-Specifically, the CC-3M, CC-12M, and YFCC datasets were sourced from [45] and rewritten using ShareCaptioner, while Recaption data were rewritten with Llama3-LLAVA1.5.
-A 1:1 mix of original captions and dense captions was used by default.
-Training Setting.
-Each configuration was trained for four epochs.
-We froze the LLM gradients to reduce computational overhead and pre-extracted text features using the LLM, storing them on disk to avoid repeated inference during training.
-Inspired by Fusemix’s efficient training strategy, we added a four-layer linear adapter after the LLM and trained with gradients enabled for the CLIP ViT model to better integrate LLM knowledge.
-Additionally, we introduced two projector layers after each encoder for dimensional alignment.
-Training the CLIP ViT-L/14-224 model required approximately 30GB of GPU memory per GPU and took 2, 9, and 45 hours on 8 H100 GPUs for the 3M, 15M, and 60M datasets, respectively.
-We employed the AdamW [26] optimizer, utilizing a total batch size of 4096, and set the learning rate to 1e-5 with a cosine decay rate of 0.05.
-For language models, we tested four configurations: Jina-Embeddings-V2, and three popular models from the Llama3 family [8]: Llama 3.2 1B, Llama 3 8B, and Mistral-Nemo 12B.
-Unless otherwise noted, experiments used EVA02 ViT-L/14 as the baseline and Llama 3 8B as the default language model.
-Methods Flickr-CN COCO-CN
-I2T@1 I2T@5 I2T@10 T2I@1 T2I@5 T2I@10 I2T@1 I2T@5 I2T@10 T2I@1 T2I@5 T2I@10
-ViT-L/14-336
-Wukong 76.1 94.8 97.5 51.7 78.9 86.3 53.4 80.2 90.1 55.2 81.0 90.6
-CN-CLIP 80.2 96.6 98.2 68.0 90.7 95.4 63.4 84.2 92.9 64.0 89.2 94.4
-JinaCLIP 3.30 9.90 15.1 0.7 3.5 6.0 2.9 8.9 13.7 1.0 4.9 8.2
-EVA02 4.40 11.8 16.7 0.94 2.9 4.8 2.7 9.8 15.2 1.0 3.7 7.3
-+LLM2CLIP 86.9 98.1 99.3 75.1 92.9 96.0 69.1 92.5 97.2 70.0 92.6 96.7
-Table 3: Retrieval Performance across Flickr30K-CN and COCO-CN.
+ViT-B/16(86M)
+ALIGN [11] 224 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
+BLIP [19] 224 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
+Long-CLIP [39] 224 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
+jina-clip-v2 [15] 224 84.4 69.8 57.1 41.9 94.5 91.1 80.4 78.0 77.6 78.2 78.8 71.8
+SigLIP [38] 224 89.1 74.7 65.8 47.8 85.8 83.5 62.8 62.4 70.1 70.6 74.7 67.8
+MetaCLIP [36] 224 85.6 70.8 59.3 41.3 90.4 86.6 68.9 63.3 70.9 71.5 75.0 66.7
+CLIP [28] 224 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
+     +LLM2CLIP-15M 224 91.8 80.7 64.8 52.4 97.8 97.6 90.1 92.1 86.1 86.9 86.1 81.9
+EVA02 [31] 224 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
+     +LLM2CLIP-15M 224 91.2 81.4 66.0 53.9 98.5 98.5 88.2 91.3 86.8 88.5 86.1 82.7
+ViT-L/14 (307M)
+Long-CLIP [39] 224 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
+MetaCLIP [36] 224 90.1 76.4 64.5 47.1 86.4 79.5 73.4 70.0 76.5 76.7 78.2 69.9
+SAIL-L-NV2 [40] 224 87.9 75.8 62.4 48.7 94.8 94.9 81.5 80.2 76.5 78.9 80.6 75.7
+CLIP [28] 224 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
+     +LLM2CLIP-15M 224 93.0 83.5 67.4 56.4 98.7 98.6 91.5 94.1 88.3 90.4 87.8 84.6
+EVA02 [31] 224 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
+     +LLM2CLIP-3M 224 94.3 84.3 65.2 56.0 98.0 98.0 91.0 94.7 87.8 90.7 87.3 84.7
+     +LLM2CLIP-15M 224 94.6 85.0 69.5 58.3 98.9 99.1 93.6 95.7 89.8 91.2 89.3 85.8
+     +LLM2CLIP-60M 224 95.9 85.1 71.7 58.5 99.2 99.2 95.2 97.0 90.1 92.0 90.4 86.3
+CLIP [28] 336 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
+     +LLM2CLIP-15M 336 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0 86.6 83.3
+     +LLM2CLIP-60M 336 93.9 84.3 68.5 54.8 98.9 99.1 96.6 97.9 89.6 91.6 91.1 84.5
+EVA02 [31] 336 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
+     +LLM2CLIP-15M 336 94.7 85.1 69.7 58.5 98.8 99.2 93.5 96.2 89.2 91.3 88.0 84.8
+     +LLM2CLIP-60M 336 95.9 85.4 72.4 58.8 98.8 99.2 95.5 97.3 90.7 92.6 91.0 86.8
+SigLIP [38] 384 93.7 81.4 72.0 53.9 90.1 90.4 76.3 74.4 77.8 79.5 82.0 75.9
+SO/14(428M)
+SigLIP [38] 224 91.0 75.2 69.8 51.8 56.1 50.0 31.5 29.1 42.3 44.8 58.1 50.2
+SigLIP2 [33] 224 93.9 82.9 72.0 55.5 90.2 87.2 75.7 74.5 77.0 78.9 81.8 75.8
+     +LLM2CLIP-60M 224 95.2 84.6 73.7 57.6 99.4 99.4 96.2 97.3 91.5 92.9 91.2 86.4
+SigLIP [38] 384 94.3 83.0 72.4 54.3 91.6 89.4 74.5 73.6 74.5 75.8 81.5 75.2
+InternVL(6B)
+InternVL [4] 224 94.2 81.6 69.4 53.2 92.4 90.5 79.6 77.7 79.2 81.6 83.0 76.9
+Table 2: Systematic comparison of model performance on multiple datasets. The light blue row indicate LLM2CLIP post trained version.
+Models Flickr-CN [16] COCO-CN [20] XM3600 [32]
+I2T T2I I2T T2I I2T T2I
+CN-CLIP 80.2 68.0 63.4 64.0
+EVA-L-224 4.4 0.9 2.6 1.0 14.0 8.0
+     +LLM2CLIP 90.6 75.6 72.0 70.1 68.3 56.0
+Siglip2 79.2 56.9 55.3 51.7 59.7 48.2
+     +LLM2CLIP 90.0 76.1 70.8 70.2 69.1 56.3
+Table 3: Multi-lingua retrieval results on different datasets.
 Methods Imagenet
@@
      +LLM2CLIP 79.80 63.15 52.37 69.92 58.35 88.55 82.76 87.75 1505.82 68.29 60.40 62.7 34.8 60.96 68.80 38.96
-Table 5: Performance of Llava 1.5 benchmarks. The best results are highlighted in bold.We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
-Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average CRA
+Table 5: Performance of Llava 1.5 benchmarks.
+The best results are highlighted in bold.
+We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
+Methods Flickr COCO ShareGPT4V Urban-1k DOCCI Average
 I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5 27.1
-+Jina Bert 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0 29.68
-+Llama3-8B 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0 5.22
-+Llama3-8B-TC(LLM2Vec) 92.7 82.1 68.1 54.6 97.7 98.2 88.9 93.8 85.0 87.8 86.5 83.3 28.42
-+Llama3-8B-CC 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2 29.46
-+Llama3.2-1B-CC 91.6 81.3 65.8 52.5 98.3 98.2 84.5 91.9 83.4 86.4 84.7 82.1 29.42
-+Mistral-Nemo-12B-CC 93.5 83.7 68.5 54.7 98.6 98.9 90.4 94.3 88.0 89.7 87.8 84.3 29.96
-Table 6: Comparison of various text encoders.
-Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
+CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
+Directly Finetune(50%) 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
+bge-en-icl 89.1 78.5 58.8 49.7 95.0 95.7 77.9 87.7 73.5 79.4 78.9 78.2
+LLM2Vec-Llama-3-8B 91.1 81.4 61.5 51.9 94.5 96.4 82.1 88.6 77.7 82.6 81.4 80.2
+NV-Embed-v2 90.4 80.0 60.5 51.6 94.5 95.7 83.3 90.0 78.3 82.1 81.4 79.9
+Qwen2.5-0.5B-CC 86.2 74.9 56.0 45.1 92.6 93.2 73.7 79.0 69.5 73.0 75.6 73.0
+Llama-3.2-1B-CC 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
+Llama-3-8B-CC 90.4 80.7 62.7 51.9 96.5 96.2 84.2 89.5 83.3 86.4 83.4 80.9
+DeepSeek-R1-Distill-Llama-8B-CC 91.7 80.9 62.2 51.9 96.7 96.3 84.3 88.3 82.5 85.2 83.5 80.5
+Llama-3.1-8B-CC 92.2 81.5 63.5 52.3 97.1 96.2 86.5 89.3 84.7 85.9 84.8 81.0
+Llama3.1-8B 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
+Table 6: Ablation for using different text encoder.
+BGE-EN-ICL, LLM2Vec-Llama-3-8B, and NV-Embed-v2 are community-released text encoders derived from LLMs.
+The suffix ”-CC” indicates encoders that have undergone our caption contrastive fine-tuning method based on the respective original LLMs.
+Methods(MLLM’s Caption Ratio) Flickr COCO ShareGPT4V Urban-1k DOCCI Average
 I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
-+ Jina-Bert 88.1 77.7 60.5 51.1 83.3 81.0 66.9 68.5 68.9 71.2 73.5 70.0
-       ++ Dense Caption 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0
-+ Llama3-8B 87.9 75.6 56.7 41.8 55.1 46.1 37.2 35.1 39.3 32.3 55.2 46.2
-       ++ Dense Caption 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0
-+ Llama3-8B-CC 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
-       ++ Dense Caption 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
-Table 7: Ablation analysis of performance variations between short simple captions and long & dense captions using different text encoders.
-We conducted experiments on the EVA02 ViT-L/14 model, replacing its original text encoder with Jina-BERT (+ Jina-BERT), the original Llama3 (+ Llama3-8B), and Llama3 fine-tuned with Caption Contrastive learning (+ Llama3-8B-CC).
-Fine-tuning was performed on datasets containing only short simple captions and datasets with a 50% probability mix of long dense captions.
-The objective of this experiment was to isolate the effects of LLMs and long dense captions, and to demonstrate that the design of LLM2CLIP enhances the efficiency of dense caption utilization.
-The results with a red background indicate that the LLM without fine-tuning has a significant negative effect on CLIP.
-Ratio Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
+CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
+Directly Finetune(50%) 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
+       +LLM2CLIP (100%) 89.2 77.2 55.6 47.6 97.6 98.0 88.5 92.5 82.6 86.9 82.7 80.4
+       +LLM2CLIP (75%) 92.0 81.3 61.5 52.0 97.7 97.8 85.1 92.2 84.0 87.7 84.1 82.2
+       +LLM2CLIP (50%) 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
+       +LLM2CLIP (25%) 91.1 81.8 61.8 53.0 97.3 97.2 83.4 90.6 83.9 86.9 83.5 81.9
+       +LLM2CLIP (0%) 91.1 81.5 60.9 53.0 96.0 94.4 72.0 85.6 82.6 85.6 80.5 80.0
+Table 7: Ablation study on the proportion of dense captions from different MLLMs.
+Here, we explore the impact of both genuine image captions and fake dense captions generated by MLLMs on the performance of  LLM2CLIP.
+Stage 1 Stage 2 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
 I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-100% 85.5 72.7 60.1 46.9 98.7 99.0 88.7 93.9 88.0 90.5 84.2 80.6
-75% 92.4 82.6 68.5 54.2 98.7 99.3 89.0 94.3 88.1 90.2 87.3 84.1
-50% 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
-25% 93.0 82.8 68.1 54.8 98.4 98.7 87.7 92.9 87.9 90.0 87.0 83.8
-0% 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
-Table 8: Comparison Experiment of Different Ratios of Dense Captions in the LLM2CLIP Training Process.
-Adaptor Layers Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
+89.2 77.7 59.0 47.3 94.3 94.3 75.0 80.5 74.0 77.5 78.3 75.5
+Linear(×$\times 1$) 88.5 77.5 59.8 48.4 95.1 94.5 77.2 84.4 75.3 78.7 79.2 76.7
+Linear(×$\times 2$) 89.6 78.1 59.6 48.5 95.5 94.7 79.7 84.0 76.1 79.0 80.1 76.8
+Linear(×$\times 4$) 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
+Linear(×$\times 4$) Linear(×$\times 4$) 89.9 78.8 59.7 48.9 95.1 95.6 79.9 84.9 78.0 80.4 80.5 77.7
+Transformer(×$\times 1$) 88.6 78.7 59.9 48.6 95.8 95.1 79.5 84.5 77.0 79.8 80.2 77.3
+Transformer(×$\times 1$) Transformer(×$\times 1$) 89.8 78.6 58.1 49.0 96.1 95.4 81.9 85.6 76.5 78.1 80.5 77.3
+[…]
+3 Methods
+3.1 Stage 1: LLM Caption Contrastive Fine-tuning
+Traditionally, LLM output layers serve as classification heads producing text tokens.
+Prior works, such as LLM2Vec [2] and NV-EMBED-v2 [17], have explored converting LLMs into universal text encoders by applying additional designs and fine-tuning.
+Inspired by these approaches, we adapt similar concepts to the novel scenario of cross-modal pretraining.
+Our key design considerations encompass model architecture, training methods, and training data, summarized as follows:
+Model Architecture:
+  • Sentence Token Representation: We map the output layer features of the LLM to represent whole sentences.
+We considered two strategies: using the [EOS] token, which can attend to the entire sentence during pretraining, or employing average pooling across all output tokens.
+Empirically, average pooling performed better and is our default setting.
+  • Bidirectional Attention: Given that the generative capability is unnecessary for our encoder-focused application, we remove the attention mask from the LLM, enabling bidirectional textual relationship modeling for enhanced comprehension.
+  • Fine-tuning Parameters: To efficiently activate stronger textual comprehension capabilities within the output features of LLMs, we apply parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA).
+  • Adaptor Design: Adding an adaptor following the LLM can potentially enhance feature separability.
+We experimented with two adaptor variants: (1) a Transformer-based adaptor composed of several latent cross-attention transformer layers following NV-EMBED-v2, and (2) a Linear adaptor consisting of multiple linear layers directly attached after extracting sentence tokens from LLM outputs.
+However, neither approach significantly boosted CLIP performance during embedding fine-tuning.
+Consequently, we opted not to use adaptors during the LLM caption contrastive fine-tuning phase for simplicity.
+Training Methods:
+  • Masked Next Token Prediction (MNTP): Inspired by masked language modeling in BERT, MNTP improves feature quality by masking and predicting specific tokens.
+Unlike BERT, predictions are always made at the position preceding the masked token, aligning with LLM’s next-token prediction convention.
+While effective in pure text tasks (as shown by LLM2Vec), MNTP alone or combined with contrastive learning underperformed relative to direct contrastive fine-tuning in multimodal scenarios.
+  • Textual Contrastive Learning: Following SimCSE, we employ textual contrastive learning to enhance feature separability, pulling positive samples closer and pushing negative samples apart in feature space.
+Two variants are possible: unsupervised (using dropout augmentation) and supervised (leveraging semantic pairs from annotated data).
+We default to supervised contrastive learning due to superior performance.
+Specifically, we generate positive pairs using two distinct captions from the same image, framed by a system prompt: ”Given a caption, retrieve a similar relevant caption.”
+Training Data: We conducted experiments with Dreamlip [41]’s caption dataset, naturally containing multiple captions per image, making it ideal for supervised textual contrastive learning.
+Additionally, to ensure the generalizability of LLM textual differentiation capabilities, we incorporated a 1.5M paired pure-text dataset curated by Echo Embeddings [29] into our contrastive training process.
+89213
+3.2 Stage 2: LLM2CLIP Post Fine-tuning
+We aim to perform post fine-tuning on pretrained CLIP since we believe the fundamental visual-language mapping has been well-established through CLIP’s extensive pretraining.
+By introducing an LLM, we expect to further enrich this representation space, potentially improving CLIP’s capabilities through a highly cost-effective post-training step.
+CLIP inherently consists of a Vision Encoder (typically a ViT) and a relatively small Text Encoder (roughly one-third the parameters of the Vision Encoder), structured as a compact autoregressive model.
+Integrating a caption-contrastive fine-tuned LLM as an additional text encoder and determining the optimal approach to enhance CLIP performance is critical.
+Here, we elaborate on model architecture, training methods, and training data.
+Model Architecture:
+  • Fine-tuning Parameters and Adaptor Design: Training CLIP typically requires large-scale datasets and significant batch sizes, resulting in considerable computational costs.
+Therefore, we aim to employ efficient Parameter-Efficient Fine-Tuning (PEFT) methods for this post-training stage.
+By default, we enable full gradients for the visual encoder to facilitate learning from LLM knowledge.
+Given the large size of the LLM, we compare two fine-tuning strategies: (1) directly fine-tuning the LLM using LoRA, and (2) freezing the LLM gradients and appending previously introduced Transformer or Linear adaptors after its output layer as learnable modules to obtain the sentence token features.
+Method (2) offers distinct advantages, including completely eliminating gradient updates for the LLM, thus significantly reducing GPU memory consumption.
+Moreover, since the adaptor is placed after the LLM, we can even precompute and store text features through offline inference, thereby reducing the inference overhead from multiple epochs to just a single pass.
+Consequently, this approach eliminates the need to load the LLM into GPU memory during training, greatly accelerating experiment iteration and training time, and drastically reducing GPU memory footprint.
+By default, we adopt method (2) with an adaptor consisting of four linear blocks.
+Detailed efficiency analysis will be presented in Section 4.4.
+Training Methods: We tested several approaches:
+  • (a) Removing the original CLIP Text Encoder and using LLM with CLIP Vision Encoder directly for cross-modal contrastive learning.
+  • (b) Retaining both LLM and original Text Encoder, applying separate CLIP losses to both encoders.
+  • (c) Adding an additional contrastive learning task between LLM and the original Text Encoder to method (b).
+  • (d) Concatenating LLM and Text Encoder output features in method (b), then performing contrastive learning with CLIP Vision Encoder.
+Extensive experiments (see Table 10) led us to adopt method (a) for its simplicity and effectiveness.
+Training Data: With the integration of LLM, our method inherently excels at processing dense and detailed image captions.
+Datasets such as DreamLIP and Recaption, which leverage Share-Captioner [3], InstructBLIP [5] and LLaVA-1.5 [23] to generate multiple dense captions per image, can be effectively utilized.
+We explored multiple caption scenarios, employing multiple positive caption examples during contrastive training and blending real and MLLM-generated dense captions.
+Ultimately, we adopted a strategy blending real captions and dense MLLM captions at a 50% ratio.
+Stage 1 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
 I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-0 91.3 79.9 64.0 50.2 97.1 97.7 88.1 90.2 81.0 86.1 84.3 80.8
-1 92.1 80.6 65.9 51.8 97.8 98.0 90.9 91.8 84.3 87.9 86.2 82.0
-4 92.1 79.9 66.0 52.2 97.4 98.0 93.0 92.1 85.0 88.0 86.7 82.0
-Table 9: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different adaptor layer numbers.
-Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks.
-The best result in each column is highlighted in bold.
-All experiments were conducted using OpenAI’s CLIP model with a ViT architecture at 224 resolution, trained on 15M data points.
-Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
+Lora, AvgPool, Bidirectional, Supervise Simcse 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
+Lora, AvgPool, Bidirectional, Un-supervise Simcse 77.2 66.0 48.8 38.2 80.2 86.2 47.7 53.5 41.9 44.7 59.2 57.7
+Lora, AvgPool, Bidirectional, MNTP 82.0 71.7 51.2 43.1 92.0 91.9 67.0 70.3 58.4 58.2 70.1 67.0
+Lora, AvgPool, Bidirectional, MNTP, Supervise Simcse 88.2 78.3 59.6 48.9 95.5 95.1 79.7 86.1 75.3 77.7 79.7 77.2
+Lora, AvgPool, Casual, Supervise Simcse 89.4 78.7 59.9 49.0 95.4 95.7 79.9 84.9 75.3 79.1 80.0 77.5
+Lora, EOS, Bidirectional, Supervise Simcse 89.6 78.6 59.4 48.5 95.3 95.4 79.8 83.8 76.0 80.0 80.0 77.3
+Frozen, Linear Adaptor, AvgPool, Bidirectional, Supervise Simcse 85.5 75.5 54.1 45.3 92.0 93.5 72.3 75.5 66.4 66.9 74.1 71.3
+Table 9: Ablation study on the training methods of LLM caption contrastive finetuning in Stage 1.
+Rows with a light blue background represent our default setting, and red text indicates content that differs from the default setting.
+Methods Training Loss Testing Text Encoder Flickr COCO ShareGPT4V Urban1k DOCCI Average
 I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-Bidirectional Attention + MNTP + Sup. SimCLR 83.3 73.5 55.3 45.6 98.1 98.0 81.5 87.9 78.8 82.6 79.4 77.5
-Bidirectional Attention + Sup. SimCLR 84.6 74.5 55.6 46.2 98.0 97.6 79.3 88.5 80.0 82.6 79.5 77.9
-Sup. SimCLR 82.0 73.5 54.6 44.8 97.6 97.6 82.2 87.9 76.9 81.5 78.7 77.1
-Table 10: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different methods.
-Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks.
-The best result in each column is highlighted in bold.
-The experiments were conducted using the Llama 3.2 1B and the EVA02 ViT-L/14 model on 3M data.
-It can be observed that MNTP has no significant effect, while bidirectional attention provides a slight performance improvement.
-4.2 Main Results
-Evaluation Setting.
-Our experiments assessed LLM2CLIP’s ability to enhance vanilla CLIP performance across multiple dimensions.
-We compared our approach against state-of-the-art (SOTA) models, including EVA [10] and OpenAI CLIP.
-Specifically, we evaluated CLIP ViT-B/16, ViT-L/14, and ViT-L/14-336, along with EVA02 ViT-B/16, EVA02 ViT-L/14, and EVA02 ViT-L/14-336, yielding a total of six comparison models.
-4.2.1 Short- and Long-Text Retrieval
+CLIP CL(CLIP-T, CLIP-V) CLIP-T 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
+Directly Finetune CL(CLIP-T, CLIP-V) CLIP-T 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
+   +LLM2CLIP a) CL(LLM, CLIP) LLM 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
+b) CL(LLM, CLIP-V)+ CL(CLIP-T, CLIP-V) CLIP-T 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
+LLM 90.1 81.3 61.3 52.4 97.1 97.4 85.6 91.1 83.8 86.9 83.6 81.8
+c) CL(LLM, CLIP-V)+CL(CLIP-T, CLIP-V)+ CL(CLIP-T, LLM) CLIP-T 89.0 77.9 58.5 47.7 89.1 88.8 67.0 75.5 66.3 70.7 74.0 72.1
+LLM 90.8 80.5 61.5 51.8 97.1 97.7 85.9 90.3 83.4 86.7 83.7 81.4
+d) CL(LLM, CLIP-V)+CL(CLIP-T, CLIP-V)+ CL(Cat(CLIP-T, LLM), CLIP-V) CLIP-T 89.6 77.9 59.1 48.1 89.4 87.5 69.1 72.5 67.0 70.6 74.8 71.3
+LLM 90.7 81.4 61.3 52.3 97.5 97.9 85.5 92.5 83.9 87.5 83.8 82.3
+Cat(CLIP-T,LLM) 91.3 82.4 62.9 53.3 97.8 97.4 86.6 92.9 84.9 88.2 84.7 82.8
+Table 10: LLM2CLIP Training Method Analysis.
+We experimented with various possibilities for post-training between the LLM and the pretrained CLIP model’s Vision Encoder (CLIP-V) and Text Encoder (CLIP-T).
+In the figure, ”CL” denotes the contrastive learning loss, and ”Cat” stands for concatenation.
+Items a–d correspond to the training methods section in Section 3.2.
+In this experiment, we performed zero-shot image-text retrieval tests on each text encoder, including the new feature projection space formed after concatenation.
+Model Strategy Training Hour Batchsize Flickr COCO ShareGPT4V Urban1k DOCCI Average
+I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I i2t t2i
+Llama-3.1-8B LLM LoRA 17 768 90.0 80.9 60.9 50.6 97.5 96.9 88.9 90.8 82.0 84.3 83.9 80.7
+LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
+Llama-3.1-8B-CC LLM LoRA 17 704 90.5 81.8 61.5 52.4 98.1 98.2 91.0 92.8 85.7 87.4 85.4 82.5
+LLM Frozen + Linear Adaptor 5.5 4096 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
+LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 93.2 82.2 64.2 54.0 98.2 98.3 87.8 93.1 86.2 88.8 85.9 83.3
+Table 11: Efficiency study of LLM2CLIP.
+We use two nodes, each equipped with 8 A100 40GB GPUs, and conduct experiments by maximizing GPU memory utilization to achieve the largest possible batch size.
+We report training hours, batch size, and retrieval accuracy for each experiment.
+During offloading, we precompute text captions offline into embeddings using the LLM.
+This approach is a one-time preprocessing step that effectively reduces the experimental overhead.
+4 Experiments
+4.1 Default Experimental Settings
+Default Settings for Caption Contrastive Fine-tuning We choose to use LLaMA 3.1 8B as the LLM integrated into CLIP pre-training, removing the causal attention mask to enable bidirectional attention.
+No additional adapters are used in caption contrastive finetune, and average pooling is employed to obtain the sentence token embedding.
+MNTP is not used by default.
+We utilize 30M DreamLIP caption data to perform caption contrastive fine-tuning on the LLM.
+Randomly sampling any two captions of an image from DreamLIP as positive pairs, we conduct supervised SimCSE training.
+Default Settings for LLM2CLIP Cross-modal Post-training After caption contrastive fine-tuning, we replace the original CLIP text encoder with the LLaMA 3.1 8B model to perform cross-modal contrastive learning post-training with the vision encoder.
+By default, we open the ViT gradients and freeze the LLM gradients, then add an adapter layer composed of four linear layers after the LLM.
+The design of the Adaptor is the same as FuseMix [35], using a simple inverted bottleneck MLP architecture.
+By default, we use 15M DreamLIP annotated subsets of CC 3M and CC 12M.
+For the 60M data setting, we use DreamLIP re-annotated CC 3M and CC 12M, YFCC 15M, and a 30M subset of the LAION dataset.
+For the 3M setting, we only use the CC 3M subset.
+We summerize our default experiments setting and implementation details in Appendix A.
+4.2 System Comparison
+In the system comparison, we used the default settings from  Section 4.1 throughout.
+However, we also conducted experiments on different pretrained SOTA CLIP models to verify that applying LLM2CLIP post training can significantly enhance their performance.
+4.2.1 Zero-Shot English Text-Image Retrieval
 Evaluation Dataset.
@@
 DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation.
+To facilitate observation, we additionally report the average Top-1 image-to-text (I2T) and text-to-image (T2I) accuracy across these five datasets.
 LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval.
-As shown in Table 2, LLM2CLIP fine-tuning yielded significant performance improvements, even surpassing models pretrained on large datasets like LAION2B.
-Compared to LongCLIP and JinaCLIP, our method achieved transformative gains, showcasing its efficiency in leveraging LLM knowledge to enhance pretrained CLIP models.
+In Table Table 2, we compared OpenAI’s CLIP, EVA02, and Siglip2, and conducted experiments using three types of visual models: ViT-B/16, ViT-L/14, and SoVit/14.
+Regardless of the model size or resolution shown in the table, LLM2CLIP yields a significant performance boost for both CLIP and EVA02.
+For example, under EVA02 at 224 resolution, the LLM2CLIP-60M design achieves an average Top-1 retrieval accuracy improvement of +12 and +14.8.
+Even for SigLIP2—which was pretrained on 40B data—our method still brings an average performance improvement of +9.4 and +10.6; even on the short-caption datasets Flickr and COCO, where SigLIP2 is already very familiar, improvements of +1 and +1.9 are observed.
+Comparatively, the improvements for long captions are even larger, as the inherent large window of LLMs fully leverages their advantage in understanding long and complex texts.
+Training Data Volume Analysis for LLM2CLIP.
+In Table Table 2, we also conducted experiments with 3M, 15M, and 60M training data volumes on CLIP and EVA02 using ViT-L/14.
+We found that, whether at 224 or 336 resolution, more data consistently leads to better retrieval performance, for both long and short text retrieval tasks.
+Furthermore, we observed that the 3M data experiment on EVA02 is already sufficient to yield a noticeable improvement in long text retrieval tasks; however, no clear gain was observed in short text retrieval.
+We speculate that with a small data volume, the model first needs to compensate for the loss incurred by disrupting the original cross-modal space—a fact evidenced by the significant difference between the 3M and 15M data experiments.
 4.2.2 Cross-Lingual Retrieval
-Evaluation Dataset.
-For Chinese retrieval tasks, we used FlickrCN [18] and CNCOCO [21], which are translated versions of Flickr30K and MS COCO-1K, respectively.
-LLAVA training and evaluation were conducted using the standard splits from LLAVA 1.5 [24].
-LLM Enables English CLIP to Learn Chinese Without Exposure to Chinese Data.
-To test LLM knowledge transfer, we conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text.
-As shown in Table 3, EVA02 CLIP and JinaCLIP showed near-zero performance, while LLM2CLIP achieved remarkable results, even outperforming models trained on extensive Chinese data like Wukong [12] and CN-CLIP [42].
-This demonstrates the power of LLMs in transferring knowledge to cross-modal spaces.
+LLMs naturally enable strong multilingual capabilities.
+As in Table 3, We conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text.
+We successfully endowed EVA02 with multilingual capabilities, which it previously lacked, and significantly enhanced the already robust multilingual performance of SigLip2.
 4.2.3 Zero-shot Imagenet Classification and Linear Probe Performance
@@
 This highlights the potential of LLM2CLIP for improving complex image reasoning and understanding tasks.
-4.3 Main Conclusions
-Directly Replacing CLIP’s Text Encoder with a Vanilla LLM is Harmful.
-Directly replacing the text encoder of EVA02 ViT-L/14 with Llama3-8B significantly degrades retrieval performance, as shown by the red-highlighted numbers in Table 7 and Table 6, with average performance drops of 10%-20%.
-Additionally, the CRA score of the vanilla Llama3-8B is particularly low, indicating that its output features lack sufficient discriminability for captions.
-These results clearly demonstrate that such an approach imposes a significant burden on CLIP’s learning process.
-Improving the Discriminability of LLM Output Features is Key to Integrating LLMs with CLIP.
-To enhance the discriminability of LLM output features, we applied Caption Contrastive fine-tuning.
-Llama3-8B-TC and Llama3-8B-CC represent the results of using pure text corpora(same as LLM2Vec) and a mix of text and augmented CC3M corpora from ShareCaptioner, respectively, both trained with supervised SimCSE loss.
-As shown in the Table 6, the contrastive learning on mixed caption corpora yields higher CRA scores compared to training on generic text, with noticeable improvements across almost all benchmarks.
-This difference stems from the fact that LLMs trained on caption-distributed data have stronger discriminability for caption features.
-These findings highlight the importance of text model discriminability for CLIP training and underscore the rationale behind the caption contrastive fine-tuning in LLM2CLIP.
-LLM’s Built-in Knowledge Directly Benefits CLIP.
-In Table 7, we observe that directly using LLM2CLIP leads to significant performance improvements even with short caption datasets, which is quite challenging, as CLIP has already seen most of this training data.
-However, the results from LLM2CLIP with +Llama3-8B-CC still show a clear performance gain, which can be attributed purely to the improvement in the text encoder.
-On the other hand, +Jina-Bert fails to achieve similar results, and in fact, leads to a performance decline.
-LLM Improves CLIP’s Ability to Handle Long and Dense Captions.
-Thanks to the LLM’s inherent long context windows and strong text understanding capabilities, CLIP can process long caption datasets more effectively without requiring any architectural modifications.
-Jina CLIP, for example, enables long-text input by replacing its text encoder with Jina BERT.
-To verify the LLM’s natural advantage with long texts, we compared our method to Jina CLIP to assess whether the LLM provides a better understanding of long captions.
-As shown in Table 7, when dense captions are added during training, LLM2CLIP demonstrates significantly greater performance improvements than Jina CLIP, proving that LLMs are better suited to processing long and dense captions.
-Larger LLMs Lead to Better Performance.
-LLM2CLIP explores the method of replacing CLIP’s text encoder with an LLM and continuing training.
-We thoroughly investigate the performance differences that various versions of LLMs bring to CLIP.
-In the experiments shown in Table 6, we swapped the text encoder in CLIP with different language models to explore their impact.
-We maintained the original LLM2CLIP design, with the only change being the initialization of the text encoder.
-As expected, within the same Llama family, the 8B model significantly outperfoWrms the 1B model.
-Mistral Memo also shows a slight improvement over the 8B model.
-It’s important to note that during CLIP training, the LLM’s gradients remain frozen, further proving that larger LLMs, with their richer knowledge, can deliver substantial performance improvements to CLIP.
-4.4 Ablation
-Impact of Different Data Ratios.
-With the integration of an LLM, LLM2CLIP enhances our ability to understand dense captions.
-In  Table 8, we present an ablation study that examines the impact of different ratios of dense captions versus original captions on CLIP’s training performance.
-The “Ratio” column represents the proportion of dense captions used during training.
-Overall, dense captions contribute to a noticeable improvement in performance, but more is not always better.
-For example, when using 0% dense captions, the performance on datasets like COCO and Flickr30K remains decent, but long-text benchmarks such as ShareGPT4V, Urban-1K, and DOCCI show poor results.
-Conversely, with 100% dense captions, performance on short-text retrieval benchmarks is the worst.
-It’s important to note that the dense captions we used were generated by the ShareCaptioner model, and there may be some distribution differences and noise compared to real data captions, which could somehow explain why using 100% dense data is suboptimal.
-Our findings indicate that the best performance is achieved when dense captions make up 50% to 75% of the training data.
-Impact of Different Data Size.
-As shown in  Table 2, larger training datasets consistently yield positive results for LLM2CLIP, with clear improvements in both long-text and short-text retrieval tasks.
-Our 60M dataset version has already pushed the limits of what CLIP could previously achieve.
-Even the 3M version, though lightweight, still delivers significant performance gains, demonstrating the efficiency of the LLM2CLIP approach.
-It’s worth noting that training on the 3M dataset takes only about 3 hours on an 8 H100 GPUs machine, yet results in a transformative leap in performance for a well-pretrained CLIP ViT model.
+4.3 Ablation Study: Stage-1 Caption Contrastive Fine-tuning
+Experimental Settings.
+For simplicity, we select Llama 3.1 1B as the default LLM to investigate different design choices for caption contrastive fine-tuning in this section.
+We use a subset of the CC 3M dataset for caption contrastive fine-tuning.
+Additionally, we explore augmenting the training data with the Wikitext-103 dataset [25] for MNTP training and the E5 dataset [29] as supplementary pure-text data for caption contrastive fine-tuning.
+Structural Design Ablation.
+We examine various adaptor structures in Table 8.
+Comparing rows 4 and 5, as well as rows 6 and 7, we observe that neither Linear nor Transformer adaptors significantly impact the performance of subsequent CLIP post-training.
+Therefore, we default to using no adaptor structure in Stage 1.
+Training Method Ablation.
+We conduct an ablation study of Stage-1 LLM caption contrastive fine-tuning methods in Table Table 9, observing the following key insights: 1.
+Freezing LLM gradients and relying solely on adaptors yields poor performance, indicating that enabling LoRA is essential.
+2.
+SimCSE is identified as the most critical loss function.
+MNTP does not significantly enhance SimCSE, and relying solely on MNTP without SimCSE severely deteriorates performance.
+3.
+Supervised SimCSE substantially outperforms unsupervised SimCSE.
+4.
+The performance difference between causal and bidirectional attention mechanisms is negligible.
+Ablation for different LLMs.
+In Table 6, our method achieves significant improvements over these encoders.
+The result highlighted in red shows that using the original Llama3.1-8B without CC fine-tuning yields notably poor performance, negatively affecting the original CLIP.
+4.4 Ablation Study: Stage-2 LLM2CLIP
 Adaptor Design.
-In Table 9, we explore the impact of the Adaptor and its layer count.
-We find that adding at least one layer of the Adaptor is beneficial, resulting in an average performance improvement of two points.
-However, increasing it to four layers provides minimal additional gains.
-Caption Contrastive Fine-tuning Ablation.
-In Table 10, we explore the impact of the key steps in LLM2Vec’s Caption Contrastive fine-tuning, namely MNTP and bidirectional attention.
-Through comparisons, we observe that MNTP offers almost no positive effect, while bidirectional attention provides some performance gains but is not decisive.
-Therefore, we conclude that the most significant contribution comes from the supervised SimCSE contrastive learning applied to the captions.
-5 Limitations and Future Work
-This paper focuses solely on the impact of LLMs on CLIP.
-In reality, there are many other areas where text can play a transformative role, such as diffusion models [34, 4, 39], Segment Anything [16] and GroundingDINO [25], both of which could potentially see significant performance improvements by applying our method.
-While LLMs possess open-world knowledge, aligning them with CLIP may require specific data to fully unlock their potential.
-LLM2CLIP is a method that requires only a small amount of data for fine-tuning in conjunction with a pretrained CLIP visual encoder, and we have already demonstrated significant performance gains.
-However, we did not intentionally select fine-tuning data based on specific characteristics.
-By focusing on aspects such as data distribution, length, or categories, we could further tailor the LLM to address CLIP’s limitations, allowing the LLM to act as a more comprehensive teacher for various tasks.
+As shown in Table 8, we explore adaptor structures in the Stage-2 LLM2CLIP cross-modal pre-training.
+Comparing Linear (×$\times 4$) and Transformer (×$\times 1$), their performances are relatively similar, and thus, we select the simpler Linear adaptor structure.
+Furthermore, in our layer-wise ablation of Linear adaptors, we find that increasing the adaptor size indeed improves performance.
+Specifically, adaptor performance improves progressively when transitioning from no adaptor to 1, 2 and 4-layer.
+Training Method Ablation.
+We examine several methods of integrating LLMs into CLIP post-training in Table Table 10, with implementation details described in Appendix A.
+By comparing experiments (a)-(b) and (b)-(c), we find that neither jointly training two losses nor aligning text embeddings significantly improves performance.
+However, experiment (d), which applies contrastive learning on the concatenated encoder outputs, achieves notable improvements.
+For simplicity, we do not adopt this design by default but suggest it as a promising direction for future work.
+Optimization Strategy Ablation and Efficiency Analysis of LLM2CLIP.
+As shown in  Table 11, we conducted experiments comparing methods such as LoRA training, frozen LLM online training, and offloading on fixed hardware.
+Experimental results demonstrate that, with identical computational resources, our proposed offloading method can increase the batch size by tenfold and extend training time by fourfold, ultimately achieving higher performance.
+Data Ratio Ablation.
+In Table Table 7, we experiment with various dataset compositions.
+The results show a clear trend: higher proportions of dense captions enhance performance in long-text image retrieval tasks, whereas lower dense caption ratios improve short-text retrieval performance.
+We select a 1:1 ratio by default, as it provides the most balanced improvement for both short and long texts.
+Interestingly, even training exclusively on real short captions (100%), our model still significantly boosts performance for both short and long-text retrieval tasks.
+This improvement can be attributed to the incorporation of LLM, given that real captions share the same distribution as the CLIP pre-training data, yet LLM integration still yields substantial gains.
 5 Conclusion
-Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation learning, allowing language — the carrier of human knowledge — to play a pivotal role in spreading the intelligence of LLMs across multiple modalities and applications.
-The core contributions of this work are threefold: First, we identified the primary reason why LLMs struggle to directly participate in multimodal representation learning — the lack of discriminability in their output features.
-Second, we introduced the caption contrastive fine-tuning method, which significantly improves feature discriminability, removing the biggest obstacle to utilizing LLMs in CLIP training.
-Third, we designed the LLM2CLIP training framework, which has proven to be an efficient and effective approach for significantly enhancing the performance of pretrained CLIP models.
+Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation, allowing language—the carrier of human knowledge—to play a pivotal role in disseminating the intelligence of LLMs across multiple modalities.
+This paper successfully proposes an efficient method to seamlessly integrate LLMs into CLIP post-training, demonstrating substantial performance improvements on state-of-the-art multimodal models, including CLIP, EVA02, and SigLip2.
+Appendix A Implementation details
+Details for Caption Contrastive Fine-tuning
+We default to using LoRA for training the LLM, with a rank of 16 and $\alpha=32$.
+Training is conducted on four nodes, each with 8 NVIDIA A100 40GB GPUs, with an effective batch size of 2048 for one epoch, using a maximum sequence length of 512 tokens for 68 hours.
+The Adam optimizer is used with a learning rate of $2.5\times 10^{-4}$ and a linear learning rate warm-up for the first 300 steps.
+We employ bfloat16 quantization, gradient checkpointing, and FlashAttention 2 to optimize GPU memory consumption.
+Details for LLM2CLIP Cross-modal Post-training
+We employ the AdamW optimizer with a total batch size of 4096 and set the learning rate to $1\times 10^{-5}$ with a cosine decay rate of 0.05.
+Details of multi encoder.
+When using multiple encoders, we need to align CLIP-V, CLIP-T, and LLM embeddings.
+We design a one-layer linear projector for each encoder, maintaining the dimensionality at 1024.
+For the concatenation head scenario, we design a 2048-dimensional head placed after the concatenation of (LLM, CLIP-T) embeddings and another after CLIP-V.
+Appendix B Additional results
+Training Data Ablation.
+In Table Table 12, we explore the impact of different data sources and volumes on Stage-1 LLM caption contrastive fine-tuning.
+Our findings include: 1.
+Comparing the combination of 3M Set1 and 1.5M E5 with 4.5M Set1 alone, adding supervised pure-text data from E5 does not provide practical benefits to CLIP’s cross-modal post-training performance and may even slightly degrade performance.
+2.
+For datasets sampled differently from the same distribution (Set1 and Set2), the best results are achieved when Stage 1 and Stage 2 utilize distinct datasets.
+The overall retrieval performance may depend primarily on the total amount of data seen across both stages; for instance, the performance using Set2 in Stage 1 closely matches that of combining Set1 and Set2.
+XM3600 Multilingual Results.
+Figure 4 plots radar charts for zero‑shot T2I and I2T retrieval across 36 languages on the XM3600 dataset.
+Although Siglip2 is already pre‑trained on a large, multilingual corpus, augmenting it with LLM2CLIP—leveraging a large language model to enhance linguistic understanding—still yields a substantial boost in cross‑lingual retrieval performance, even when the fine‑tuning is performed exclusively on English data.

### v5

@@
-LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation
+LLM2CLIP: Powerful Language Model Unlocks Richer Cross-Modality Representation
 Abstract
-CLIP is a foundational multimodal model that aligns image and text features into a shared representation space via contrastive learning on large-scale image-text pairs.
-Its effectiveness primarily stems from the use of natural language as rich supervision.
-Motivated by the remarkable advancements in large language models (LLMs), this work explores how LLMs’ superior text understanding and extensive open-world knowledge can enhance CLIP’s capability, especially for processing longer and more complex image captions.
-We propose an efficient post-training strategy that integrates LLMs into pretrained CLIP.
-To address the challenge posed by the autoregressive nature of LLMs, we introduce a caption-to-caption contrastive fine-tuning framework, significantly enhancing the discriminative quality of LLM outputs.
-Extensive experiments demonstrate that our approach outperforms LoRA-based methods, achieving nearly fourfold faster training with superior performance.
-Furthermore, we validate substantial improvements over state-of-the-art models such as CLIP, EVA02, and SigLip2 across various zero-shot multimodal retrieval tasks, cross-lingual retrieval tasks, and multimodal language model pretraining.
+CLIP is a seminal multimodal model that maps images and text into a shared representation space by contrastive learning on billions of image–caption pairs.
+Inspired by the rapid progress of large language models (LLMs), we investigate how the superior linguistic understanding and broad world knowledge of LLMs can further strengthen CLIP—particularly in handling long, complex captions.
+We introduce an efficient fine-tuning framework that embeds an LLM into a pretrained CLIP while incurring almost the same training cost as regular CLIP fine-tuning.
+Our method first “embedding-izes” the LLM for the CLIP setting, then couples it to the pretrained CLIP vision encoder through a lightweight adaptor trained on only a few million image–caption pairs.
+With this strategy we achieve large performance gains—without large-scale retraining—over state-of-the-art CLIP variants such as EVA02 and SigLIP-2.
+The LLM-enhanced CLIP delivers consistent improvements across a wide spectrum of downstream tasks, including linear-probe classification, zero-shot image–text retrieval with both short and long captions (in English and other languages), zero-shot/supervised image segmentation, object detection, and used as tokenizer for multimodal large-model benchmarks.
+Code & Models: https://aka.ms/llm2clip.
 1 Introduction
@@
 After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training.
-We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision.
-Figure 2: By adjusting the output space of the LLM, we enable the LLM to more effectively act as a private tutor for CLIP, allowing the fine-tuning process of LLM2CLIP to efficiently adjust the pretrained CLIP’s cross-modal space.
-CLIP [28] is one of the most influential multimodal foundational models today.
-What makes CLIP particularly powerful as a vision encoder?
-The core of its strength lies in its unprecedented ability to align visual pretraining with natural language—the carrier of human knowledge.
-Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text.
-This alignment with language is what sets CLIP apart and unlocks its vast potential.
-However, since CLIP’s introduction, large language models (LLMs) have advanced significantly.
-Models like GPT-4 [1] and Llama [6] now demonstrate remarkable language capabilities, yet these advancements have not translated into corresponding improvements in visual representation learning.
-This raises the question: Can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
+We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer textual supervision.
+CLIP (Radford et al.
+2021; Tschannen et al.
+2025) has emerged as one of the most influential cross-modal foundation models in recent years.
+Trained on hundreds of millions to tens of billions of image–text pairs via contrastive pre-training, it embeds vision and language into a shared representation space.
+As a retriever, CLIP underpins zero-shot classification, detection, segmentation, and—most notably—image–text retrieval.
+As a feature extractor, it supports a broad spectrum of cross-modal applications, from image/video understanding to text-to-image and text-to-video generation.
+For instance, Multi-modality large language models such as LLaVA (Li et al.
+2024a) and Qwen-VL (Bai et al.
+2023) rely on CLIP’s visual features, while image/video generative model like Stable-Diffusion-3 (Esser et al.
+2024) and Wan (Kong et al.
+2024) leverage its text encoder.
+Nevertheless, as we push toward broader generality and higher task complexity, the representational capacity inherited from CLIP’s original paradigm is gradually becoming inadequate.
+[…]
 The potential benefits of incorporating LLMs into CLIP are clear.
@@
 This allows them to expand on caption information during training, increasing the efficiency of the learning process.
-This paper successfully explores a efficient method of leveraging LLM capabilities in CLIP pretraining.
-LLMs, as text generation models, are primarily pretrained to capture the distribution of natural language, thus their outputs naturally lie close to textual spaces.
-However, what CLIP pretraining actually requires is a robust text embedding model.
-Previous works such as LLM2Vec [2] and NV-Embed-v2 [17] have demonstrated that fine-tuning LLMs significantly enhances their ability as text encoders.
-Nevertheless, it remains unclear how directly applying LLMs affects CLIP performance.
-Our preliminary results indicate that directly integrating LLMs into CLIP pretraining can be detrimental because LLMs struggle to differentiate effectively between image descriptions, as illustrated by example cases in Figure 3.
-Moreover, as shown in Table 1, we conducted caption-to-caption retrieval experiments on the COCO 5K test set (each image has five captions; we used the first two captions as positive pairs) and observed that large language models like Llama 3 8B achieve very low accuracy, barely distinguishing between different captions of the same image.
-Language Model Top 1 Acc.
-CLIP-L/14 25.2
-EVA02-L/14 27.1
-Llama3-8B 5.2
-Llama3.2-1B 5.6
-Llama3-8B-CC 29.5
-Llama3.2-1B-CC 29.4
-Table 1: Comparison of top-1 retrieval accuracy for various language models on the MS COCO [22] 5K test set.
-Figure 3: Real examples of top-1 results from the caption-to-caption retrieval experiment in MS COCO 5K test set. Before fine-tuning, Llama3’s results were often completely unrelated.
-We designed a caption contractive (CC) fine-tuning strategy, applying lightweight fine-tuning to the output tokens of LLM using image captioning dataset as in  Figure 2.
-The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions.
-We utilized a supervised SimCSE [8, 2] contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner [3] were treated as positive pairs, pulling them closer.
-In contrast, all other captions formed a negative sample set that the model learned to push away.
-Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 5.2% to 29.5%.
-This successful fine-tuning process enables us to more effectively harness the capabilities of LLMs for CLIP training.
-Building on fine-tuning of LLMs, we designed an efficient approach, LLM2CLIP, to enhance the performance of pretrained CLIP.
-As in  Figure 1, This method uses an LLM to further fine-tune the pretrained CLIP, allowing the LLM to impart knowledge and capabilities to CLIP’s vision encoder that were lacking during its initial pretraining.
-LLM2CLIP consists of two stages: in the first stage, we apply caption contrastive (CC) fine-tuning on the LLM with a small dataset to adjust the LLM’s output space.
-In the second stage, we fine-tune the CLIP vision encoder with the CC-fine-tuned LLM, with LLM gradients frozen, to further improve CLIP’s performance.
-The main contributions of this paper are:
-  1. 1.
-    We identify the core limitation preventing LLMs from fully leveraging their capabilities within the CLIP training framework—the insufficient separability of LLM output features.
-  2. 2.
-    We introduce a straightforward approach by applying caption-to-caption contrastive fine-tuning on the LLM’s output layer, enhancing the separability of the output space.
-This approach shows experimentally that involving LLMs in the CLIP training process can provide substantial benefits.
-  3. 3.
-    We successfully verified that LLM2CLIP can substantially enhance the performance of CLIP and EVA02 across multiple tasks, including short-text retrieval, long-text retrieval, cross-language retrieval, and multimodal large language model pretraining, through a simple post-training strategy.
-Remarkably, LLM2CLIP can further improve the short-text retrieval accuracy of state-of-the-art large-scale pretrained models like SigLip2 [33] by +1 and +1.9 points, achieve impressive gains of +14.8 and +15.8 points in long-text retrieval, and deliver significant improvements of +11.9 and +15.2 points in multilingual tasks.
+Bringing LLMs into the CLIP framework, however, raises two key challenges: 1.
+Feature separability.
+Vanilla LLM embeddings are not sufficiently discriminative for contrastive training; existing “LLM-as-embedding” recipes target pure-text tasks.
+We therefore devise a CLIP-specific embeddingization strategy for LLMs that yields far more separable caption features.
+2.
+Training cost.
+CLIP pre-training is already expensive; naively fine-tuning an LLM jointly would be prohibitive.
+We propose a lightweight fine-tuning procedure that injects LLM power into a pretrained CLIP at almost no extra cost.
+This paper introduces LLM2CLIP, an efficient fine–tuning framework that augments the feature space of a pre-trained CLIP with an embedding-tuned LLM, thereby importing LLM capability at very low cost.
+Injecting a vanilla LLM into CLIP is problematic: as shown by the cases in Figure 2, a COCO test reveals that raw LLM embeddings have poor separability for image captions.
+To remedy this, we perform caption-contrastive(CC) fine-tuning on the LLM with a set of high-quality caption datasets, revisiting several design choices—e.g.
+employing average pooling to aggregate token features, enabling bidirectional attention, and training with a supervised SimCSE contrastive loss.
+Another key challenge when introducing LLMs into CLIP is the presence of two text encoders—the original CLIP text encoder and the newly added LLM—which were never aligned during pre-training.
+We conduct experiments to explore efficient integration of the LLM into the CLIP architecture.
+As shown in Figure 1, we adopt a cost-effective solution: freeze all LLM gradients, treat its sentence embeddings as fixed features, and append a small learnable adaptor trained with the CLIP visual encoder.
+The original CLIP text encoder is discarded during both training and inference.
+Ablation studies indicate that alternative strategies—combining CLIP and LLM embeddings via triplet contrastive losses or concatenating features—provide marginal or negative gains compared to simply replacing the CLIP text encoder.
+This minimal design yields an LLM-enhanced cross-modal space while keeping compute cost nearly identical to standard CLIP fine-tuning.
+Empirical results confirm that LLM2CLIP yields substantial improvements—even with only a few million training examples—boosting the performance of the original CLIP across a variety of downstream tasks.
+Our contributions are as follows:
+  • We empirically demonstrate that injecting LLM capability into CLIP brings significant performance gains.
+  • We design (i) a caption-contrastive fine-tuning recipe that turns an LLM into a effective embedding model for CLIP, and (ii) a efficient fine-tuning method that couples this embedding with a pretrained CLIP.
+  • Extensive experiments reveal that the resulting LLM-enhanced CLIP markedly improves several state-of-the-art models—including EVA02 and SigLIP-2—on a wide spectrum of multimodal benchmarks: short/long-text and cross-lingual image retrieval, zero-shot classification, detection, segmentation, and even as the visual encoder inside LLaVA-1.5.
+Remarkably, LLM2CLIP lifts SigLIP-2’s short-caption retrieval by +1.0/+1.9, long-caption retrieval by +14.8/+15.8, and multilingual tasks by +11.9/+15.2.
+Figure 2: Real examples of top-1 results from the caption-to-caption retrieval experiment in MS COCO 5K test set. Before fine-tuning, Llama3’s results were often unrelated.
 2 Related Works
@@
 Several works have explored the integration of LLMs into CLIP.
-JinaCLIP [14] employed Jina-embeddings-v2 [9] as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts.
-Though achieving similar visual performance to EVA-CLIP [30], its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs.
-E5-V [12] took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B [18] while freezing the ViT gradient to focus on adjusting the MLLM output.
-However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP.
-MATE [10] designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs.
+JinaCLIP (Koukounas et al.
+2024a) employed Jina-embeddings-v2 (Günther et al.
+2023) as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts.
+Though achieving similar visual performance to EVA-CLIP (Sun et al.
+2023a), its text encoder is far behind ours, limiting the potential benefits from LLMs.
+MATE (Jang et al.
+2024) designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs.
 They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks.
-However, they did not recognize the critical issue we propose in this paper: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training.
-Our work aims to thoroughly explore the impact of LLMs on CLIP as a foundational model, going beyond the influence on long-text retrieval.
+However, they did not recognize the critical issue we propose: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training.
+This paper presents the first comprehensive study of an efficient strategy for incorporating LLMs into CLIP fine-tuning, thereby boosting performance.
 CLIP meets Longer Captions.
-It has been widely recognized that the quality of CLIP’s text embedding is coarse and limited to only 77 tokens.
-Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly.
-For instance, DCI [34] hired annotators to expand captions based on SAM [13]’s target hints.
-LaCLIP [7] adopted ChatGPT, Bard, and human rewriting.
-DreamLIP [41] leveraged ShareCaptioner [3] and InstructBLIP to augment 30M captions.
-Recap-DataComp-1B [21] used Llama3-trained LLAVA1.5 to extend 1B captions.
-In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones [34], splitting captions into multiple segments [7, 41], or finetuning model’s positional encoding to support longer token inputs [39].
-LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions.
-It not only resolves the limition of token length but also allows for better understanding using LLMs’ open-world knowledge.
-As a result, our model achieves a breakthrough in performance.
-Methods Resolution Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
+CLIP’s text embedding is widely recognized as coarse and limited to 77 tokens.
+Many works have attempted to extend caption length and retrain CLIP accordingly, including DCI (Urbanek et al.
+2024) with human annotation, LaCLIP (Fan et al.
+2024) using ChatGPT and Bard, DreamLIP (Zheng et al.
+2024) leveraging ShareCaptioner (Chen et al.
+2023), and Recap-DataComp-1B (Li et al.
+2024b) using Llama3-trained LLAVA1.5.
+To handle longer captions, these methods employ workarounds such as summarization (Urbanek et al.
+2024), segmentation (Fan et al.
+2024; Zheng et al.
+2024), or positional encoding finetuning (Zhang et al.
+2024a).
+In comparison, LLM2CLIP leverages LLMs as the text encoder, enabling comprehensive understanding of long and dense captions while utilizing LLMs’ open-world knowledge.
+Method Res Flickr COCO SG4V Urban DOCCI Avg
 I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
 ViT-B/16 (86M)
-ALIGN [11] 224 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
-BLIP [19] 224 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
-Long-CLIP [39] 224 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
-jina-clip-v2 [15] 224 84.4 69.8 57.1 41.9 94.5 91.1 80.4 78.0 77.6 78.2 78.8 71.8
-SigLIP [38] 224 89.1 74.7 65.8 47.8 85.8 83.5 62.8 62.4 70.1 70.6 74.7 67.8
-MetaCLIP [36] 224 85.6 70.8 59.3 41.3 90.4 86.6 68.9 63.3 70.9 71.5 75.0 66.7
-CLIP [28] 224 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
+ALIGN (Jia et al. 2021) 224 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
+BLIP (Li et al. 2022a) 224 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
+Long-CLIP (Zhang et al. 2024a) 224 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
+jina-clip-v2 (Koukounas et al. 2024b) 224 84.4 69.8 57.1 41.9 94.5 91.1 80.4 78.0 77.6 78.2 78.8 71.8
+SigLIP (Zhai et al. 2023) 224 89.1 74.7 65.8 47.8 85.8 83.5 62.8 62.4 70.1 70.6 74.7 67.8
+MetaCLIP (Xu et al. 2023) 224 85.6 70.8 59.3 41.3 90.4 86.6 68.9 63.3 70.9 71.5 75.0 66.7
+CLIP (Radford et al. 2021) 224 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
 +LLM2CLIP-15M 224 91.8 80.7 64.8 52.4 97.8 97.6 90.1 92.1 86.1 86.9 86.1 81.9
-EVA02 [31] 224 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
+EVA02 (Sun et al. 2023b) 224 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
 +LLM2CLIP-15M 224 91.2 81.4 66.0 53.9 98.5 98.5 88.2 91.3 86.8 88.5 86.1 82.7
 ViT-L/14 (307M)
-Long-CLIP [39] 224 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
-MetaCLIP [36] 224 90.1 76.4 64.5 47.1 86.4 79.5 73.4 70.0 76.5 76.7 78.2 69.9
-SAIL-L-NV2 [40] 224 87.9 75.8 62.4 48.7 94.8 94.9 81.5 80.2 76.5 78.9 80.6 75.7
-CLIP [28] 224 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
+Long-CLIP (Zhang et al. 2024a) 224 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
+MetaCLIP (Xu et al. 2023) 224 90.1 76.4 64.5 47.1 86.4 79.5 73.4 70.0 76.5 76.7 78.2 69.9
+L-NV2 (Zhang et al. 2025) 224 87.9 75.8 62.4 48.7 94.8 94.9 81.5 80.2 76.5 78.9 80.6 75.7
+CLIP (Radford et al. 2021) 224 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
 +LLM2CLIP-15M 224 93.0 83.5 67.4 56.4 98.7 98.6 91.5 94.1 88.3 90.4 87.8 84.6
-EVA02 [31] 224 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
+EVA02 (Sun et al. 2023b) 224 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
 +LLM2CLIP-3M 224 94.3 84.3 65.2 56.0 98.0 98.0 91.0 94.7 87.8 90.7 87.3 84.7
@@
 +LLM2CLIP-60M 224 95.9 85.1 71.7 58.5 99.2 99.2 95.2 97.0 90.1 92.0 90.4 86.3
-CLIP [28] 336 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
+CLIP (Radford et al. 2021) 336 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
 +LLM2CLIP-15M 336 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0 86.6 83.3
 +LLM2CLIP-60M 336 93.9 84.3 68.5 54.8 98.9 99.1 96.6 97.9 89.6 91.6 91.1 84.5
-EVA02 [31] 336 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
+EVA02 (Sun et al. 2023b) 336 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
 +LLM2CLIP-15M 336 94.7 85.1 69.7 58.5 98.8 99.2 93.5 96.2 89.2 91.3 88.0 84.8
 +LLM2CLIP-60M 336 95.9 85.4 72.4 58.8 98.8 99.2 95.5 97.3 90.7 92.6 91.0 86.8
-SigLIP [38] 384 93.7 81.4 72.0 53.9 90.1 90.4 76.3 74.4 77.8 79.5 82.0 75.9
+SigLIP (Zhai et al. 2023) 384 93.7 81.4 72.0 53.9 90.1 90.4 76.3 74.4 77.8 79.5 82.0 75.9
 SO/14 (428M)
-SigLIP [38] 224 91.0 75.2 69.8 51.8 56.1 50.0 31.5 29.1 42.3 44.8 58.1 50.2
-SigLIP2 [33] 224 93.9 82.9 72.0 55.5 90.2 87.2 75.7 74.5 77.0 78.9 81.8 75.8
+SigLIP (Zhai et al. 2023) 224 91.0 75.2 69.8 51.8 56.1 50.0 31.5 29.1 42.3 44.8 58.1 50.2
+SigLIP2 (Tschannen et al. 2025) 224 93.9 82.9 72.0 55.5 90.2 87.2 75.7 74.5 77.0 78.9 81.8 75.8
 +LLM2CLIP-60M 224 95.2 84.6 73.7 57.6 99.4 99.4 96.2 97.3 91.5 92.9 91.2 86.4
-SigLIP [38] 384 94.3 83.0 72.4 54.3 91.6 89.4 74.5 73.6 74.5 75.8 81.5 75.2
+SigLIP (Zhai et al. 2023) 384 94.3 83.0 72.4 54.3 91.6 89.4 74.5 73.6 74.5 75.8 81.5 75.2
 InternVL (6B)
-InternVL [4] 224 94.2 81.6 69.4 53.2 92.4 90.5 79.6 77.7 79.2 81.6 83.0 76.9
-Table 2: Systematic comparison of model performance on multiple datasets. The light blue row indicate LLM2CLIP post trained version.
-Models Flickr-CN [16] COCO-CN [20] XM3600 [32]
+InternVL (Chen et al. 2024) 224 94.2 81.6 69.4 53.2 92.4 90.5 79.6 77.7 79.2 81.6 83.0 76.9
+VLM2Vec (7B)
+VLM2Vec* (Jiang et al. 2024) 1344 91.6 79.8 61.3 51.5 93.9 91.0 90.9 92.0 80.5 86.0 83.6 80.1
+Table 1: Systematic comparison of model performance on multiple datasets. * means reproduced by ourselves.
+3 Methods
+3.1 Stage 1: LLM Caption Contrastive Fine-tuning
+Traditionally, an LLM’s output layer functions as a classification head that produces discrete text tokens.
+Recent studies—such as LLM2Vec (BehnamGhader et al.
+2024) and NV-EMBED-v2 (Lee et al.
+2024)—have shown that, with additional design tweaks, an LLM can be repurposed as a text-embedding model.
+Yet these efforts do not account for CLIP’s cross-modal pre-training regime.
+Here, we advance the discussion from three fronts—model architecture, training methods and training data—explicitly targeting the seamless integration of LLMs with CLIP.
+Model Architecture:
+  • Sentence Token Representation: We map the output layer features of the LLM to represent whole sentences.
+We considered two strategies: using the [EOS] token, which can attend to the entire sentence during pretraining, or employing average pooling across all output tokens.
+Empirically, average pooling performed better and is our default setting.
+  • Bidirectional Attention: Given that the generative capability is unnecessary for our encoder-focused application, we remove the attention mask from the LLM, enabling bidirectional textual relationship modeling for enhanced comprehension.
+  • Fine-tuning Parameters: To efficiently activate stronger textual comprehension capabilities within the output features of LLMs, we apply parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA).
+  • Adaptor Design: Adding an adaptor following the LLM can potentially enhance feature separability.
+We experimented with three adaptor variants: (1) a Transformer-based adaptor composed of several latent cross-attention transformer layers following NV-EMBED-v2, (2) a Linear adaptor consisting of multiple linear layers directly attached after extracting sentence tokens from LLM outputs and (3) without using adaptor.
+Training Methods:
+  • Masked Next Token Prediction (MNTP): Inspired by masked language modeling in BERT, MNTP improves feature quality by masking and predicting specific tokens.
+Unlike BERT, predictions are always made at the position preceding the masked token, aligning with LLM’s next-token prediction convention.
+While effective in pure text tasks (as shown by LLM2Vec), we find out that MNTP alone or combined with contrastive learning underperformed relative to direct contrastive fine-tuning in multimodal scenarios.
+  • Textual Contrastive Learning: Following SimCSE, we employ textual contrastive learning to enhance feature separability, pulling positive samples closer and pushing negative samples apart in feature space.
+Two variants are possible: unsupervised (using dropout augmentation) and supervised (leveraging semantic pairs from annotated data).
+We default to supervised contrastive learning due to superior performance.
+Specifically, we generate positive pairs using two distinct captions from the same image, framed by a system prompt: ”Given a caption, retrieve a similar relevant caption.”
+Training Data:
+To leverage LLM embeddings for CLIP fine-tuning, we fine-tune the LLM on caption corpora extracted from image–text datasets, aligning its embedding space with downstream multimodal tasks.
+Experiments use the Dreamlip (Zheng et al.
+2024) captions, which provide multiple captions per image, suitable for supervised textual contrastive learning.
+To preserve general language understanding, we also incorporate 1.5M pure-text pairs from Echo Embeddings (Springer et al.
+2024) into a mixed contrastive training setup.
+3.2 Stage 2: LLM2CLIP Post Fine-tuning
+We aim to perform post fine-tuning on pretrained CLIP since we believe the fundamental visual-language mapping has been well-established through CLIP’s extensive pretraining.
+By introducing an LLM, we expect to further enrich this representation space, potentially improving CLIP’s capabilities through a highly cost-effective fine-tuning step.
+CLIP inherently consists of a Vision Encoder (typically a ViT) and a relatively small Text Encoder (roughly one-third the parameters of the Vision Encoder), structured as a compact autoregressive model.
+Integrating a caption-contrastive fine-tuned LLM as an additional text encoder and finding an effective approach to enhance CLIP performance is critical.
+Here, we also elaborate on model architecture, training methods and training data.
+Model Architecture: Training CLIP typically requires large-scale datasets and substantial batch sizes, resulting in considerable computational costs.
+Therefore, we employ efficient Parameter-Efficient Fine-Tuning (PEFT) methods for this stage.
+By default, we enable full gradients for the visual encoder to facilitate learning from LLM knowledge.
+Given the large size of the LLM, we compare two strategies: (1) fine-tuning the LLM with LoRA, and (2) freezing LLM gradients and appending previously introduced Transformer or Linear adaptors after its output layer as learnable modules to obtain sentence token features.
+Method (2) offers clear advantages: it completely eliminates LLM gradient updates, significantly reducing GPU memory usage.
+Moreover, since the adaptor is placed after the LLM, text features can be precomputed and stored via offline inference, reducing inference overhead from multiple epochs to a single pass.
+As a result, this approach avoids loading the LLM into GPU memory during training, accelerating experimentation and reducing training time and memory footprint.
+By default, we adopt method (2) with an adaptor of four linear blocks.
+Detailed efficiency analysis is provided in Section 4.4.
+Training Methods: After introducing an LLM into CLIP, we are left with two text encoders.
+How should they be used most effectively?
+We tested several approaches: (a) Removing the original CLIP Text Encoder and using LLM with CLIP Vision Encoder directly for cross-modal contrastive learning.
+(b) Retaining both LLM and original Text Encoder, applying separate CLIP losses to both encoders.
+(c) Adding an additional contrastive learning task between LLM and the original Text Encoder to method (b).
+(d) Concatenating LLM and Text Encoder output features in method (b), then performing contrastive learning with CLIP Vision Encoder.
+Our extensive experiments led us to adopt method (a) for its simplicity and effectiveness.
+Training Data: With the integration of LLM, our method inherently excels at processing dense and detailed image captions.
+Datasets such as DreamLIP and Recaption, which leverage Share-Captioner (Chen et al.
+2023), InstructBLIP (Dai et al.
+2023) and LLaVA-1.5 (Liu et al.
+2023a) to generate multiple dense captions per image, can be effectively utilized.
+We explored multiple caption scenarios, employing multiple positive caption examples during contrastive training and blending real and MLLM-generated dense captions.
+Ultimately, we adopted a strategy blending real captions and dense MLLM captions at a 50% ratio.
+4 Experiments
+4.1 Default Experimental Settings
+Default Settings for Stage-1 We choose to use LLaMA 3.1 8B as the LLM integrated into CLIP pre-training, removing the causal attention mask to enable bidirectional attention.
+No additional adapters are used in caption contrastive finetune, and average pooling is employed to obtain the sentence token embedding.
+MNTP is not used by default.
+We utilize 30M DreamLIP caption data to perform caption contrastive fine-tuning on the LLM.
+Randomly sampling any two captions of an image from DreamLIP as positive pairs, we conduct supervised SimCSE training.
+Default Settings for Stage-2 After caption contrastive fine-tuning, we replace the original CLIP text encoder with the LLaMA-3.1-8B to perform cross-modal contrastive learning fine-tuning with the vision encoder.
+By default, we open the ViT gradients and freeze the LLM gradients, then add an adapter layer composed of four linear layers after the LLM.
+The design of the Adaptor is the same as FuseMix (Vouitsis et al.
+2023), using a simple inverted bottleneck MLP architecture.
+By default, we use 15M DreamLIP annotated subsets of CC 3M and CC 12M.
+For the 60M data setting, we use DreamLIP re-annotated CC 3M and CC 12M, YFCC 15M, and a 30M subset of the LAION dataset.
+For the 3M setting, we only use the CC 3M subset.
+4.2 System Comparison
+In the system comparison, we used the default settings from  Section 4.1.
+However, we also conducted experiments on different pretrained SOTA CLIP models to verify that applying LLM2CLIP post training can significantly enhance their performance.
+Zero-Shot English Text-Image Retrieval For short-text retrieval, we used the MSCOCO (Lin et al.
+2014) 5K test set and the Flickr (Young et al.
+2014) 1K test set.
+For long-text retrieval, we employed a 1K subset of ShareGPT4V (Chen et al.
+2023) and the Urban1K (Zhang et al.
+2024a) dataset from LongCLIP (Zhang et al.
+2024a), along with the DOCCI (Onoe et al.
+2024) dataset.
+The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU (Ordonez et al.
+2011), and MS COCO.
+Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions.
+DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation.
+To facilitate observation, we additionally report the average Top-1 image-to-text (I2T) and text-to-image (T2I) accuracy across these five datasets.
+LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval.
+In  Table 1, we compared OpenAI’s CLIP, EVA02, and SigLIP2, and conducted experiments using three types of visual models: ViT-B/16, ViT-L/14, and SoViT/14.
+Regardless of the model size or resolution shown in the table, LLM2CLIP yields a significant performance boost for both CLIP and EVA02.
+For example, under EVA02 at 224 resolution, the LLM2CLIP-60M design achieves an average Top-1 retrieval accuracy improvement of +12 and +14.8.
+Even for SigLIP2—which was pretrained on 40B data—our method still brings an average performance improvement of +9.4 and +10.6; even on the short-caption datasets Flickr and COCO, where SigLIP2 is already very familiar, improvements of +1 and +1.9 are observed.
+Comparatively, the improvements for long captions are even larger, as the inherent large window of LLMs fully leverages their advantage in understanding long and complex texts.
+Notably, LLM2CLIP even surpasses InternVL—which contains 6 B trainable parameters—and VLM2Vec, whose pipeline fine-tunes the entire LLM with LoRA.
+Models Flickr-CN COCO-CN XM3600
 I2T T2I I2T T2I I2T T2I
@@
 Methods Imagenet
-0-shot Linear
-CLIP L/14-336 76.6 84.8
-+LLM2CLIP 74.1 85.2
-CLIP L/14 75.5 83.9
-+LLM2CLIP 71.8 84.4
-CLIP B16 70.2 79.9
-+LLM2CLIP 67.3 80.6
-Table 4: Imagenet 0-shot classification and linear probe results.
-MODEL VQA Datasets Pope Metrics MM Benchmarks Seed Benchmarks
-VQAv2 GQA VizWiz SQA-IMG TextVQA Random Adv. Popular MME MMBench MMBench-CN LlavaBench MMVet All IMG Video
+0-shot* 0-shot Linear
+CLIP L/14-336 74.9 76.6 84.8
++LLM2CLIP 74.6 75.8 85.2
+CLIP L/14 73.7 75.5 83.9
++LLM2CLIP 73.2 74.3 84.4
+Table 3: Zero-shot classification and linear probe performance on ImageNet. *0-shot uses the class template ‘a photo of the {classname}’ only.
+Method Zero-shot Seg. mIOU OV-COCO Det. COCO val2017
+COCO-S ADE VOC City Novel Base All APbb/APseg
+EVA02 12.9 11.5 21.0 13.5 24.7 53.6 46.0 45.0/38.2
++LLM2CLIP 15.3 15.8 29.1 20.1 28.9 54.7 48.0 45.6/38.7
+Table 4: Zero-shot/supervised segmentation, open-vocabulary detection benchmarks. COCO val2017 is supervised results using CLIP’s visual encoder.
+Model VQA Pope $M$ Seed
+V2 GQA Vz SQA TV R A P MME MB MC LB MV All I V
 Llava (Paper) 78.5 62.0 50.0 66.8 58.2 87.3 86.1 84.2 1510.7 64.3 58.3 65.4 31.1 58.6 66.1 37.3
@@
 +LLM2CLIP 79.80 63.15 52.37 69.92 58.35 88.55 82.76 87.75 1505.82 68.29 60.40 62.7 34.8 60.96 68.80 38.96
-Table 5: Performance of Llava 1.5 benchmarks.
-The best results are highlighted in bold.
-We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
-Methods Flickr COCO ShareGPT4V Urban-1k DOCCI Average
-I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
-CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
-Directly Finetune(50%) 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
-bge-en-icl 89.1 78.5 58.8 49.7 95.0 95.7 77.9 87.7 73.5 79.4 78.9 78.2
-LLM2Vec-Llama-3-8B 91.1 81.4 61.5 51.9 94.5 96.4 82.1 88.6 77.7 82.6 81.4 80.2
-NV-Embed-v2 90.4 80.0 60.5 51.6 94.5 95.7 83.3 90.0 78.3 82.1 81.4 79.9
-Qwen2.5-0.5B-CC 86.2 74.9 56.0 45.1 92.6 93.2 73.7 79.0 69.5 73.0 75.6 73.0
-Llama-3.2-1B-CC 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
-Llama-3-8B-CC 90.4 80.7 62.7 51.9 96.5 96.2 84.2 89.5 83.3 86.4 83.4 80.9
-DeepSeek-R1-Distill-Llama-8B-CC 91.7 80.9 62.2 51.9 96.7 96.3 84.3 88.3 82.5 85.2 83.5 80.5
-Llama-3.1-8B-CC 92.2 81.5 63.5 52.3 97.1 96.2 86.5 89.3 84.7 85.9 84.8 81.0
-Llama3.1-8B 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
-Table 6: Ablation for using different text encoder.
-BGE-EN-ICL, LLM2Vec-Llama-3-8B, and NV-Embed-v2 are community-released text encoders derived from LLMs.
-The suffix ”-CC” indicates encoders that have undergone our caption contrastive fine-tuning method based on the respective original LLMs.
+Table 5: Performance of Llava 1.5 benchmarks. For + LLM2CLIP we replace Llava’s CLIP Vit-L/14 with our finetuned version.
+Fine-tuning Data Volume Analysis.
+In Table 1, we experimented with 3M, 15M, and 60M training data on CLIP and EVA02 using ViT-L/14.
+More data consistently improves retrieval performance for both long and short text tasks at 224 and 336 resolutions.
+With 3M data, EVA02 shows noticeable improvement in long text retrieval but minimal gain in short text retrieval, suggesting the model first compensates for disrupting the original cross-modal space before achieving broader improvements.
+Zero-Shot Multilingual Text-Image Retrieval As in Table 2, We conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text.
+We successfully endowed EVA02 with multilingual capabilities, which it previously lacked, and significantly enhanced the already robust multilingual performance of SigLIP2.
+Zero-shot & Linear Probe ImageNet Classification We evaluate our model on ImageNet using two standard protocols: (1) Zero-shot classification, where we report accuracy using the prompt template “a photo of the {classname}”, following the CLIP methodology.
+We also evaluate the average performance over 80 handcrafted prompt variants, as commonly practiced in prior work.
+(2) Linear probing, where we freeze the visual encoder and train a linear classifier using fixed hyperparameters across all experiments (batch size 1024, learning rate 0.1, momentum 0.9, weight decay 0, 50 epochs with SGD).
+As shown in Table 3, we conduct experiments using CLIP ViT-L/14 at both 224 and 336 resolutions.
+We observe that after applying LLM2CLIP, zero-shot performance on ImageNet shows a modest drop.
+However, the accuracy from linear probing improves over the original CLIP baseline.
+This suggests that while the quality of the learned visual features remains strong (as linear probing only tests CLIP visual encoder’s representation power using a supervised head), the overall multimodal alignment in the shared space may slightly deteriorate since zero-shot classification drops, especially for fine-grained category separation.
+We hypothesize that the drop in zero-shot performance could be mitigated by increasing the amount of LLM2CLIP fine-tuning data, which remains to be verified in future work.
+Nonetheless, this trade-off echoes a common pattern observed in recent CLIP fine-tuning studies, such as LongCLIP (Zhang et al.
+2024a) and CLIP-MoE (Zhang et al.
+2024b), where performance on head classes improves while long-tail categories suffer due to distribution imbalance—particularly under limited-scale fine-tuning.
+The contrast with retrieval results is instructive: retrieval benefits greatly because its vocabulary is broad and frequent, whereas classification demands uniform discriminability across all (often obscure) class nouns.
+Zero-Shot / Supervised Segmentation and Object Detection As our linear-probe results already suggest, the visual features produced by the encoder become noticeably stronger after LLM2CLIP training.
+Table 4 further confirms this on both zero-shot and fully supervised object detection and segmentation benchmarks.
+In the zero-shot setting—which relies on the text encoder—LLM2CLIP improves CLIP’s multimodal representations; in the supervised setting—where only the vision encoder is finetuned—it also enhances the purely visual, low-level feature extractor.
+We conjecture that these gains stem from the LLM’s ability to parse dense captions more accurately, especially spatial terms and object-to-object relations, and to inject that knowledge into the joint vision–language space.
+Multimodal Large Language Models Performance Following LLAVA1.5 (Liu et al.
+2023b), we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B.
+Pretraining included 558K image-caption pairs and 665K visual instruction samples.
+We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction.
+As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks.
+This highlights the potential of LLM2CLIP for improving CLIP visual encoder’s abilities for complex image reasoning and understanding.
+Stage1 Avg
+I2T T2I
+Lora, AvgPool, Bidirectional, Supervise Simcse 80.4 77.9
+Lora, AvgPool, Bidirectional, Un-supervise Simcse 59.2 57.7
+Lora, AvgPool, Bidirectional, MNTP 70.1 67.0
+Lora, AvgPool, Bidirectional, MNTP, Supervise Simcse 79.7 77.2
+Lora, AvgPool, Casual, Supervise Simcse 80.0 77.5
+Lora, EOS, Bidirectional, Supervise Simcse 80.0 77.3
+Frozen, Linear Adaptor, AvgPool, Bidirectional, Supervise Simcse 74.1 71.3
+Table 6: Ablation study on the training methods of LLM caption contrastive finetuning in Stage 1. red text indicates content that differs from the default setting.
+Method Avg
+I2T T2I
+CLIP 74.4 72.0
+Directly Finetune (50%) 74.5 72.3
+bge-en-icl 78.9 78.2
+LLM2Vec-Llama-3-8B 81.4 80.2
+NV-Embed-v2 81.4 79.9
+VLM2Vec 78.2 77.1
+bge-m3-XLM-R 65.0 63.6
+jina-v3-XLM-R 73.6 71.0
+e5 (XLM-R) 74.0 71.7
+Qwen2.5-0.5B-$C$ 75.6 73.0
+Llama-3.2-1B-$C$ 80.4 77.9
+Llama-3-8B-$C$ 83.4 80.9
+DeepSeek-R1-Distill-Llama-8B-$C$ 83.5 80.5
+Llama-3.1-8B-$C$ 84.8 81.0
+Llama3.1-8B 66.5 62.5
+Table 7: Ablation for using different text encoder. ”-CC” indicates encoders with caption contrastive fine-tuning.
+4.3 Ablation Study: Stage-1 Caption Contrastive Fine-tuning
+For simplicity, we select Llama 3.1 1B as the default LLM to investigate different design choices for caption contrastive fine-tuning in this section.
+We use a subset of the CC 3M dataset for caption contrastive fine-tuning.
+Additionally, we explore augmenting the training data with the Wikitext-103 dataset (Merity et al.
+2016) for MNTP training and the E5 dataset (Springer et al.
+2024) as supplementary pure-text data for caption contrastive fine-tuning.
+Stage1 Stage2 Avg
+I2T T2I
+78.3 75.5
+Linear($\times 1$) 79.2 76.7
+Linear($\times 2$) 80.1 76.8
+Linear($\times 4$) 80.4 77.9
+Linear($\times 4$) Linear($\times 4$) 80.5 77.7
+Transformer($\times 1$) 80.2 77.3
+Transformer($\times 1$) Transformer($\times 1$) 80.5 77.3
+Table 8: Ablation experiments for adaptor design on Stage 1 caption contrastive finetune and Stage 2 LLM2CLIP post training.
+The 4-layer Linear Adaptor has 67.1M parameters, while the single layer Transformer Adaptor has 67.6M.
+Architecture Design Ablation.
+We examine various adaptor architecture in Table A7.
+Comparing rows 4 and 5, as well as rows 6 and 7, we observe that neither Linear nor Transformer adaptors significantly impact the performance of subsequent CLIP fine-tuning.
+Therefore, we default to using no adaptor in Stage 1.
+Methods Training Loss Testing Text Encoder Average
+I2T T2I
+CLIP CL(CLIP-T, CLIP-V) CLIP-T 74.4 72.0
+Directly Finetune CL(CLIP-T, CLIP-V) CLIP-T 74.5 72.3
++LLM2CLIP a) CL(LLM, CLIP) LLM 83.9 82.1
+b) CL(LLM, CLIP-V)+ CL(CLIP-T, CLIP-V) CLIP-T 74.4 72.0
+LLM 83.6 81.8
+c) CL(LLM, CLIP-V) +CL(CLIP-T, CLIP-V) +CL(CLIP-T, LLM) CLIP-T 74.0 72.1
+LLM 83.7 81.4
+d) CL(LLM, CLIP-V) +CL(CLIP-T, CLIP-V)+ CL(Cat(CLIP-T, LLM), CLIP-V) CLIP-T 74.8 71.3
+LLM 83.8 82.3
+Cat(CLIP-T,
+LLM)
+84.7 82.8
+Table 9: LLM2CLIP Training Method Analysis.
+We experimented with various possibilities for fine-tuning between the LLM and the pretrained CLIP model’s Vision Encoder (CLIP-V) and Text Encoder (CLIP-T).
+”CL” denotes the contrastive learning loss, and ”Cat” means concatenation.
+Items a–d correspond to the a–d in Section 3.2.
+Training Method Ablation.
+We conduct an ablation study of Stage-1 LLM caption contrastive fine-tuning methods in  Table A5, observing the following key insights: 1.
+Freezing LLM gradients and relying solely on adaptors yields poor performance, indicating that enabling LoRA is essential.
+2.
+SimCSE is identified as the most critical loss function.
+MNTP does not significantly enhance SimCSE, and relying solely on MNTP without SimCSE severely deteriorates performance.
+3.
+Supervised SimCSE substantially outperforms unsupervised SimCSE.
+4.
+The performance difference between causal and bidirectional attention is negligible.
+Ablation across different LLM backbones.
+Table A6 compares a variety of text-embedding models plugged into our pipeline.
+Our approach delivers substantial gains over all contenders.
+The line highlighted in red shows that plain Llama 3.1-8B—without our caption-contrastive (CC) fine-tuning—performs very poorly and even harms the original CLIP, underscoring the necessity of the CC stage.
+We also evaluated several state-of-the-art embedding models, including bge, jina-v3, and the VLM2Vec embedding obtained from a multimodal LLM fine-tune.
+In every case, the LLM2CLIP variant still achieves the best results, confirming the effectiveness of our training recipe.
+4.4 Ablation Study: Stage-2 LLM2CLIP
+Adaptor Design.
+As shown in Table A7, we explore adaptor structures in the Stage-2 LLM2CLIP cross-modal pre-training.
+Comparing Linear ($\times 4$) and Transformer ($\times 1$), their performances are relatively similar, and thus, we select the simpler Linear adaptor structure.
+Furthermore, in our layer-wise ablation of Linear adaptors, we find that increasing the adaptor size indeed improves performance.
+Specifically, adaptor performance improves progressively when transitioning from no adaptor to 1, 2 and 4-layer.
+Training-method ablation.
+We explored multiple ways to integrate an LLM during CLIP fine-tuning (implementation details appear in Table A8.
+Comparing experiments (a)–(b) and (b)–(c) shows that neither (i) optimising two separate losses nor (ii) explicitly aligning the two text encoders brings any meaningful benefit.
+This observation motivates our choice to replace, rather than reuse, CLIP’s original text encoder in the main pipeline.
+Experiment (d)—which performs contrastive learning on the concatenated outputs of both encoders—does deliver a noticeable uplift; for the sake of simplicity, we do not adopt it as our default, but regard it as a promising future work.
+5 Conclusion
+We have revisited the problem of upgrading CLIP with large-language-model knowledge and proposed a lightweight pipeline that first turns an LLM into a discriminative caption-embedding module and then inserts it into CLIP via a cost-effective fine-tuning stage.
+With only million-scale training pairs and a compute budget essentially identical to standard CLIP fine-tuning, our method delivers sizeable gains—on top of already SOTA baselines—in image–text retrieval, image classification (zero-shot & linear probe), zero-shot/supervised detection/segmentation, and even as the visual encoder inside a multimodal LLM.
+Acknowledgments
+This work is partially supported by the National Natural Science Foundation of China (NSFC Granted No. 62276190).
+Appendix A Reproducibility Statement
+To ensure the reproducibility of our work, we are committed to making all the training and testing code, along with the datasets used in our experiments, publicly available in our code repository.
+Additionally, we will open-source our LLM2CLIP-enhanced versions of SigLip2, EVA02, and OpenAI’s CLIP to facilitate applications within the open-source community requiring high-quality embedding models.
+We will provide comprehensive documentation alongside the codebase to facilitate straightforward replication of our experiments.
+All hyperparameter settings, training configurations, and preprocessing steps will be thoroughly documented and made accessible to the research community.
+Additionally, we will include detailed environment setup instructions, dependency requirements, and step-by-step execution guidelines to minimize barriers for reproduction.
+Our commitment to open science ensures that researchers can easily validate, extend, and build upon our contributions.
+Appendix B Models and Checkpoints
+Table A2 lists all the models and checkpoints used in our experiments, ensuring transparency and facilitating reproducibility.
+It includes both vision and text encoders, with corresponding links to their publicly available implementations or weights.
+Appendix C Experiment Details
+C.1 Training Setup
+Settings for Stage-1
+We choose to use LLaMA 3.1 8B as the LLM integrated into CLIP pre-training, removing the causal attention mask to enable bidirectional attention.
+No additional adapters are used in caption contrastive finetune, and average pooling is employed to obtain the sentence token embedding.
+MNTP is not used by default.
+Additionally, to ensure the generalizability of LLM textual differentiation capabilities, we incorporated a 1.5M paired pure-text dataset curated by Echo Embeddings (Springer et al.
+2024) into our contrastive training process.Our models are trained with LoRA ($r=16$, $\alpha=32$), bfloat16 quantization, gradient checkpointing, and FlashAttention-2 for memory efficiency.
+Training uses the AdamW optimizer (learning rate $2\times 10^{-4}$, 300-step linear warm-up), sequence length 512, effective batch size 2048, and 1 epoch over 30M(1.5m paired pure-text and captions dataset) samples on 32 NVIDIA A100 GPUs.
+Settings for Stage-2
+We employed the AdamW optimizer, utilizing a total batch size of 4096, and set the learning rate to 1e-5 with a cosine decay rate of 0.05.
+For every training step we randomly sampled a single caption from the DreamLIP caption data.
+All models were trained for 4 epochs by default.
+The embedding dimension for both modalities was set to 1280.
+C.2 Evaluation
+Multilingual Retrieval
+We evaluate our models on three cross-modal retrieval benchmarks that support multilingual and cross-lingual evaluation: Flickr-CN, COCO-CN, and XM3600.
+The results are shown in Table 2 of the main text..
+Flickr-CN (Lan et al.
+2017) and COCO-CN (Li et al.
+2019) are Chinese adaptations of the popular Flickr30K and MS-COCO datasets, respectively.
+These datasets contain images paired with Chinese captions, enabling evaluation of cross-modal retrieval performance in Chinese.
+Crossmodal-3600(XM-3600)
+(Thapliyal et al.
+2022) is a geographically-diverse multilingual dataset comprising 3,600 images annotated with human-generated reference captions in 36 languages .
+The images were carefully selected from across the world, covering regions where the 36 languages are natively spoken.
+Open-Vocabulary Semantic Segmentation
+CLIP-like models enable patch-level semantic matching by aligning visual and textual representations, facilitating open-vocabulary semantic segmentation.
+Following MaskCLIP (Zhou et al.
+2022), we compute cosine similarity between patch embeddings and sentence embeddings to generate segmentation masks.
+We evaluated our models on four benchmark datasets: ADE20K (Zhou et al.
+2017), COCO-Stuff164k (Caesar et al.
+2018), VOC20 (Everingham et al.
+2015), and Cityscapes (Cordts et al.
+2016) using mean Intersection over Union (mIoU) to measure segmentation accuracy.
+The results are shown in Table 4 of the main text.
+Open-Vocabulary Object Detection
+To assess the fine-grained localization capabilities of our approach, we integrate LLM2CLIP as the backbone network for open-vocabulary detection tasks.
+Following established protocols from prior work (Wu et al.
+2023), we adopt a two-stage detection framework for evaluation.
+We conduct experiments on the OV-COCO benchmark and report box Average Precision (AP) at IoU threshold 0.5 across three categories: base classes (seen during training), novel classes (unseen during training), and all classes combined.
+The novel class performance particularly demonstrates the model’s generalization ability to previously unseen object categories.
+The strong results on novel classes, as shown in Table 4 of the main paper, demonstrate our method’s effectiveness in recognizing and localizing previously unseen objects, highlighting its robust generalization capabilities for open-vocabulary detection tasks.
+Supervised Instance Segmentation.
+To evaluate the dense prediction capabilities of our method on supervised instance segmentation tasks, we employ LLM2CLIP as the backbone for ViTDet (Li et al.
+2022b) within the Mask R-CNN (He et al.
+2017) framework.
+The models are trained on the COCO dataset using 4 A100 machines with 2 images per GPU, resulting in a total batch size of 64.
+We train for 25 epochs using the AdamW optimizer with a learning rate of 0.0001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of 0.1.
+Our approach achieves results with 45.6 mAP for object detection and 38.7 mAP for instance segmentation, as shown in Table 4 of the main text, demonstrating the effectiveness of our cross-modal pre-training in improving dense prediction tasks.
+These results highlight the model’s ability to transfer learned representations to complex downstream tasks requiring both object localization and precise pixel-level segmentation.
+Comparisons on General Multimodal Benchmarks.
+We evaluate LLM2CLIP (EVA02-L-336) as a visual feature extractor for multimodal large language models, comparing it against the standard CLIP baseline.
+Our experiments are conducted using LLaVA-v1.5-7B, which originally employs CLIP as its visual encoder.
+To ensure fair and rigorous comparison, we maintain all parameter configurations identical to the original LLaVA setup and utilize the same training data provided by the LLaVA framework.
+The comprehensive evaluation results on various multimodal understanding benchmarks((i.e., VQA (Goyal et al.
+2017) (Hudson and Manning 2019) (Gurari et al.
+2018) (Lu et al.
+2022) (Singh et al.
+2019), POPE (Li et al.
+2023b), MMBench (Liu et al.
+2024), SEED (Li et al.
+2023a)) are presented in Table 5 of the main text.
+Figure A1: Comprehensive evaluation results on cross-modal retrieval tasks.
+The radar plot illustrates the performance of different models across multiple benchmarks.
+Our methods consistently achieve higher scores across all evaluation metrics compared to baseline models.
+Appendix D More Analysis
+Quantitative Analysis of the Original LLM’s Feature Separability.
+As shown in Table A1, we conducted caption-to-caption retrieval experiments on the COCO 5K test set (each image has five captions; we used the first two captions as positive pairs).
+We observed that large language models, such as Llama 3 8B, achieve very low accuracy, struggling to distinguish effectively between different captions of the same image.
+This result highlights the necessity of the embedding-oriented training (caption contrastive fine-tuning) explored in this work for adapting LLMs specifically for the CLIP scenario.
+Language Model Top 1 Acc.
+CLIP-L/14 25.2
+EVA02-L/14 27.1
+Llama3-8B 5.2
+Llama3.2-1B 5.6
+Llama3-8B-$C$ 29.5
+Llama3.2-1B-$C$ 29.4
+Table A1: Comparison of top-1 retrieval accuracy for various language models on the MS COCO (Lin et al. 2014) 5K test set.
+Model Link
+Vision Encoders
+ViT-B/16 https://huggingface.co/openai/clip-vit-base-patch16
+ViT-L/14 https://huggingface.co/openai/clip-vit-large-patch14
+ViT-L/14-336 https://huggingface.co/openai/clip-vit-large-patch14-336
+EVA-L-224 https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02˙CLIP˙L˙psz14˙s4B.pt
+EVA-L-336 https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02˙CLIP˙L˙336˙psz14˙s6B.pt
+SigLIP SO/14 https://huggingface.co/google/siglip-so400m-patch14-224
+SigLIP2 SO/14 https://huggingface.co/google/siglip2-so400m-patch14-224
+Text Encoders
+bge-en-icl https://huggingface.co/BAAI/bge-en-icl
+LLM2Vec-Llama-3-8B https://huggingface.co/McGill-NLP/LLM2Vec-Meta-Llama-3-8B-Instruct-mntp
+NV-Embed-v2 https://huggingface.co/nvidia/NV-Embed-v2
+VLM2Vec https://huggingface.co/TIGER-Lab/VLM2Vec-LLaVa-Next
+bge-m3-XLM-R https://huggingface.co/BAAI/bge-m3
+jina-v3-XLM-R https://huggingface.co/jinaai/jina-embeddings-v3
+e5 https://huggingface.co/intfloat/multilingual-e5-large
+Qwen2.5-0.5B https://huggingface.co/Qwen/Qwen2.5-0.5B
+Llama-3.2-1B https://huggingface.co/meta-llama/Llama-3.2-1B
+Llama-3-8B https://huggingface.co/meta-llama/Meta-Llama-3-8B
+DeepSeek-R1-Distill-Llama-8B https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B
+Llama-3.1-8B https://huggingface.co/meta-llama/Llama-3.1-8B
+Table A2: Models and Checkpoints used in our experiments, including both vision and text encoders.
+(a) Multilingual zero-shot image-to-text retrieval (Recall@1)
+(b) Multilingual zero-shot text-to-image retrieval (Recall@1)
+Figure A2: Multilingual zero-shot retrieval results on Crossmodal-3600.
 Methods(MLLM’s Caption Ratio) Flickr COCO ShareGPT4V Urban-1k DOCCI Average
@@
      + LLM2CLIP (0%) 91.1 81.5 60.9 53.0 96.0 94.4 72.0 85.6 82.6 85.6 80.5 80.0
-Table 7: Ablation study on the proportion of dense captions from different MLLMs.
-Here, we explore the impact of both genuine image captions and fake dense captions generated by MLLMs on the performance of  LLM2CLIP.
-Stage 1 Stage 2 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
+Table A3: Ablation study on the proportion of dense captions from different MLLMs.
+Here, we explore the impact of both original short image captions and synthetic dense captions generated by MLLMs on the performance of LLM2CLIP.
+Model Strategy Training Hour Batchsize Flickr COCO ShareGPT4V Urban1k DOCCI Average
 I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I i2t t2i
-89.2 77.7 59.0 47.3 94.3 94.3 75.0 80.5 74.0 77.5 78.3 75.5
-Linear(×$\times 1$) 88.5 77.5 59.8 48.4 95.1 94.5 77.2 84.4 75.3 78.7 79.2 76.7
-Linear(×$\times 2$) 89.6 78.1 59.6 48.5 95.5 94.7 79.7 84.0 76.1 79.0 80.1 76.8
-Linear(×$\times 4$) 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
-Linear(×$\times 4$) Linear(×$\times 4$) 89.9 78.8 59.7 48.9 95.1 95.6 79.9 84.9 78.0 80.4 80.5 77.7
-Transformer(×$\times 1$) 88.6 78.7 59.9 48.6 95.8 95.1 79.5 84.5 77.0 79.8 80.2 77.3
-Transformer(×$\times 1$) Transformer(×$\times 1$) 89.8 78.6 58.1 49.0 96.1 95.4 81.9 85.6 76.5 78.1 80.5 77.3
-[…]
-3 Methods
-3.1 Stage 1: LLM Caption Contrastive Fine-tuning
-Traditionally, LLM output layers serve as classification heads producing text tokens.
-Prior works, such as LLM2Vec [2] and NV-EMBED-v2 [17], have explored converting LLMs into universal text encoders by applying additional designs and fine-tuning.
-Inspired by these approaches, we adapt similar concepts to the novel scenario of cross-modal pretraining.
-Our key design considerations encompass model architecture, training methods, and training data, summarized as follows:
-Model Architecture:
-  • Sentence Token Representation: We map the output layer features of the LLM to represent whole sentences.
-We considered two strategies: using the [EOS] token, which can attend to the entire sentence during pretraining, or employing average pooling across all output tokens.
-Empirically, average pooling performed better and is our default setting.
-  • Bidirectional Attention: Given that the generative capability is unnecessary for our encoder-focused application, we remove the attention mask from the LLM, enabling bidirectional textual relationship modeling for enhanced comprehension.
-  • Fine-tuning Parameters: To efficiently activate stronger textual comprehension capabilities within the output features of LLMs, we apply parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA).
-  • Adaptor Design: Adding an adaptor following the LLM can potentially enhance feature separability.
-We experimented with two adaptor variants: (1) a Transformer-based adaptor composed of several latent cross-attention transformer layers following NV-EMBED-v2, and (2) a Linear adaptor consisting of multiple linear layers directly attached after extracting sentence tokens from LLM outputs.
-However, neither approach significantly boosted CLIP performance during embedding fine-tuning.
-Consequently, we opted not to use adaptors during the LLM caption contrastive fine-tuning phase for simplicity.
-Training Methods:
-  • Masked Next Token Prediction (MNTP): Inspired by masked language modeling in BERT, MNTP improves feature quality by masking and predicting specific tokens.
-Unlike BERT, predictions are always made at the position preceding the masked token, aligning with LLM’s next-token prediction convention.
-While effective in pure text tasks (as shown by LLM2Vec), MNTP alone or combined with contrastive learning underperformed relative to direct contrastive fine-tuning in multimodal scenarios.
-  • Textual Contrastive Learning: Following SimCSE, we employ textual contrastive learning to enhance feature separability, pulling positive samples closer and pushing negative samples apart in feature space.
-Two variants are possible: unsupervised (using dropout augmentation) and supervised (leveraging semantic pairs from annotated data).
-We default to supervised contrastive learning due to superior performance.
-Specifically, we generate positive pairs using two distinct captions from the same image, framed by a system prompt: ”Given a caption, retrieve a similar relevant caption.”
-Training Data: We conducted experiments with Dreamlip [41]’s caption dataset, naturally containing multiple captions per image, making it ideal for supervised textual contrastive learning.
-Additionally, to ensure the generalizability of LLM textual differentiation capabilities, we incorporated a 1.5M paired pure-text dataset curated by Echo Embeddings [29] into our contrastive training process.
-89213
-3.2 Stage 2: LLM2CLIP Post Fine-tuning
-We aim to perform post fine-tuning on pretrained CLIP since we believe the fundamental visual-language mapping has been well-established through CLIP’s extensive pretraining.
-By introducing an LLM, we expect to further enrich this representation space, potentially improving CLIP’s capabilities through a highly cost-effective post-training step.
-CLIP inherently consists of a Vision Encoder (typically a ViT) and a relatively small Text Encoder (roughly one-third the parameters of the Vision Encoder), structured as a compact autoregressive model.
-Integrating a caption-contrastive fine-tuned LLM as an additional text encoder and determining the optimal approach to enhance CLIP performance is critical.
-Here, we elaborate on model architecture, training methods, and training data.
-Model Architecture:
-  • Fine-tuning Parameters and Adaptor Design: Training CLIP typically requires large-scale datasets and significant batch sizes, resulting in considerable computational costs.
-Therefore, we aim to employ efficient Parameter-Efficient Fine-Tuning (PEFT) methods for this post-training stage.
-By default, we enable full gradients for the visual encoder to facilitate learning from LLM knowledge.
-Given the large size of the LLM, we compare two fine-tuning strategies: (1) directly fine-tuning the LLM using LoRA, and (2) freezing the LLM gradients and appending previously introduced Transformer or Linear adaptors after its output layer as learnable modules to obtain the sentence token features.
-Method (2) offers distinct advantages, including completely eliminating gradient updates for the LLM, thus significantly reducing GPU memory consumption.
-Moreover, since the adaptor is placed after the LLM, we can even precompute and store text features through offline inference, thereby reducing the inference overhead from multiple epochs to just a single pass.
-Consequently, this approach eliminates the need to load the LLM into GPU memory during training, greatly accelerating experiment iteration and training time, and drastically reducing GPU memory footprint.
-By default, we adopt method (2) with an adaptor consisting of four linear blocks.
-Detailed efficiency analysis will be presented in Section 4.4.
-Training Methods: We tested several approaches:
-  • (a) Removing the original CLIP Text Encoder and using LLM with CLIP Vision Encoder directly for cross-modal contrastive learning.
-  • (b) Retaining both LLM and original Text Encoder, applying separate CLIP losses to both encoders.
-  • (c) Adding an additional contrastive learning task between LLM and the original Text Encoder to method (b).
-  • (d) Concatenating LLM and Text Encoder output features in method (b), then performing contrastive learning with CLIP Vision Encoder.
-Extensive experiments (see Table 10) led us to adopt method (a) for its simplicity and effectiveness.
-Training Data: With the integration of LLM, our method inherently excels at processing dense and detailed image captions.
-Datasets such as DreamLIP and Recaption, which leverage Share-Captioner [3], InstructBLIP [5] and LLaVA-1.5 [23] to generate multiple dense captions per image, can be effectively utilized.
-We explored multiple caption scenarios, employing multiple positive caption examples during contrastive training and blending real and MLLM-generated dense captions.
-Ultimately, we adopted a strategy blending real captions and dense MLLM captions at a 50% ratio.
-Stage 1 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
+Llama-3.1-8B LLM LoRA 17 768 90.0 80.9 60.9 50.6 97.5 96.9 88.9 90.8 82.0 84.3 83.9 80.7
+LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
+Llama-3.1-8B-$C$ LLM LoRA 17 704 90.5 81.8 61.5 52.4 98.1 98.2 91.0 92.8 85.7 87.4 85.4 82.5
+LLM Frozen + Linear Adaptor 5.5 4096 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
+LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 93.2 82.2 64.2 54.0 98.2 98.3 87.8 93.1 86.2 88.8 85.9 83.3
+Table A4: Efficiency study of LLM2CLIP.
+We use two nodes, each equipped with 8 A100 40GB GPUs, and conduct experiments by maximizing GPU memory utilization to achieve the largest possible batch size.
+We report training hours, batch size, and retrieval accuracy for each experiment.
+During offloading, we precompute text captions offline into embeddings using the LLM.
+This approach is a one-time preprocessing step that effectively reduces the experimental overhead.
+Stage1 Flickr COCO ShareGPT4V Urban-1k DOCCI Avg
 I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
@@
 Frozen, Linear Adaptor, AvgPool, Bidirectional, Supervise Simcse 85.5 75.5 54.1 45.3 92.0 93.5 72.3 75.5 66.4 66.9 74.1 71.3
-Table 9: Ablation study on the training methods of LLM caption contrastive finetuning in Stage 1.
+Table A5: Ablation study on the training methods of LLM caption contrastive finetuning in Stage 1.
 Rows with a light blue background represent our default setting, and red text indicates content that differs from the default setting.
+Method Flickr COCO ShareGPT4V Urban-1k DOCCI Avg
+I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
+CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
+Directly Finetune (50%) 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
+bge-en-icl 89.1 78.5 58.8 49.7 95.0 95.7 77.9 87.7 73.5 79.4 78.9 78.2
+LLM2Vec-Llama-3-8B 91.1 81.4 61.5 51.9 94.5 96.4 82.1 88.6 77.7 82.6 81.4 80.2
+NV-Embed-v2 90.4 80.0 60.5 51.6 94.5 95.7 83.3 90.0 78.3 82.1 81.4 79.9
+bge-m3-XLM-R 80.7 70.3 51.4 42.0 84.5 86.5 56.8 63.4 51.7 55.9 65.0 63.6
+jina-v3-XLM-R 84.4 73.7 56.3 45.6 90.0 90.9 70.8 74.1 66.7 70.6 73.6 71.0
+e5 (XLM-R) 86.4 75.3 56.4 46.2 88.4 88.5 71.1 77.3 67.5 71.2 74.0 71.7
+VLM2VEC 91.6 79.8 61.3 51.5 93.9 91.0 90.9 92.0 80.5 86.0 83.6 80.1
+VLM2VEC (finetune) 90.2 79.3 60.0 50.1 89.8 91.4 76.9 85.8 74.1 78.7 78.2 77.1
+Qwen2.5-0.5B-$C$ 86.2 74.9 56.0 45.1 92.6 93.2 73.7 79.0 69.5 73.0 75.6 73.0
+Llama-3.2-1B-$C$ 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
+Llama-3-8B-$C$ 90.4 80.7 62.7 51.9 96.5 96.2 84.2 89.5 83.3 86.4 83.4 80.9
+DeepSeek-R1-Distill-Llama-8B-$C$ 91.7 80.9 62.2 51.9 96.7 96.3 84.3 88.3 82.5 85.2 83.5 80.5
+Llama-3.1-8B-$C$ 92.2 81.5 63.5 52.3 97.1 96.2 86.5 89.3 84.7 85.9 84.8 81.0
+Llama3.1-8B 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
+Table A6: Ablation for using different text encoder.
+BGE-EN-ICL, LLM2Vec-Llama-3-8B, and NV-Embed-v2 are community-released text encoders derived from LLMs.
+The suffix ”-CC” indicates encoders that have undergone our caption contrastive fine-tuning method based on the respective original LLMs.
+Stage 1 Stage 2 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
+I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
+89.2 77.7 59.0 47.3 94.3 94.3 75.0 80.5 74.0 77.5 78.3 75.5
+Linear($\times 1$) 88.5 77.5 59.8 48.4 95.1 94.5 77.2 84.4 75.3 78.7 79.2 76.7
+Linear($\times 2$) 89.6 78.1 59.6 48.5 95.5 94.7 79.7 84.0 76.1 79.0 80.1 76.8
+Linear($\times 4$) 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
+Linear($\times 4$) Linear($\times 4$) 89.9 78.8 59.7 48.9 95.1 95.6 79.9 84.9 78.0 80.4 80.5 77.7
+Transformer($\times 1$) 88.6 78.7 59.9 48.6 95.8 95.1 79.5 84.5 77.0 79.8 80.2 77.3
+Transformer($\times 1$) Transformer($\times 1$) 89.8 78.6 58.1 49.0 96.1 95.4 81.9 85.6 76.5 78.1 80.5 77.3
+Table A7: Ablation experiments for adaptor design on Stage 1 caption contrastive finetune and Stage 2 LLM2CLIP post training.
+In parentheses, ($\times\textit{l}$) denotes an adaptor with $\times\textit{l}$ layers.
+The 4-layer Linear Adaptor has 67.1M parameters, while the single layer Transformer Adaptor has 67.6M parameters.
+This experiment is conducted using llama 3.1 1B and OpenAI CLIP ViT-L 224, with CC 3M data utilized in both Stage 1 and Stage 2.
+The light blue row indicate our default setting.
 Methods Training Loss Testing Text Encoder Flickr COCO ShareGPT4V Urban1k DOCCI Average
@@
 Cat(CLIP-T,LLM) 91.3 82.4 62.9 53.3 97.8 97.4 86.6 92.9 84.9 88.2 84.7 82.8
-Table 10: LLM2CLIP Training Method Analysis.
-We experimented with various possibilities for post-training between the LLM and the pretrained CLIP model’s Vision Encoder (CLIP-V) and Text Encoder (CLIP-T).
+Table A8: LLM2CLIP Training Method Analysis.
+We experimented with various possibilities for fine-tuning between the LLM and the pretrained CLIP model’s Vision Encoder (CLIP-V) and Text Encoder (CLIP-T).
 In the figure, ”CL” denotes the contrastive learning loss, and ”Cat” stands for concatenation.
@@
 In this experiment, we performed zero-shot image-text retrieval tests on each text encoder, including the new feature projection space formed after concatenation.
-Model Strategy Training Hour Batchsize Flickr COCO ShareGPT4V Urban1k DOCCI Average
-I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I i2t t2i
-Llama-3.1-8B LLM LoRA 17 768 90.0 80.9 60.9 50.6 97.5 96.9 88.9 90.8 82.0 84.3 83.9 80.7
-LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
-Llama-3.1-8B-CC LLM LoRA 17 704 90.5 81.8 61.5 52.4 98.1 98.2 91.0 92.8 85.7 87.4 85.4 82.5
-LLM Frozen + Linear Adaptor 5.5 4096 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
-LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 93.2 82.2 64.2 54.0 98.2 98.3 87.8 93.1 86.2 88.8 85.9 83.3
-Table 11: Efficiency study of LLM2CLIP.
-We use two nodes, each equipped with 8 A100 40GB GPUs, and conduct experiments by maximizing GPU memory utilization to achieve the largest possible batch size.
-We report training hours, batch size, and retrieval accuracy for each experiment.
-During offloading, we precompute text captions offline into embeddings using the LLM.
-This approach is a one-time preprocessing step that effectively reduces the experimental overhead.
-4 Experiments
-4.1 Default Experimental Settings
-Default Settings for Caption Contrastive Fine-tuning We choose to use LLaMA 3.1 8B as the LLM integrated into CLIP pre-training, removing the causal attention mask to enable bidirectional attention.
-No additional adapters are used in caption contrastive finetune, and average pooling is employed to obtain the sentence token embedding.
-MNTP is not used by default.
-We utilize 30M DreamLIP caption data to perform caption contrastive fine-tuning on the LLM.
-Randomly sampling any two captions of an image from DreamLIP as positive pairs, we conduct supervised SimCSE training.
-Default Settings for LLM2CLIP Cross-modal Post-training After caption contrastive fine-tuning, we replace the original CLIP text encoder with the LLaMA 3.1 8B model to perform cross-modal contrastive learning post-training with the vision encoder.
-By default, we open the ViT gradients and freeze the LLM gradients, then add an adapter layer composed of four linear layers after the LLM.
-The design of the Adaptor is the same as FuseMix [35], using a simple inverted bottleneck MLP architecture.
-By default, we use 15M DreamLIP annotated subsets of CC 3M and CC 12M.
-For the 60M data setting, we use DreamLIP re-annotated CC 3M and CC 12M, YFCC 15M, and a 30M subset of the LAION dataset.
-For the 3M setting, we only use the CC 3M subset.
-We summerize our default experiments setting and implementation details in Appendix A.
-4.2 System Comparison
-In the system comparison, we used the default settings from  Section 4.1 throughout.
-However, we also conducted experiments on different pretrained SOTA CLIP models to verify that applying LLM2CLIP post training can significantly enhance their performance.
-4.2.1 Zero-Shot English Text-Image Retrieval
-Evaluation Dataset.
-For short-text retrieval, we used the MSCOCO [22] 5K test set and the Flickr [37] 1K test set.
-For long-text retrieval, we employed datasets from LongCLIP, including a 1K subset of ShareGPT4V [3], the Urban1K [39] dataset, and the DOCCI [26] dataset.
-The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU [27], and MS COCO.
-Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions.
-DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation.
-To facilitate observation, we additionally report the average Top-1 image-to-text (I2T) and text-to-image (T2I) accuracy across these five datasets.
-LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval.
-In Table Table 2, we compared OpenAI’s CLIP, EVA02, and Siglip2, and conducted experiments using three types of visual models: ViT-B/16, ViT-L/14, and SoVit/14.
-Regardless of the model size or resolution shown in the table, LLM2CLIP yields a significant performance boost for both CLIP and EVA02.
-For example, under EVA02 at 224 resolution, the LLM2CLIP-60M design achieves an average Top-1 retrieval accuracy improvement of +12 and +14.8.
-Even for SigLIP2—which was pretrained on 40B data—our method still brings an average performance improvement of +9.4 and +10.6; even on the short-caption datasets Flickr and COCO, where SigLIP2 is already very familiar, improvements of +1 and +1.9 are observed.
-Comparatively, the improvements for long captions are even larger, as the inherent large window of LLMs fully leverages their advantage in understanding long and complex texts.
-Training Data Volume Analysis for LLM2CLIP.
-In Table Table 2, we also conducted experiments with 3M, 15M, and 60M training data volumes on CLIP and EVA02 using ViT-L/14.
-We found that, whether at 224 or 336 resolution, more data consistently leads to better retrieval performance, for both long and short text retrieval tasks.
-Furthermore, we observed that the 3M data experiment on EVA02 is already sufficient to yield a noticeable improvement in long text retrieval tasks; however, no clear gain was observed in short text retrieval.
-We speculate that with a small data volume, the model first needs to compensate for the loss incurred by disrupting the original cross-modal space—a fact evidenced by the significant difference between the 3M and 15M data experiments.
-4.2.2 Cross-Lingual Retrieval
-LLMs naturally enable strong multilingual capabilities.
-As in Table 3, We conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text.
-We successfully endowed EVA02 with multilingual capabilities, which it previously lacked, and significantly enhanced the already robust multilingual performance of SigLip2.
-4.2.3 Zero-shot Imagenet Classification and Linear Probe Performance
-Evaluation Dataset.
-We conduct zero-shot retrieval testing on Imagenet, following the CLIP method.
-We also perform linear probing, fine-tuning the model with the exact same hyperparameters: batch size of 1024, learning rate of 0.1, momentum 0.9, weight decay 0 and total 50 epochs using SGD.
-Imagenet Performance Better In Linear Probe.
-Our method shows a slight drop in performance on Imagenet for zero-shot classification.
-We hypothesize that this is due to insufficient alignment of the data, resulting in some categories having very low word frequency, as well as the LLM’s limited ability to distinguish short captions.
-However, in the linear probe experiment, we observe that after supplementing the category information, LLM2CLIP actually leads to a greater improvement.
-This demonstrates that LLM2CLIP can indeed provide better visual features.
-One of the current limitations is that scaling the dataset to a larger size might be necessary to resolve the zero-shot classification issue.
-4.2.4 Multimodal Large Language Models Performance
-Evaluation Dataset.
-Following LLAVA1.5 [24], we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B.
-Pretraining included 558K image-caption pairs and 665K visual instruction samples.
-We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction.
-LLM2CLIP Improves VLLM Performance.
-As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks.
-This highlights the potential of LLM2CLIP for improving complex image reasoning and understanding tasks.
-4.3 Ablation Study: Stage-1 Caption Contrastive Fine-tuning
-Experimental Settings.
-For simplicity, we select Llama 3.1 1B as the default LLM to investigate different design choices for caption contrastive fine-tuning in this section.
-We use a subset of the CC 3M dataset for caption contrastive fine-tuning.
-Additionally, we explore augmenting the training data with the Wikitext-103 dataset [25] for MNTP training and the E5 dataset [29] as supplementary pure-text data for caption contrastive fine-tuning.
-Structural Design Ablation.
-We examine various adaptor structures in Table 8.
-Comparing rows 4 and 5, as well as rows 6 and 7, we observe that neither Linear nor Transformer adaptors significantly impact the performance of subsequent CLIP post-training.
-Therefore, we default to using no adaptor structure in Stage 1.
-Training Method Ablation.
-We conduct an ablation study of Stage-1 LLM caption contrastive fine-tuning methods in Table Table 9, observing the following key insights: 1.
-Freezing LLM gradients and relying solely on adaptors yields poor performance, indicating that enabling LoRA is essential.
-2.
-SimCSE is identified as the most critical loss function.
-MNTP does not significantly enhance SimCSE, and relying solely on MNTP without SimCSE severely deteriorates performance.
-3.
-Supervised SimCSE substantially outperforms unsupervised SimCSE.
-4.
-The performance difference between causal and bidirectional attention mechanisms is negligible.
-Ablation for different LLMs.
-In Table 6, our method achieves significant improvements over these encoders.
-The result highlighted in red shows that using the original Llama3.1-8B without CC fine-tuning yields notably poor performance, negatively affecting the original CLIP.
-4.4 Ablation Study: Stage-2 LLM2CLIP
-Adaptor Design.
-As shown in Table 8, we explore adaptor structures in the Stage-2 LLM2CLIP cross-modal pre-training.
-Comparing Linear (×$\times 4$) and Transformer (×$\times 1$), their performances are relatively similar, and thus, we select the simpler Linear adaptor structure.
-Furthermore, in our layer-wise ablation of Linear adaptors, we find that increasing the adaptor size indeed improves performance.
-Specifically, adaptor performance improves progressively when transitioning from no adaptor to 1, 2 and 4-layer.
-Training Method Ablation.
-We examine several methods of integrating LLMs into CLIP post-training in Table Table 10, with implementation details described in Appendix A.
-By comparing experiments (a)-(b) and (b)-(c), we find that neither jointly training two losses nor aligning text embeddings significantly improves performance.
-However, experiment (d), which applies contrastive learning on the concatenated encoder outputs, achieves notable improvements.
-For simplicity, we do not adopt this design by default but suggest it as a promising direction for future work.
-Optimization Strategy Ablation and Efficiency Analysis of LLM2CLIP.
-As shown in  Table 11, we conducted experiments comparing methods such as LoRA training, frozen LLM online training, and offloading on fixed hardware.
-Experimental results demonstrate that, with identical computational resources, our proposed offloading method can increase the batch size by tenfold and extend training time by fourfold, ultimately achieving higher performance.
+Detailed Results for X3600 Multilingual Task.
+In Figure 2(a) and Figure 2(b), we visualize the I2T and T2I results for 36 different languages.
+It can be observed that our method outperforms the original SigLip2 text encoder, despite the latter being trained with 10% of its training samples from 12 billion alt-texts covering 109 languages.
+In contrast, our method LLM2CLIP was trained exclusively on 15 million English-only samples.
+This clearly demonstrates the inherent multilingual advantage gained by incorporating an LLM.
 Data Ratio Ablation.
-In Table Table 7, we experiment with various dataset compositions.
+In  Table A3, we experiment with various dataset compositions.
 The results show a clear trend: higher proportions of dense captions enhance performance in long-text image retrieval tasks, whereas lower dense caption ratios improve short-text retrieval performance.
@@
 This improvement can be attributed to the incorporation of LLM, given that real captions share the same distribution as the CLIP pre-training data, yet LLM integration still yields substantial gains.
-5 Conclusion
-Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation, allowing language—the carrier of human knowledge—to play a pivotal role in disseminating the intelligence of LLMs across multiple modalities.
-This paper successfully proposes an efficient method to seamlessly integrate LLMs into CLIP post-training, demonstrating substantial performance improvements on state-of-the-art multimodal models, including CLIP, EVA02, and SigLip2.
-Appendix A Implementation details
-Details for Caption Contrastive Fine-tuning
-We default to using LoRA for training the LLM, with a rank of 16 and $\alpha=32$.
-Training is conducted on four nodes, each with 8 NVIDIA A100 40GB GPUs, with an effective batch size of 2048 for one epoch, using a maximum sequence length of 512 tokens for 68 hours.
-The Adam optimizer is used with a learning rate of $2.5\times 10^{-4}$ and a linear learning rate warm-up for the first 300 steps.
-We employ bfloat16 quantization, gradient checkpointing, and FlashAttention 2 to optimize GPU memory consumption.
-Details for LLM2CLIP Cross-modal Post-training
-We employ the AdamW optimizer with a total batch size of 4096 and set the learning rate to $1\times 10^{-5}$ with a cosine decay rate of 0.05.
-Details of multi encoder.
-When using multiple encoders, we need to align CLIP-V, CLIP-T, and LLM embeddings.
-We design a one-layer linear projector for each encoder, maintaining the dimensionality at 1024.
-For the concatenation head scenario, we design a 2048-dimensional head placed after the concatenation of (LLM, CLIP-T) embeddings and another after CLIP-V.
-Appendix B Additional results
-Training Data Ablation.
-In Table Table 12, we explore the impact of different data sources and volumes on Stage-1 LLM caption contrastive fine-tuning.
-Our findings include: 1.
-Comparing the combination of 3M Set1 and 1.5M E5 with 4.5M Set1 alone, adding supervised pure-text data from E5 does not provide practical benefits to CLIP’s cross-modal post-training performance and may even slightly degrade performance.
-2.
-For datasets sampled differently from the same distribution (Set1 and Set2), the best results are achieved when Stage 1 and Stage 2 utilize distinct datasets.
-The overall retrieval performance may depend primarily on the total amount of data seen across both stages; for instance, the performance using Set2 in Stage 1 closely matches that of combining Set1 and Set2.
-XM3600 Multilingual Results.
-Figure 4 plots radar charts for zero‑shot T2I and I2T retrieval across 36 languages on the XM3600 dataset.
-Although Siglip2 is already pre‑trained on a large, multilingual corpus, augmenting it with LLM2CLIP—leveraging a large language model to enhance linguistic understanding—still yields a substantial boost in cross‑lingual retrieval performance, even when the fine‑tuning is performed exclusively on English data.
+However, when training exclusively on synthesized long texts, although retrieval performance on long-text tasks can be significantly improved, performance on short-text retrieval tasks tends to decline.
+We hypothesize two possible reasons for this phenomenon: First, relying entirely on synthesized data might introduce significant biases in the training samples.
+Second, when the learned feature space is constructed exclusively from long texts, the model may prioritize global semantic matching and overlook fine-grained information such as specific words, potentially negatively affecting short-text retrieval.
+Therefore, we believe that for contrastive learning tasks focusing solely on long texts, it is important to consider including both global and local levels of text granularity during training like FLIP (Yao et al.
+2021), or introducing sufficient hard examples to encourage the model to attend to low-level semantic details.
+Optimization Strategy Ablation and Efficiency Analysis of LLM2CLIP.
+As shown in  Table A4, we conducted experiments comparing methods such as LoRA training, frozen LLM online training, and offloading on fixed hardware.
+Experimental results demonstrate that, with identical computational resources, our proposed offloading method can increase the batch size by tenfold and extend training time by fourfold, ultimately achieving higher performance.
+Complete Tables and Additional Visualizations.
+In Tables Table A5Table A6Table A7, and Table A8, we provide additional detailed results corresponding to Table 6, 7, 8, and 9 in the main text, facilitating a clear analysis of the differences in retrieval performance between long and short texts.
+The conclusions remain consistent with those presented in the main text.
+Additionally, Figure A1 illustrates radar plots highlighting the performance improvements achieved by LLM2CLIP on SigLip2 and EVA02.